A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1920–1929

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

χ^{2} [176, 398, 843, 177]. n [439]. P [176].

1 [156]. 1832 [38]. 1856 [850]. 1875 [779]. 19 [244, 338, 121]. 1913 [54]. 1914 [287, 187]. 1917 [898]. 1918 [668, 424, 269]. 1920 [156, 182]. 1921 [157, 156, 641, 185]. 1922 [382]. 1923 [425, 368, 452]. 1924 [828, 592, 610, 662, 514]. 1925 [661]. 1926 [607]. 1928 [851].

2 [78, 236, 334, 418, 503, 603, 704, 794, 890]. 20 [90, 444, 445, 266]. 21 [240, 241, 345, 775, 448]. 22 [407, 560, 367]. 23 [426].

50 [666].

9 [78].

Abglaises [805]. Abroad [89, 613]. Academic [341, 685]. Accidents [27, 676, 605]. Account [178]. Accounts [396, 760, 142]. Achille [453, 87]. Achinstein [829]. Activity [652]. Acts [283]. Actuarial [286, 361]. Actuary [444]. Actuels [220]. Acworth [472, 489, 560, 59, 494]. Adam [806, 34]. Adams [558]. Additions [95, 23, 45, 94, 171, 114, 132, 150, 170, 256, 192, 210, 231, 255, 351, 279, 303, 326, 350, 434, 373, 393, 412, 433, 523, 459, 477, 497, 522, 622, 546, 571, 598, 621, 724, 650, 673, 697, 723, 812, 748, 767, 788, 811, 904, 835, 859, 881, 903].

Adelphi [78, 236, 334, 418, 503, 603, 704, 794, 890]. Aditions [70]. Adjustment [530]. Administration [203, 592, 295, 226]. Administrative [549]. Advancement [185]. Adventure [584]. Advertising [272, 737]. Affairs [403]. Africa [852]. After [864, 162, 240, 33, 86, 287, 275]. Against [202]. Age [632, 331, 752]. Agger [11]. Agraire [227]. Agrarian [898]. Agricole [452]. Agricultural [354, 250, 740, 382, 383, 508, 509, 585, 426, 461]. Agriculture [226, 173, 681]. Al [264]. Alan [668]. Albert [198, 471]. Alexander [178, 596]. Alexandrovitch [596]. Alf [336]. Alfred [242, 901, 312, 679, 286, 31, 554, 457]. Alfredo [381, 102]. Algernon [281, 298]. alle [381]. Allen [609]. Allowances [583, 565]. Ally [425]. Alson [630]. Altri [239]. Alzada [895]. Amalgamation [405, 610]. Ambedkar [581]. America [782, 585]. American [202, 428, 741, 275, 321, 338, 396, 711, 848, 538]. among [369]. Amoroso [218]. Amount [338]. Analysis [356, 89, 817, 777, 464, 454, 425, 154, 781]. Anderson [55]. Andrews [274, 429]. Angas [221]. Angell [707]. Anglais [242]. Annals [686, 534]. Annuaire [452]. Annual [95, 171, 256, 351, 434, 523, 622, 724, 812, 904, 285, 79, 78, 157, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890]. Annuitants [444, 445]. Antonio [443]. Any [888]. Appendices [103]. Appendix [733]. Application [821, 755, 199, 556, 378, 177, 713]. Applications [676]. Applicazioni [381]. Applied [487]. Appliquee [360, 158]. Appointed [203]. Approximation [793]. Arable [887]. Arithmetic [439]. Armand [360, 158]. Arnaune [288]. Arne [420]. Arnold [539]. Arrays [265]. Art [541]. Arte [381]. Arthur [77, 741, 54, 666, 734, 14, 662, 182, 558]. Articles [22, 44, 69, 93, 113, 131, 149, 169, 191, 209, 230, 254, 278, 302, 325, 349, 372, 392, 411, 432, 458, 476, 496, 521, 545, 570, 597, 620, 649, 672, 696, 722, 747, 766, 787, 810, 834, 858, 880, 902]. Arts [237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890]. Asher [829]. Ashley [13, 638]. Ashton [516]. Aspects [243, 728, 104, 549]. Association [699, 177]. Asymptotic [439]. Athelstane [544]. Attacks [27]. Auguste [288]. Augustus [857]. Austin [582]. Australia [781, 17]. Austrian [513]. Automobile [802]. aux [227]. Axel [446].

B

[518, 668, 274, 581, 709, 806, 188, 166, 807, 55, 221, 386, 510, 39, 585, 140, 639, 517, 488, 558]. Back [837, 861, 905]. Bahadurji [616]. Baines [544]. Balance [512, 847]. Bank [105, 448]. Banker [141]. Bankers [387]. Banking [55, 342, 339, 610, 851, 657, 514, 11, 35, 611]. Banks [777, 801]. Barbara [38]. Barker [57]. Barnett [743]. Barometer [366]. Barometri [869]. Barton [780]. Based [483, 13]. Bases [517]. Bateman [901]. Batten [635]. Bauernbewegung [898]. be [480]. Bearing [421, 480, 159]. Beatrice [223, 58, 294]. Beckhart [510]. Bedeutung [336]. Beef [588]. Before [162, 33]. Behaviour [804]. Being [687, 90, 294]. Bellerby [559]. Benefits [361]. Benham [781]. Benjamin [429]. Benn [536]. Beoefening [844]. Bernard [294, 245]. Berridge [367]. Bertillon [253]. Bertram [582]. Bertrand [778]. Between [551, 398, 888, 525]. Beveridge [800]. Bevölkerungstheorie [400]. Bews [742]. Bhatnagar [517]. Biology [137, 204, 535]. Birth [90, 260]. Birth-Rate [90, 260]. Births [865, 823, 847]. Bledisloe [250]. Board [849, 606, 270]. Boards [406]. Boddington [755, 199, 556]. Bodio [112]. Bombay [537, 537, 346]. Bonar [467].

Book [183, 405, 518, 536, 593, 637, 761, 825, 853, 870, 897, 242, 40, 62, 63, 64, 80, 125, 126, 161, 163, 180, 202, 222, 223, 225, 243, 245, 272, 312, 314, 365, 366, 385, 427, 428, 453, 468, 470, 490, 516, 589, 616, 636, 661, 668, 679, 741, 780, 798, 827, 828, 830, 845, 867, 873, 875, 895, 380, 8, 60, 82, 89, 103, 138, 162, 164, 203, 219, 224, 268, 274, 313, 363, 364, 404, 406, 469, 511, 537, 540, 559, 562, 565, 566, 581, 583, 592, 611, 612, 613, 630, 632, 642, 643, 663, 667, 688, 693, 707, 709, 716, 736, 739, 773, 778, 781, 802, 804, 806, 821, 829, 868, 869, 894, 184, 204, 421, 441, 54, 179, 264, 286, 359, 362, 443]. Book [444, 483, 820, 831, 41, 109, 128, 188, 360, 402, 422, 512, 607, 849, 83, 106, 107, 143, 146, 166, 206, 226, 227, 239, 250, 273, 295, 310, 318, 337, 379, 381, 399, 400, 706, 740, 800, 807, 818, 844, 896, 9, 11, 35, 37, 55, 86, 121, 160, 220, 221, 240, 241, 248, 267, 271, 287, 288, 289, 291, 293, 340, 342, 344, 345, 346, 386, 387, 407, 408, 424, 450, 472, 473, 489, 510, 560, 561, 564, 580, 587, 606, 609, 640, 664, 665, 666, 682, 692, 711, 715, 731, 732, 734, 735, 738, 758, 774, 782, 797, 822, 824, 826, 848, 871, 892, 7, 30, 102, 119, 217, 218, 265, 336, 420, 531, 605, 819, 59, 127, 776, 852, 53, 777, 14, 16]. Book [31, 32, 84, 88, 139, 144, 145, 270, 319, 339, 361, 388, 445, 539, 608, 610, 662, 683, 684, 689, 691, 742, 743, 757, 760, 762, 775, 779, 801, 846, 850, 851, 275, 320, 321, 369, 442, 486, 506, 535, 557, 590, 658, 659, 823, 10, 17, 34, 38, 39, 57, 58, 81, 247, 292, 451, 513, 554, 555, 563, 591, 681, 759, 783, 898, 33, 205, 294, 317, 382, 383, 384, 452, 464, 487, 491, 507, 508, 509, 515, 532, 533, 538, 578, 579, 584, 585, 586, 588, 629, 644, 660, 718, 753, 755, 756, 796, 803, 36, 61, 85, 104, 105, 122, 123, 124, 140, 142, 165, 181, 182, 186, 198, 199, 200, 244, 246, 338, 341, 343, 401, 403, 423, 449, 454, 467]. Book [471, 505, 534, 541, 582, 631, 634, 635, 639, 641, 687, 710, 713, 714, 730, 754, 893, 249, 87, 108, 141, 185, 201, 266, 290, 367, 389, 425, 426, 429, 517, 556, 799, 805, 315, 316, 368, 638, 690, 708, 712, 717, 872, 874, 876, 12, 13, 15, 56, 187, 269, 615, 633, 733, 737, 847, 90, 120, 158, 159, 238, 311, 484, 446, 447, 448, 465, 466, 485, 488, 514, 558, 614, 680, 685, 686]. Bordes [513]. Bouniatian [245]. Bowie [293]. Bowker [806]. Bowley [564, 662, 54, 422, 658, 182, 465]. Bravais [336]. Bravais-Pearsonschen [336]. Bresciani [869]. Bretagne [615]. Brigham [321]. Brij [64, 468]. Britain [668, 709, 806, 106, 318, 140, 635, 56, 479, 175, 405, 825, 876, 680]. British [202, 780, 406, 540, 581, 291, 731, 757, 779, 851, 140, 799, 315, 13, 574, 173, 548, 827, 800, 240, 241, 609, 39, 140]. Brockway [294]. Brown [667]. Brownlee [671]. Bruins [358]. Brunner [853]. Brussels [138, 138]. Budgets [346, 609]. Buer [715]. Building [518]. Bunting [164]. Burden [500]. Burdett [68]. Bureau [338, 343]. Burgess [892, 801]. Burnett [537]. Burnett-Hurst [537]. Burns [180, 612]. Burt [320, 644]. Business [366, 291, 386, 471, 893, 56, 488, 558, 487, 199, 343, 713, 754, 638, 686]. Businesses [577]. Buying [688, 829].

C

[518, 825, 63, 125, 180, 661, 875, 89, 540, 613, 773, 781, 804, 806, 869, 362, 483, 41, 206, 844, 715, 777, 144, 684, 691, 321, 81, 554, 759, 464, 533, 796, 123, 181, 244, 338, 754, 893, 690]. Calcolo [30]. Calculation [332, 176]. Calculus [286, 379]. Can [143]. Canada [117, 580, 512]. Canadian [2]. Cannan [735]. Capacity [219, 562, 500]. Capital [258, 779, 85, 123, 200, 368, 164]. Capitalism [61, 690]. Capitalist [536]. Caradog [217, 660, 756]. Carl [732, 321]. Carr [756, 238]. Carr-Saunders [756, 238]. Cartels [757]. Carver [56, 56]. Case [61]. Cassel [482, 287, 851]. Castelnuovo [30]. Casualties [234]. Causal [506]. Causation [605]. Caused [408]. Causes [733]. Cecil [487]. Census [6, 155, 417, 333, 331, 395, 461, 397]. Centenary [388]. Central [777]. Century [159, 842, 48, 81]. Certain [728]. Change [242, 54, 288]. Changes [3, 50, 116, 752, 260]. Chapin [225]. Chapters [105]. Characteristics [361]. Charbon [345]. Charles [224, 344, 17, 300]. Charters [777]. Chief [90]. Children [369]. Chile [342]. Chiozza [126]. Christian [607]. Christopher [853]. Circles [166]. Circulation [850]. Clair [37]. Clapham [187, 680]. Clarke [295, 299, 249]. Class [346, 790]. Classes [537, 48]. Clay [872]. Clearing [141]. Cleona [760]. Climatic [446]. Clothing [876, 272]. Club [388, 250]. Co [63, 223, 89, 224, 585, 63, 89, 613, 14, 509]. Co-Operation [63, 89, 63, 89, 613, 14]. Co-Operative [223, 224, 585, 509]. Coal [798, 799, 315, 316, 490, 830, 271, 586]. Coal-Mining [798]. Coefficient [866]. Coefficients [235, 308]. Coghlan [619]. Cohen [202]. Cohn [267]. Cole [741]. Collapse [625]. Coller [584]. Collis [184]. Combinations [714]. Commerce [13, 615, 272, 199, 312, 755, 556]. Commercial [142]. Commerciali [266]. Commission [90]. Commissioners [444]. Committee [5, 428, 103, 203, 656, 824, 508, 358, 294]. Commodities [74, 118, 196, 377, 463, 550, 654, 841, 551, 284, 772]. Commons [274]. Communities [307, 104]. Comparative [468]. Comparing [887, 307]. Comparison [414, 396, 135, 575]. Comparisons [305]. Compilation [776, 820]. Complementi [486]. Compliation [199]. Components [677]. Composition [587]. Computation [285]. Comstock [895]. con [443]. Concerning [601]. Conciliation [241]. Conclusions [483]. Conditions [537, 398, 107, 39, 13, 202]. Conference [138]. Confessions [536]. Conflicts [830].

Conjoncture [245]. Connues [615].

Consequences [243, 867, 32, 338, 488, 140]. Considerazioni [869]. Constitution [777, 616]. Constitutions [752]. Construction [361]. Consumer [223, 688, 802, 224]. Contingencies [286]. Contingency [176, 177]. Continuous [428]. Contracts [165]. Contribution [758, 605]. Contributions [602, 329]. Control [436, 106, 270, 491, 874, 15, 56, 365, 800, 871, 127]. Controlled [52]. Controversy [159, 692]. Coolidge [485]. Coombs [828]. Correa [332, 159]. Correction [308]. Correlated [628, 439]. Correlation [551, 440, 552, 803, 793, 866, 308, 152, 336, 730]. Correlations [888, 525]. Corrigenda [769]. Corrigendum [882, 906]. Corso [337]. Cost [183, 163, 490, 50, 643, 797, 790]. Costing [16]. Costs [135, 206, 272, 386]. Cotton [314, 750, 109, 188, 270, 614]. Council [490, 759, 710, 78, 156, 236, 334, 418, 503, 603, 704, 794, 890]. Countries [614, 797, 53]. Cours [845]. Course [379, 217, 629]. Covariation [446]. Cox [82]. Crawford [21]. Credit [312, 802, 37, 55, 288, 201, 8, 868, 387, 125]. Credit-Power [125]. Criminal [732]. Crises [245]. Crisis [561, 14, 426]. Criticism [707]. Crop [49, 727]. Crop-Cycles [49]. Crops [887]. Crossroads [384]. Crowell [165]. Crown [513]. Crowther [637]. Crum [630, 742, 737]. Currencies [340, 86]. Currency [625, 138, 480, 339, 201, 514, 8, 563]. Current [19, 43, 66, 92, 111, 130, 148, 168, 190, 208, 229, 252, 277, 297, 323, 348, 371, 391, 410, 431, 456, 475, 493, 520, 543, 568, 595, 618, 646, 670, 695, 720, 745, 764, 785, 809, 833, 855, 878, 900, 219, 194, 469, 403]. Currie [630]. Curve [530]. Curves [659, 701, 730]. Cycle [246]. Cycles [49, 343, 754, 829, 386, 893, 447, 488, 558, 367]. Cyclical [775, 448, 737, 661]. Cyril [775, 320, 644]. Czuber [359, 400].

D [897, 611, 188, 217, 270, 247, 660, 756, 142, 633]. Dahlberg [823]. Dalton [290, 104]. Daniels [109, 660]. Danish [548]. d’Apres [805]. d’Apres-Guerre [805]. Darmois [818]. Darwin [665]. Data [212, 659]. David [274, 362, 342]. Davison [875]. Dawes [896, 897]. Dawson [589]. Day [806, 635, 294, 637]. Days [715]. Dearle [643]. Death [330, 28, 307, 204]. Death-Rate [28]. Death-Rates [330, 307]. Deaths [847]. Debates [13]. Debt [656, 444, 99, 451, 85, 500]. Debts [451, 425]. Deceit [822]. December [78, 156, 236, 334, 418, 503, 603, 704, 890]. Declaration [13]. Decomposition [677]. d’Économie [631, 703]. Deflation [627, 122]. del [470]. Delinquent [644]. Delisle [180]. Della [102, 453, 87]. Demand [821, 247, 538]. Democracy [125]. Demografia [239]. Demografica [266]. Denmark [63, 354, 887]. Dennis [692]. Density [330, 28]. Department [222, 226, 368]. Depreciation [864, 513]. Depression [750, 407, 742]. Depuis [242, 831]. Derived [283]. Description [89]. Determination [758]. dettate [266]. Deutsche [663]. Development [540, 771, 53, 591, 783, 175, 187]. Developments [291, 139]. Deviates [888]. Deviations [628]. Devine [40, 586]. Difference [152]. Differences [771, 803]. Dilema [799]. Diminished [564]. Diminishing [681]. Disability [361]. Discharged [40]. Discount [510]. Discourse [640]. Discrepancey [398]. Discussion [529, 882, 282, 839, 233]. Disease [282, 27, 816]. Distribution [3, 272, 216, 728, 375, 54, 235, 888, 793, 338, 641, 528, 866, 439]. Distributions [27, 4]. District [577]. Districts [101]. do [525]. Dobb [783]. Dobbs [876]. Doctrines [385]. Documents [138, 851, 615]. Does [141].

Dopo [470, 590]. Dorothy [406]. Double [358, 481, 889]. Douglas [125, 566]. Dr [483, 896, 267]. Dr. [253]. Draft [125]. Drawings [793]. Droite [845]. Dublin [580]. Dubrowski [898]. Dudfield [495]. Due [573, 440]. Dumas [408]. Dundas [404]. d’une [227]. Dunn [711]. Durante [590]. During [3, 47, 864, 162, 106, 72, 240, 75, 51, 26, 39, 315, 448]. Duty [140]. Dynamic [198].

Early [109, 715, 691]. Earning [537]. Earnings [407]. Earth [867]. East [384]. Economic [22, 44, 69, 93, 113, 131, 149, 169, 191, 209, 230, 254, 278, 302, 325, 349, 372, 392, 411, 432, 458, 476, 496, 521, 545, 570, 597, 620, 649, 672, 696, 722, 747, 766, 787, 810, 834, 858, 880, 902, 64, 243, 138, 537, 540, 630, 781, 806, 607, 344, 782, 601, 32, 762, 846, 292, 451, 591, 783, 338, 343, 403, 541, 634, 108, 13, 187, 446, 447, 680, 82, 363, 362, 57, 33]. Economica [266, 453]. Economici [869]. Economics [825, 636, 473, 144, 319, 608, 487, 660, 635, 639, 717, 490, 759, 710, 422, 740, 489, 145, 292, 383, 198, 638, 802, 41, 429, 558]. Economique [245]. Economiques [245]. Economist [735]. Economy [388, 181, 505, 517]. Edgar [184]. Edgeworth [569, 505]. Edinburgh [178]. Edmund [901]. Edward [40, 16, 384, 586, 635, 687]. Edwin [581, 802, 735, 122]. Effects [73, 201, 55, 344]. Eight [157]. Eight-Seventh [157]. Eighteenth [842]. Eighteenth-Century [842]. Eighth [237, 236]. Eighty [79, 78, 156, 237, 236, 335, 334]. Eighty-Eighth [237, 236]. Eighty-Ninth [335, 334]. Eighty-Seventh [156]. Eighty-Sixth [79, 78]. Einaudi [358]. Elborne [16]. Elbourne [687]. Elderton [445, 775, 534, 730]. Element [501].

Elementary [212, 309]. Elementi [7]. Elementos [443]. Elements [421, 489, 658, 533, 660]. Elkin [777]. Elliptic [466]. Ellis [57, 731]. Emanuel [359]. Emil [739]. Emmott [647]. Empire [13, 540]. Employees [293]. Employment [780, 528, 863, 407]. end [108]. Ended [236, 334, 418, 503, 603, 704, 890, 78, 156]. Ending [78, 156, 236, 334, 418, 503, 603, 704, 794, 890]. Enfield [426]. Engberg [759]. Engineering [428, 827, 135, 796]. England [560, 666, 509, 756, 269, 573, 887, 414, 462, 48, 727, 105, 714, 15, 146, 515]. English [109, 610, 200, 396, 294, 205, 294, 718, 105]. Enquiries [791]. Enquiry [314, 415, 294]. Enstrom [446]. Enterprise [608]. Entwicklung [739]. Erdman [273]. Erich [663]. Ernest [299]. Ernle [205, 718]. Ernst [336]. Error [440, 502]. Errors [628]. Especial [152]. Especially [159]. Esquisse [227]. Essai [245]. Essay [414]. Essays [362, 64]. Estadistica [443]. Estimates [727]. Estimation [159]. Ethel [534]. Ethic [607]. Eugenic [665]. Eugenics [534]. Europe [227, 213, 771, 847]. European [752]. Evasion [481, 889]. Eve [515]. Event [676]. Everett [799]. Eversley [786]. Evidence [90]. Evolution [581, 483, 807, 238]. Examination [653]. Examinations [261]. Excess [140]. Exchange [162, 136, 287, 339, 482, 780, 221]. Exchanges [423, 201]. Expansions [4]. Expenditure [761, 761]. Expenses [365]. Experience [444, 15]. Experiment [415]. Experimental [177]. Experiments [355, 491]. Experts [138]. External [285].

F [80, 225, 873, 736, 804, 286, 337, 706, 807, 248, 289, 587, 609, 774, 127, 689, 691, 775, 681, 200, 246, 505, 316, 717, 269, 446]. F.R.S [483]. Factor [607, 605]. Factors [436, 846, 506]. Factory [592]. Faculté [845]. Fall [248]. False [292]. Families [790].

Family [848, 566, 565, 583]. Farmer [759]. Farming [205, 718, 548, 206]. Farr [28]. Fatigue [606, 329]. Fay [89, 613, 806, 81]. Feb [765]. Federal [510, 195]. Federated [428]. Fedselen [557]. Felice [399]. Fenelon [473]. Fenner [294]. Fernand [682]. Fertility [576, 48, 260]. Field [225]. Fifteenth [353]. Fifth [891, 890]. Fifty [2]. Figures [842]. Filadelfo [119, 486]. Final [183]. Finance [761, 616, 138, 469, 581, 35, 240, 689, 105, 140, 201, 425, 290, 364, 773, 450, 34, 591, 124, 290, 12]. Finances [424]. Financial [78, 156, 236, 334, 418, 503, 603, 704, 794, 890, 138, 561, 194, 36, 140, 269, 358, 161, 560, 449]. Finanziaria [337]. Finaziaria [218]. Findlay [314, 35, 346, 450, 633]. First [504, 503, 664, 217, 629]. Fiscal [269]. Fisher [668, 693, 420, 531, 819, 311]. Fishery [462]. Fisk [105, 425]. Fit [355, 235, 843]. Fitzgerald [714]. Flier [424]. Flores [443]. Fluctuations [661, 136, 775, 285, 737, 448, 684]. Flux [363, 423]. Fondamenti [453]. Food [83, 800, 26, 52, 83, 491, 84, 317]. Forces [198]. Ford [637]. Forecast [454]. Forecasting [713, 306, 600, 471]. Foreign [136, 221, 287, 711, 797, 689, 451, 423, 376, 201, 162]. Formula [659]. Formulae [235, 466]. Forrester [188, 585]. Forschungsmethoden [359]. Forsyth [464]. Foster [471]. Foundations [383]. Fourth [795, 794]. Franc [682]. France [761, 55, 187, 615, 49, 761, 188, 561]. Frances [20]. Francis [582, 484, 197, 879, 569]. Franco [239]. Frank [60, 898, 584, 108]. Frankel [852]. Franklin [165]. Frederic [781]. Frederick [402, 487, 533, 244, 338]. Free [80]. French [344]. Frequency [216, 416, 27, 4, 888, 262, 439, 265, 730]. Frisch [631]. Front [624, 726, 813, 838, 862, 884]. Fry [796]. Function [441, 793]. Functions [616, 466].

Fundamental [139]. Fundamentals [532]. Fundamentos [443]. Further [791]. Future [867, 14, 339].

G [593, 62, 161, 243, 314, 616, 668, 8, 268, 313, 562, 632, 642, 868, 184, 441, 483, 879, 379, 818, 35, 160, 287, 346, 450, 473, 451, 629, 660, 341, 634, 517, 685]. G.C.S.I. [746]. Galton [197, 484]. Gaston [761]. Gavett [629]. Geldes [739]. General [79, 78, 157, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890, 866]. Generali [486]. Generalized [502]. Generating [447]. Genius [731]. George [156, 867, 109, 318, 561, 782, 743, 746, 140, 708]. German [625]. Germany [187, 864, 26, 589, 896]. Gesetz [264]. get [525]. Gibson [240]. Gide [224, 344]. Gideonse [897]. Gilbert [690]. Giorgio [7, 590, 266]. Glover [67]. Gold [664, 770, 73, 511, 267]. Goldenweiser [510]. Gompertz [378]. Gonnard [385]. Gonner [362]. Goodness [355, 235, 843]. Goodwill [142]. Gordon [63, 871]. Governing [438]. Government [593, 263, 444, 249, 56, 469, 294, 180, 165]. Graduating [659]. Graduation [361]. Granet [825]. Graphical [817]. Graunt [729]. Gray [15]. Great [405, 825, 156, 668, 709, 106, 88, 140, 439, 876, 56, 806, 479, 75, 51, 175]. Greenwood [184]. Gregory [851, 162, 268, 511, 664, 850]. Griffith [632]. Grossen [264]. Groundwork [422]. Groves [848]. Growing [614]. Growth [436, 577, 576, 573, 633]. Guerra [239, 470, 590, 87]. Guerre [242, 683, 805]. Guidance [606]. Guido [30]. Guilds [62, 14]. Guillermo [342]. Guldberg [336]. Gunnar [823]. Guy [825]. Guyot [765].

H [125, 161, 365, 366, 616, 164, 274, 566, 583, 592, 611, 716, 820, 188, 800, 240, 472, 510, 564, 640, 692, 265, 777, 339, 445, 608, 247, 783, 317, 464, 491, 579, 584, 893, 389, 187, 447, 680]. Habits [688]. Haig [140]. Halbwachs [831]. Hall [237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890, 448]. Hamburger [894]. Hamilton [779, 746, 874]. Hammer [806]. Hammond [38, 38, 39]. Handbook [579]. Handley [867]. Happenings [27]. Harald [408]. Harold [82, 451]. Harris [593]. Harrison [741, 709]. Harry [897, 442]. Hartley [387, 36, 61]. Hartshorne [822]. Harvey [105, 425]. Hastings [386]. Havelock [731]. Hawtrey [8, 313, 868, 634]. Hayem [615]. Health [184, 421, 675, 824, 715, 734]. Heaton [128]. Heinrich [389]. Held [157, 237, 236, 335, 419, 418, 604, 603, 705, 704, 891, 79, 78, 156, 334, 504, 503, 795, 794, 890]. Helen [274, 799]. Henderson [247, 270]. Henri [761, 227, 615]. Henry [637, 68, 856, 821, 250, 273, 740, 826, 84, 513, 753, 449, 872, 12, 286]. her [761, 451]. Herbert [128, 852]. Hereditary [823]. Hermann [757]. Heterotypic [216]. Hettinger [471]. Hexter [488]. Higgs [449, 12]. High [582, 122]. Hirst [609]. Histoire [227, 615, 385]. Historical [655, 13, 59]. History [707, 53, 691, 850, 58, 292, 142, 108, 269, 680, 274]. Hjalmar [738]. Hoagland [274]. Hobhouse [294]. Hobson [62, 319]. Hodges [60]. Hogg [564]. Holland [199, 141]. Holland-Martin [141]. Holmes [140]. Home [89, 613]. Homer [742]. Hon. [619]. Honorary [281]. Horace [578, 120]. Hour [428]. Hours [314, 798, 407]. House [141]. Household [429]. Housing [330, 537]. Howard [15]. Hubert [849, 270, 247, 541]. Hudson [386]. Hugh [583, 820, 822, 104, 290]. Human [605, 535, 238]. Humbert [427]. Humphreys [281, 298]. Hunter [717]. Hurry [166]. Hurst [537]. Hypergeometric [553]. Hypothesis [398].

I. [13]. I.S.O [281]. III [138]. Illustrated [2, 756]. Illustrations [177]. Impact [782]. Imperial [25]. Incidence [375, 101]. Including [295]. Income [3, 882, 709, 728, 54, 839, 814, 662, 338, 244, 338, 641]. Incomes [104]. Inconnues [615]. Increase [867, 103, 580]. Indebtedness [512]. Index [551, 332, 501, 357, 75, 159, 311]. Index-Numbers [551, 332, 501, 357, 75, 159, 311]. India [468, 203, 562, 155, 417, 243, 562, 581, 642, 339]. Indian [64, 616, 203, 339, 517, 35]. Indices [652]. Individual [628]. Inductive [512]. Industrial [906, 516, 589, 184, 652, 143, 807, 606, 715, 840, 396, 676, 292, 515, 175, 863, 872, 397, 329, 826, 127, 684, 714]. Industrie [615]. Industriekonjunktur [736]. Industries [828, 775, 128]. Industry [428, 750, 188, 291, 758, 742, 591, 528, 315, 712, 13, 15, 653, 125, 180, 314, 428, 219, 41, 109, 14, 539, 757, 639, 316, 798, 31, 34, 759]. Inequality [104]. Inevitable [454]. Infant [865, 506, 733]. Inflation [627, 9]. Influence [551, 88]. Influenza [121]. Information [1, 71, 96, 151, 172, 232, 257, 328, 352, 413, 435, 498, 524, 599, 623, 698, 725, 789, 883, 657]. Ingalls [338, 403]. Inhabitants [867]. Inheritance [549]. Inleiding [844]. Inquiry [27, 346]. Insolera [337, 119, 486]. Instalment [802]. Institute [490, 606, 759, 710, 353, 527]. Institution [156]. Instituto [266]. Insurance [283, 675, 691, 202, 361]. Insured [361]. Intelligence [732, 321]. Inter [425]. Inter-Ally [425]. Interest [870]. Intermediate [499]. International [138, 707, 512, 305, 414, 271, 760, 452, 353, 527, 774, 382]. Interpolation [379, 706]. Interpretation [176, 816, 234]. Interrelation [375]. Introduccion [443]. Introduction [630, 850, 464, 578, 485, 892, 145]. Introductory [363]. Invalidity [601]. Inversion [402]. Investigation [906, 212, 840, 198, 441]. Investments [711].

Iron [47, 742, 516]. Irving [693, 629, 311]. Isador [490, 799]. Isidoro [443]. Issue [777]. Italia [590]. Italy [602, 75, 51]. IV [138].

J [536, 245, 873, 164, 274, 364, 404, 559, 562, 716, 483, 143, 295, 706, 9, 271, 291, 424, 580, 587, 609, 892, 14, 88, 319, 445, 608, 10, 38, 57, 513, 383, 140, 165, 244, 471, 249, 389, 316, 690, 187, 269, 485, 680]. Jacob [512]. Jacques [253]. Jagtiani [472]. James [63, 707, 293, 648, 467, 775]. Jamieson [166]. Japan [712]. Jean [805]. Jean-Pierre [805]. Jenks [779]. Jerome [893, 442]. Jevon [339]. Jevons [339]. Jeze [761]. John [67, 780, 830, 274, 667, 179, 295, 32, 145, 659, 401, 236, 334, 418, 503, 603, 704, 794, 890, 671]. Jones [660, 217, 608, 756, 186]. Joseph [202, 610, 721]. Joshi [562, 642]. Josiah [219, 469, 358, 607, 139, 662, 762, 846, 687, 708]. Julien [615]. Julin [158, 360]. June [157, 79, 78, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890].

K.C.M.G. [901, 619]. Karin [870]. Karl [534, 484]. Katzenellenbaum [514]. Kaufkraft [739]. Kelley [380]. Kemmerer [122]. Kent [533]. Keynes [179, 32, 401]. Khambata [562]. Kiernan [140, 364]. King [244, 338, 407, 198]. Kingdom [89, 182, 728, 72, 627, 49, 249]. Kinley [342]. Kirkaldy [240, 241, 34]. Kirkus [776]. Kisch [777]. Knauth [338, 244]. Knibbs [867]. Knowles [540]. Kock [870]. koeffizienten [336]. Korrelations [336]. Korrelations-koeffizienten [336]. Kuczynski [847]. Kuznets [661].

L [63, 380, 274, 540, 184, 54, 422, 226, 221, 564, 871, 662, 658, 38, 563, 515, 532, 579, 182, 15, 447, 465, 485]. Labour [222, 314, 828, 274, 829, 258, 241, 480, 886, 39, 81, 368, 636, 667, 743, 427, 537, 107, 666, 292, 124]. Labourer [38]. Labours [484]. Lamb [588].

Lancashire [750, 806]. Land [887, 404]. Land-Value [404]. Lands [593]. Large [585]. Large-Scale [585]. Lassar [267]. Lassar-Cohn [267]. Laughlin [10]. Laurence [10]. Lavington [200, 246]. Law [28, 146, 502, 378, 76]. Lawrence [248, 248]. Laws [821, 541, 438, 295]. Lawson [127]. Layton [582, 713]. Lazard [805]. League [358]. Leake [142]. Leather [639]. l’Economique [220]. Lecture [607]. Lectures [219, 872]. Legislation [592, 39]. Legitimate [260]. Lehfeldt [340]. Leland [779]. Lemus [443]. Leo [126, 451]. Leonard [630, 665, 14, 742, 633, 737]. Lester [681, 755, 199, 556]. Letter [77]. Letters [484]. Leurence [682]. Level [611]. Levels [885]. Lever [641]. Levy [123, 757]. Lewis [753, 760]. Lezioni [119, 218, 266]. Liberty [82]. Library [95, 23, 45, 70, 94, 171, 114, 132, 150, 170, 256, 192, 210, 231, 255, 351, 279, 303, 326, 350, 434, 373, 393, 412, 433, 523, 459, 477, 497, 522, 622, 546, 571, 598, 621, 724, 650, 673, 697, 723, 812, 748, 767, 788, 811, 904, 835, 859, 881, 903, 224]. Life [137, 444, 233, 408, 886, 361, 76, 234, 484, 344, 762, 846, 286, 81]. Life-Table [137]. Life-Tables [233, 76]. Light [139]. Limits [863]. Lincoln [448]. Liquor [56]. List [95, 23, 45, 70, 94, 171, 114, 132, 150, 170, 256, 192, 210, 231, 255, 351, 279, 303, 326, 350, 434, 373, 393, 412, 433, 523, 459, 477, 497, 522, 622, 546, 571, 598, 621, 724, 650, 673, 697, 723, 812, 748, 767, 788, 811, 904, 835, 859, 881, 903]. Litman [106]. Live [72]. Livelihood [564]. Lives [361]. Living [50, 797, 790, 643, 183, 163]. Llewellyn [849, 541]. Lloyd [582, 491]. Load [758]. Local [294, 249, 593]. Loch [300]. Lodewijk [894]. Logic [677]. Logistic [701]. London [78, 289, 333, 212, 309, 886, 101, 260]. Long [686]. Lord [250, 205, 718, 647, 746, 786]. Loria [453, 470, 87]. Loss [234]. Losses [408]. Louis [580]. Lubin [490, 799]. Lucas [63]. Lucien [683, 803, 158]. Luigi [218, 112]. Lynden [292].

M [245, 612, 143, 55, 345, 424, 472, 489, 715, 605, 59, 127, 506, 39, 384, 491, 756, 471, 534, 158, 238]. M. [77]. Macassey [292]. Macaulay [244, 338]. Macgregor [188]. Machinery [743]. Majesty [824]. Major [184]. Makeham [76]. Making [311, 710]. Mally [336]. Malthus [632, 467]. Mankind [580, 384]. Manufacture [741]. Manufacturing [828]. Many [593, 53]. March [803, 683, 803, 158]. Marcus [301]. Margaret [564, 591]. Maritime [615]. Mark [864, 822, 738]. Market [642, 688, 52, 685, 289, 801, 200]. Marketing [273, 509, 585, 687]. Markets [341]. Marriage [576, 848]. Marshall [312, 457, 679, 31, 554]. Martin [141]. Matematica [443, 337, 218, 443]. Matematiche [486]. Maternal [733]. Mathematical [483, 422, 336, 420, 464, 579, 466, 485, 753]. Mathematics [892, 660, 379]. Mathematique [818]. Mathematische [400]. Mathieson [146]. Matter [624, 726, 837, 861, 905, 813, 838, 862, 884]. Matthews [141]. Maturing [99]. Mauldon [717]. Maurice [831, 783, 641, 488]. Mawas [242]. Max [896]. May [822, 874]. Maynard [179, 32, 401]. Mean [888, 285]. Means [628, 439]. Measurement [50, 116, 885, 465]. Measures [398]. Meat [72, 318]. Meconnues [615]. Meerwarth [555]. Meeting [174, 259, 437, 526, 626, 815, 79, 78, 157, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890]. Memoires [615]. Memoranda [138]. Memorials [554]. Mental [601, 320, 369]. Merchandising [272]. Mess [592, 826]. Metals [117]. Method [441, 530, 887, 152, 177, 380, 268, 442, 629]. Methode [803]. Methods [630, 416, 602, 531, 819, 396, 487, 307, 578, 120]. Metodo [381]. Metodologia [443]. Metodologica [399, 119]. Mexican [898]. Middle [790, 48]. Middle-Class [790]. Middleton [317]. Migration [667, 666, 779, 285, 893]. Milk [507, 273, 587, 507]. Mill [314]. Miller [591]. Mills [804, 487]. Milnes [636]. Miner [659, 490]. Mines [60]. Minetola [262]. Minimum [153, 716]. Mining [125, 798, 315]. Ministers [178]. Ministry [824, 491]. Mint [117]. Minutes [103]. Misstatement [331]. Misura [102]. Mitchell [494, 244, 338, 661, 754, 893]. Mittelman [274]. Modern [895, 602, 139, 846, 33, 104, 159, 680, 778]. Modernen [264]. Moffit [515]. Molecular [100, 215]. Moments [553]. Monetaire [87, 242]. Monetaires [805]. Monetarie [470]. Monetary [778, 401, 313, 559, 342]. Money [642, 55, 289, 702, 801, 341, 685, 247, 37, 312, 642, 287, 339, 10]. Monnaie [288, 160]. Monopolies [757]. Monopoly [871]. Montagu [593]. Month [499]. Moore [447]. Moral [601]. Moralities [307]. Morbidity [175]. Morgan [291]. Morphologie [245]. Morrow [637, 186]. Morse [733]. Mortalidad [443]. Mortalities [440]. Mortality [573, 865, 361, 445, 101, 378, 733, 506, 444]. Mortara [7, 590, 266]. Moscow [52]. Motor [853, 573]. Moulton [161, 451]. Mouvement [683]. Movement [864, 610, 223]. Movements [146]. Moylan [159]. Mr. [529, 332, 530, 357]. Multiple [27, 262, 866]. Municipalities [616]. Murchison [732]. Murphy [324]. Murray [140]. Mutton [588].

N [562, 642, 539, 563, 56]. Narain [64, 468]. Nation [83, 140]. National [3, 882, 364, 656, 54, 444, 839, 814, 606, 662, 85, 124, 338, 343, 12, 675, 90, 448, 62, 635]. Nationalisation [60]. Nationalization [126]. Nationalokonomie [555]. Nations [358]. Natural [580]. naturali [381]. Nature [27, 525, 465]. Need [665, 635]. Neison [879]. nel [266]. Netherlands [424, 397]. Newbold [605]. Newman [184]. Newmarch [219, 850]. Newsholme [421, 734]. Niceforo [381, 102]. Nicholson [9, 721]. Nineteenth [81, 48]. Ninetienth [418]. Ninetieth [419]. Ninety [504, 503, 604, 603, 705, 704, 795, 794, 891, 890]. Ninety-Fifth [891, 890]. Ninety-First [504, 503]. Ninety-Fourth [795, 794]. Ninety-Second [604, 603]. Ninety-Third [705, 704]. Ninth [335, 334]. No [138, 361]. Noel [633, 281, 298]. Nogaro [778]. Nonsense [525]. Nonsense-Correlations [525]. Nora [636]. Normal [888, 793]. Norman [643]. North [272]. Northern [847]. Nos [138]. Note [178, 216, 333, 655, 309, 843]. Notes [19, 43, 66, 92, 111, 130, 148, 168, 190, 208, 229, 252, 277, 297, 323, 348, 371, 391, 410, 431, 456, 475, 493, 520, 543, 568, 595, 618, 646, 670, 695, 720, 745, 764, 785, 809, 833, 855, 878, 900, 214]. Notz [739]. November [98]. Number [178, 888]. Numbers [551, 332, 501, 357, 72, 75, 159, 311]. Numerical [379]. Nutrition [633].

O [408]. Oakley [445]. Obituary [746]. O’Brien [63]. Observation [398]. Observational [659]. Observations [379, 575, 600, 439]. Occupational [601]. Occupations [906, 840]. Occurrence [27]. Occurring [448]. Office [491, 824]. Official [5, 771, 2, 479, 679]. Ogburn [848]. Old [875]. Oldham [577]. Olivier [345]. Om [557]. one [178]. Operation [63, 89, 510, 63, 89, 613, 14]. Operative [223, 224, 585, 509]. Optimum [758]. Organisation [25]. Organization [97, 161, 807, 291]. Organized [734, 11]. Origin [691]. Orwin [206]. Oswald [37, 244, 338]. Other [18, 42, 65, 91, 110, 129, 147, 167, 189, 207, 228, 251, 276, 296, 322, 347, 370, 390, 409, 430, 455, 474, 492, 519, 542, 567, 594, 617, 645, 669, 694, 719, 744, 763, 784, 808, 832, 854, 877, 899, 361, 341, 872, 685]. Our [143]. Outflanking [164]. Outlook [140]. Overseas [540]. Ownership [59].

P [536, 562, 642, 879, 11, 871, 445, 142, 198, 876]. Page [710, 13, 13]. Paired [793]. Paish [708]. Palan [445]. Palin [775, 730]. Panandikar [243]. Pandemic [121]. Paper [29, 529]. Papers [679, 505]. Papier [160]. Papiergeld [267]. Parenthood [90]. Paris [831, 13, 101]. Parliamentary [13]. Partial [888]. Particular [27]. Particularly [676]. Past [205, 718]. Pasvolsky [451]. Path [480]. Paton [633]. Patrick [714]. Patton [630, 681]. Paul [566]. Peace [84, 13, 32]. Peake [341, 685]. Pearl [83, 204, 83, 535, 659]. Pearson [534, 484]. Pearsonschen [336]. Peddie [690]. Peel [561, 782]. Pendant [683]. Pensions [40]. Penty [14]. People [178, 338]. Percy [198]. Periodical [24, 46, 115, 133, 193, 211, 280, 304, 327, 374, 394, 460, 478, 547, 572, 651, 674, 769, 749, 768, 836, 860]. Periodicals [22, 44, 69, 93, 113, 131, 149, 169, 191, 209, 230, 254, 278, 302, 325, 349, 372, 392, 411, 432, 458, 476, 496, 521, 545, 570, 597, 620, 649, 672, 696, 722, 747, 766, 787, 810, 834, 858, 880, 902]. Periodicities [446]. Periodicity [816]. Periodiques [245]. Periodogram [356, 154]. Periods [428]. Peripezie [470, 87]. Perlman [274]. Permanent [361]. Persons [471]. Petersen [408]. Pethick [248]. Petroleum [538]. Petty [729]. Phenomena [33, 446, 465]. Philip [589, 275, 124, 141]. Photograph [281]. Physiology [37]. Physique [758]. Pierre [805]. Pigou [773, 144, 684, 554, 123, 181]. Plan [897, 896]. Point [823]. Policies [875]. Policy [611, 342, 510, 852, 292, 563, 404].

Polit [896]. Political [181, 505, 388]. Politique [345, 805]. Poor [295, 101]. Population [436, 468, 867, 330, 576, 831, 842, 715, 888, 438, 90, 238, 385, 632, 820, 580]. Populations [285, 752]. Position [817, 539]. Possibilities [867]. Post [829, 750, 700, 135, 425]. Post-War [829, 750, 700, 135, 425]. Potential [614]. Pound [29]. pour [615]. Poverty [564, 751, 564, 166, 633]. Power [164, 829, 125]. Practical [676, 713]. Practice [873, 771, 583, 389]. Practices [272]. Pre [700, 135]. Pre-War [700, 135]. Precious [117]. Precision [551, 628]. Preparation [199]. Prepared [222]. Present [867, 806, 539, 614, 205, 718]. Presentation [199]. Presented [156, 236, 334, 418, 503, 603, 704, 78, 794, 890]. Prevalence [816]. Price [263, 611, 375, 116, 106]. Prices [74, 118, 196, 377, 463, 550, 654, 841, 222, 213, 551, 707, 375, 284, 772, 248, 850, 75, 51, 122, 306, 600, 52, 850, 804, 894, 10, 106, 508, 182, 198]. Primer [740, 12]. Principal [361]. Principes [360, 158]. Principles [363, 139, 290]. Prison [294]. Prisons [294]. Private [608]. Prix [683]. Probabilades [443]. Probabilità [30]. Probabilities [420]. Probability [286, 501, 179, 485, 796]. Problem [853, 655, 871, 140, 471, 687, 754, 872, 159, 238, 152, 16, 634, 635]. Problème [703, 631]. Problèmes [220]. Problems [668, 219, 469, 632, 580, 734, 775, 194, 17, 563, 120, 64, 421, 36, 90]. Proceedings [79, 157, 98, 237, 174, 335, 259, 419, 437, 504, 526, 604, 626, 705, 795, 815, 891, 103]. Produce [509]. Production [354, 395, 72, 135, 317, 507, 397, 541]. Productive [652]. Productivity [887, 548]. Products [507]. professé [845]. Professions [615]. Professor [482, 552, 262]. Professors [358]. Profits [222, 293, 386, 140]. Progress [89, 2, 53, 173]. Prohibition [275, 693]. Properity [781]. Proportionate [440]. Proposal [85]. Prospects [539, 17]. Prosperity [407, 742, 759, 708]. Protest [735]. Provinces [309]. Provincial [581]. Provision [472]. Provisional [155]. Psychology [606]. Pubblica [590]. Public [138, 773, 421, 450, 212, 105, 500, 290, 425]. Publications [18, 42, 65, 91, 110, 129, 147, 167, 189, 207, 228, 251, 276, 296, 322, 347, 263, 370, 390, 409, 430, 455, 474, 492, 519, 542, 567, 594, 617, 645, 669, 694, 719, 744, 763, 784, 808, 832, 854, 877, 899]. Publicity [396]. Purchase [585]. Purdom [518]. Pure [703, 631]. Purpose [465]. Purposes [499]. Purposive [792]. Putnam [318].

Quelques [615]. Questions [159].

R [405, 825, 856, 830, 8, 89, 274, 313, 537, 559, 581, 583, 613, 802, 806, 868, 188, 250, 340, 640, 848, 892, 531, 819, 84, 506, 81, 585, 199, 244, 338, 634, 141, 266, 426, 315, 717, 847]. Raffalovich [77]. Ragnar [631]. Rail [825]. Railway [489, 479, 59, 852, 405, 776]. Railways [203, 472]. Rainfall [213]. Ramachandra [639]. Randolph [801]. Random [440]. Ratan [41]. Rate [867, 28, 90, 260]. Rates [330, 865, 309, 733, 307, 870]. Rau [639]. Rawlley [41]. Raymond [204, 83, 535, 659]. Raynes [830]. Re [707]. Re-Stateement [707]. Reading [678]. Readings [120]. Rechenfehler [736]. Reciprocity [80]. Reconstruction [560, 451, 85, 313]. Records [361]. Recovery [625, 451]. Recruited [309]. Redemption [85]. Redford [666]. Redmayne [315]. Reed [537]. Rees [291, 269]. Reference [202, 224, 802, 27, 285, 591, 152]. Reform [665, 401, 449]. Reformation [402]. Refrigerated [588]. Régime [227]. Reginald [495]. Register [368].

Registration [282, 771]. Regression [235]. Regulation [668, 427]. Relating [851, 505, 234, 653]. Relation [386, 361, 451, 140, 76]. Relations [258, 872]. Relationships [848]. Rene [385]. Renton [403]. Reparations [897, 221]. Repeated [27, 676]. Replacement [241]. Reply [482]. Report [183, 6, 78, 156, 236, 334, 418, 503, 603, 704, 794, 890, 222, 358, 656, 294, 90, 314, 103, 203, 121, 346, 824, 508, 585, 588]. Reports [851]. Representative [27, 777, 752]. Representing [416]. Republic [108]. Requisites [585]. Rer [896]. Research [222, 233, 606, 531, 819, 338, 368, 343, 225]. Reserve [510, 195, 801]. Restoration [340]. Results [415]. Retail [272, 661, 688]. Retailers [366]. Retarded [369]. Retrospect [175]. Return [511]. Returns [24, 46, 115, 133, 193, 211, 280, 304, 327, 374, 394, 460, 478, 547, 572, 651, 674, 769, 749, 768, 836, 860, 331, 681]. Review [183, 405, 518, 536, 593, 637, 761, 825, 853, 870, 897, 242, 40, 62, 63, 64, 80, 125, 126, 161, 163, 180, 202, 222, 223, 225, 243, 245, 272, 312, 314, 365, 366, 385, 427, 428, 453, 468, 470, 490, 516, 589, 616, 636, 661, 668, 679, 741, 780, 798, 827, 828, 830, 845, 867, 873, 875, 895, 380, 8, 60, 82, 89, 103, 138, 162, 164, 203, 219, 224, 268, 274, 313, 363, 364, 404, 406, 469, 511, 537, 540, 559, 562, 565, 566, 581, 583, 592, 611, 612, 613, 630, 632, 642, 643, 663, 667, 688, 693, 707, 709, 716, 736, 739, 773, 778, 781, 802, 804, 806, 821, 829, 868, 869, 894, 136, 184, 204, 421, 441, 54, 179, 264, 286, 359, 362]. Review [443, 444, 483, 820, 831, 41, 109, 128, 188, 360, 402, 422, 512, 607, 849, 83, 106, 107, 143, 146, 166, 206, 226, 227, 239, 250, 273, 295, 310, 318, 337, 379, 381, 399, 400, 706, 740, 800, 807, 818, 844, 896, 9, 11, 35, 37, 55, 86, 121, 160, 220, 221, 240, 241, 248, 267, 271, 287, 288, 289, 291, 293, 340, 342, 344, 345, 346, 386, 387, 407, 408, 424, 450, 472, 473, 489, 510, 560, 561, 564, 580, 587, 606, 609, 640, 664, 665, 666, 682, 692, 711, 715, 731, 732, 734, 735, 738, 758, 774, 782, 797, 822, 824, 826, 848, 871, 892, 7, 30, 102, 119, 217, 218, 265, 336, 420, 531, 605, 819, 59, 127, 776, 852, 53, 777, 14]. Review [16, 31, 32, 84, 88, 139, 144, 145, 270, 319, 339, 361, 388, 445, 539, 608, 610, 662, 683, 684, 689, 691, 742, 743, 757, 760, 762, 775, 779, 801, 846, 850, 851, 275, 320, 321, 369, 442, 486, 506, 535, 557, 590, 658, 659, 823, 10, 17, 34, 38, 39, 57, 58, 81, 247, 292, 451, 513, 554, 555, 563, 591, 681, 759, 783, 898, 33, 205, 294, 317, 382, 383, 384, 452, 464, 487, 491, 507, 508, 509, 515, 532, 533, 538, 578, 579, 584, 585, 586, 588, 629, 644, 660, 718, 753, 755, 756, 796, 803, 36, 61, 85, 104, 105, 122, 123, 124, 140, 142, 165, 181, 182, 186, 198, 199, 200, 244, 246, 338, 341, 343, 401, 403, 423, 449, 454]. Review [467, 471, 505, 534, 541, 582, 631, 634, 635, 639, 641, 687, 710, 713, 714, 730, 754, 893, 249, 87, 108, 141, 185, 201, 266, 290, 367, 389, 425, 426, 429, 517, 556, 799, 805, 315, 316, 368, 638, 690, 708, 712, 717, 872, 874, 876, 12, 13, 15, 56, 187, 269, 615, 633, 733, 737, 847, 675, 90, 120, 158, 159, 238, 311, 484, 446, 447, 448, 465, 466, 485, 488, 514, 558, 614, 680, 685, 686]. Revival [589]. Revolution [898, 105, 516, 715, 783, 898, 515]. Rew [250, 856, 740, 84]. Rhodes [530]. Ricardo [362]. Rice [659]. Rich [101]. Richard [21]. Richardson [716]. Rickman [842]. Riding [528]. Rietz [579, 753]. Riforma [453]. Rignano [549]. Rise [248]. River [717]. Road [473, 708]. Robert [827, 711, 140, 186, 733, 847]. Robertson [611, 247]. Robinson [379, 379]. Rogowski [663]. Role [472]. Roma [266]. Rome [108]. Ronald [875]. Rooms [78]. Roscoe [145]. Rowe [873, 316]. Royal [79, 78, 157, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890]. Rubin [301]. Rubio [443]. Rudolf [555]. Russell [88, 759]. Russia [451, 591, 107]. Russian [451, 451, 783, 514]. Russischen [898].

S

[156, 62, 243, 516, 661, 827, 143, 206, 852, 14, 262, 591, 783, 898, 198, 315, 712, 876, 514]. sa [803]. Saggi [239]. Sailors [40]. Sakulare [739]. Salaires [683]. Sales [272]. Salute [590]. Sample [415, 791, 888, 790]. Samples [655]. Sampling [440, 866, 525]. Samuel [637, 408]. Saposs [274]. Sargent [271]. Satellite [518]. Sauerbeck [857]. Saunders [756, 238]. Savorgnan [239]. Scale [585, 678]. Schacht [738]. Scheme [125, 402, 175, 549]. Schmidt [736]. Scholastic [320, 369]. School [272, 212, 309]. Schott [310]. Schultz [821]. Science [402, 450, 185]. Scientific [441, 402]. Scientifici [453]. Scienze [381]. Scotland [178, 331, 887, 63]. Sea [234]. Seárov [33]. Seasonal [285, 737]. Second [604, 603, 842, 90]. Secret [582]. Secretary [203, 281]. Secrist [365, 366, 578, 120]. See [615, 227]. Seksualproporsjonen [557]. Select [103, 851]. Selection [792]. Selig [274]. Seligman [581, 802, 358]. Selling [802, 365]. Sells [406]. Separate [677]. Sept [765]. Sequel [564]. Series [677, 842, 553, 525]. Sering [896]. Service [607]. servir [615]. Session [353, 527]. Sessional [78, 156, 236, 334, 418, 503, 603, 704, 890]. Sets [575]. Setting [754]. Seventeenth [159]. Seventeenth-Century [159]. Seventh [157, 156]. Sex [752]. Seymour [780]. Shadow [867]. Shah [616, 562]. Sharing [293]. Shaw [294]. Sherrington [825]. Shield [9, 721]. Shields [807]. Shift [428]. Shimmin [529, 539]. Shipbuilding [775, 376]. Shipping [775, 574, 234, 88, 775]. Shirley [580, 324]. Shirras [314, 35, 346, 450]. Short [379, 269]. Sickness [906, 212, 309, 840]. Sidney [222, 223, 58, 294, 876]. Siecle [831, 615].

Siècles [227]. Sigmund [310]. Silk [41]. Silver [73]. Simnett [405]. Simon [661, 106]. Simple [659]. Since [668, 783]. Sir [619, 68, 67, 901, 856, 358, 299, 300, 494, 21, 544, 648]. Sixteenth [527]. Sixth [79, 78]. Sketch [59]. Skilled [38]. Slavko [33]. Small [655]. Smith [63, 806, 849, 88, 541]. Smith-Gordon [63]. Smithsonian [466]. Snowden [124]. Social [225, 806, 421, 607, 14, 756, 198, 48, 717, 465, 295, 488]. sociali [381]. Socialisation [389]. Socialism [690]. Societies [428, 224]. Society [237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890, 79, 78, 157, 156, 237, 236, 335, 334, 419, 418, 504, 503, 604, 603, 705, 704, 795, 794, 891, 890, 734]. Soddy [402]. Soldiers [40]. Solution [635]. Some [3, 355, 771, 791, 601, 194, 341, 549, 73, 600, 329, 653, 685, 752, 243, 668, 336, 846, 104]. Sometimes [525]. Soper [265]. Sound [164, 480]. Sources [361, 641]. South [852]. Soviet [107, 563]. Spalding [289, 689]. Spaulding [709]. Special [202, 802, 821, 285, 591]. Spicer [827]. Spurgeon [286]. St [37]. Stabilisation [682, 738]. Stability [559]. Stabilization [864, 375, 508, 513]. Stacey [874]. Staff [490, 224, 759, 710]. Stamp [219, 469, 358, 607, 139, 662, 762, 846, 687, 708]. Standard [770, 628, 664]. Standardization [307]. Standpunkt [264]. Stanley [537, 339]. State [203, 472, 59, 850, 491, 584, 85, 62, 895, 612]. Stateement [707]. States [661, 828, 106, 226, 407, 580, 585, 140, 244, 338, 641, 367, 56, 195, 627, 274, 709, 55, 710, 733]. Statesmanship [57]. Stationary [824]. Statistica [119, 266, 7, 399]. Statistical [677, 79, 78, 22, 44, 69, 93, 157, 156, 113, 131, 149, 169, 237, 236, 191, 209, 230, 254, 335, 334, 278, 302, 325, 349, 419, 418, 372, 392, 411, 432, 504, 503, 458, 476, 496, 521, 604, 603, 545, 570, 597, 620, 705, 704, 649, 672, 696, 722, 795, 794, 747, 766, 787, 810, 891, 890, 834, 858, 880, 902, 353, 219, 441, 305, 602, 233, 33, 507, 578, 629, 527, 499, 675, 120, 159, 380, 821, 531, 819, 442, 487]. Statistico [381]. Statistics [5, 138, 421, 820, 771, 100, 215, 261, 699, 2, 414, 283, 479, 336, 776, 601, 25, 53, 574, 676, 382, 756, 97, 175, 462, 199, 751, 234, 13, 653, 199, 630, 892, 217, 658, 464, 532, 533, 579, 660, 753, 341, 685, 755, 556]. Statistiek [844]. Statistik [555, 310]. Statistique [845, 360, 452, 803, 158, 803, 818]. Statistische [264]. Statistischen [359]. Statutes [851]. Steel [516, 47]. Steffensen [706]. Stenning [389]. Stephen [294]. Stevens [783]. Stewart [300]. Still [865]. Still-Births [865]. Stimulus [762]. Stochastisch [264]. Stochastisch-Statistische [264]. Stock [72, 702]. Story [250]. Strakosch [513]. Strand [78]. Street [156, 236, 334, 418, 503, 603, 704, 794, 890]. Strobel [389]. Structure [756]. Stuart [225, 844]. Students [286, 660]. Studi [266]. Studies [361, 822, 535]. Study [870, 268, 364, 537, 716, 773, 802, 512, 146, 226, 291, 606, 731, 605, 777, 275, 321, 285, 292, 451, 140, 341, 717, 329, 238, 525, 448, 685, 468, 363]. Subercaseaux [342, 160]. Submerged [657]. Submitted [358]. Sugar [821]. Suggestion [164]. sui [869]. Summation [701]. Sumner [274]. Superiore [266]. Supplementary [214]. Supplies [84]. Supply [427, 821, 538, 26, 247]. Supplying [318]. Survey [886, 507, 756, 338, 806]. Surveyor [156]. Survive [143]. Sweden [414]. Sweeney [580]. Sykes [610]. System [143, 510, 195, 294, 448, 406, 510]. Système [242]. Systems [778].

T [853, 40, 516, 616, 162, 268, 364, 511, 562, 379, 664, 16, 850, 851, 317, 586, 140, 582, 687, 713, 56]. Table [137]. Tables [176, 233, 361, 76, 13, 177, 466]. Tackled [687]. Taken [480, 90]. Talbot [632]. Tannenbaum [898]. Tariff [80, 885, 710]. Tariffs [268]. Taussig [80, 774]. Tawney [640]. Tax [709, 481, 889]. Taxable [219, 562, 500]. Taxation [358, 481, 656, 889, 700, 139, 140, 549, 500, 163, 895, 186]. Taylor [690, 692]. Teachers [212, 309]. Tenney [108]. Teoria [381]. Terrace [78]. Test [355, 356, 843]. Tests [369, 320]. Textile [577, 539, 528, 653]. Their [365, 668, 375, 386, 53, 199, 421, 820, 776, 755, 556, 446]. Them [888]. Theorie [245]. Theories [805]. Theorique [360, 158]. Theory [50, 355, 707, 483, 512, 552, 4, 336, 420, 262, 33, 389, 873]. Third [705, 704]. Thomas [640, 850, 710]. Thornton [796, 140]. Thorp [686]. Thr [781]. Thrift [83, 56]. Thurstone [532]. Time [36, 499, 15, 152, 525]. Time-Correlation [152]. Time-Series [525]. Time-Unit [499]. Timothy [619]. To-day [294, 637]. To-Morrow [637, 186]. Tocher [587]. Tooke [850]. tot [844]. Total [361]. Totals [155]. Towns [518]. Tract [401]. Tracts [336]. Trade [80, 661, 138, 164, 406, 699, 221, 770, 58, 591, 588, 246, 376, 306, 600, 712, 512, 849, 271, 774, 31, 689, 868, 507]. Trades [47]. Trading [584]. Training [40]. Transfert [897]. Transition [146]. Transport [825, 853, 473]. Traqces [831]. Treatise [179, 379]. Treatment [142]. Trend [733]. Trenerry [691]. Triumph [126]. Truchy [761]. True [292]. Trufant [471]. Truman [380]. Trust [871]. Trusts [757, 291]. Tschuprow [264, 552, 596]. Tuesday [79, 157, 237, 335, 419, 504, 705, 795, 891]. Tuneside [826]. Turner [145]. Turroni [869]. Twelve [428]. Twelve-Hour [428]. Twin [823]. Twins [823]. Two [275, 575, 307].

U.S.A [687]. Udny [441, 483, 357].

Unemployed [875]. Unemployment [202, 817, 283, 454, 635, 367, 248, 283, 319, 343]. Union [563]. Unionism [164, 58]. Unit [499]. United [661, 828, 89, 274, 709, 106, 226, 55, 407, 580, 585, 140, 182, 244, 338, 710, 249, 367, 56, 733, 728, 72, 195, 627, 49]. University [272]. Unrest [480]. Unskilled [828]. Untried [416]. Unwin [109]. Upon [375, 640, 480, 585, 101, 344, 88]. Use [776]. Used [440]. Useful [4]. Uses [796]. Usury [640]. Uyehara [712].

V [138]. Vaccination [692, 824]. Valley [717]. Value [233, 142, 404]. Values [439]. Vanderblue [742]. Variability [894]. Variate [152]. Variate-Difference [152]. Variations [72, 51, 587]. Various [906, 840, 641, 48]. ved [557]. Vedel [408]. Vedel-Petersen [408]. Vehicles [573]. Venn [383]. Verrijn [844]. Very [439]. Vibart [583]. Vicious [166]. View [823]. Viner [512]. Vinici [399]. Vita [102]. Vital [421, 771, 414, 751]. Vocational [606]. Voies [831]. Vol [218, 847]. Volkseinkommen [663]. Volume [1, 71, 96, 151, 172, 232, 257, 328, 352, 413, 435, 498, 524, 599, 623, 698, 725, 789, 883, 388]. Vyle [690].

W [405, 156, 80, 873, 363, 667, 707, 109, 240, 241, 248, 289, 560, 609, 711, 774, 848, 892, 59, 777, 445, 775, 801, 34, 515, 660, 199, 244, 338, 423, 582, 713, 730, 141, 316, 448]. W.C [78, 236, 334, 418, 503, 603, 704, 794, 890]. Wade [17]. Wadia [562, 642]. Wage [537, 716]. Wage-Earning [537]. Wages [314, 490, 636, 668, 828, 153, 864, 182, 198, 827, 582, 874, 222, 798, 873, 566, 612, 316]. Wales [756, 573, 887, 414, 462, 48, 509]. Walker [710]. Wallace [713]. Wallis [198]. Walre [513]. Walsh [332]. Walter [122, 338, 403, 582]. Walton [874]. Wanlass [226]. War [761, 243, 103, 829, 750, 55, 240, 344, 408, 609, 700, 88, 135, 33, 491, 123, 140, 165, 338, 201, 425, 13, 3, 47, 72, 75, 51, 73, 26, 162, 106, 84, 39, 317, 181, 315, 86, 424, 36, 15, 56]. Ward [143]. Warren [471]. Wars [33]. Wealth [103, 219, 562, 715, 338, 751, 123, 219, 562]. Webb [223, 58, 294, 222, 223, 58, 876]. Webster [178]. Week [499]. Welding [659]. Welfare [144]. Wesley [661, 892, 244, 338, 754, 893]. West [528]. Westergaard [408]. Western [272, 847, 213]. Wheat [213, 134]. Which [398, 436]. White [404]. Whitney [828]. Whittaker [379]. Whitwell [275]. Whole [273]. Wholesale [74, 118, 196, 377, 463, 550, 654, 841, 661, 284, 772, 75, 51]. Widening [688]. Wider [378]. Willard [686]. Willford [407, 244, 338]. William [630, 146, 226, 800, 472, 489, 689, 742, 850, 494, 471, 713, 367, 638, 13, 737]. Willis [483]. Wilson [640, 275, 648]. Wissenschaft [264]. Withers [387, 36, 61]. Wolfe [427]. Wolfenden [820]. Women [758, 153]. Wood [20]. Woodbury [506, 733]. Wool [741, 539, 528, 653]. Woolf [14]. Woollen [128]. Work [225, 428, 798, 467]. Work-Periods [428]. Worker [368, 184]. Workers [876, 531, 819]. Working [203, 346, 283]. World [867, 106, 340, 702, 134, 667]. Worst [693]. Worsted [128].

XIII [138]. XIXe [227]. XVII [615]. XVIIIe [227].

Year [78, 156, 178, 236, 334, 418, 503, 603, 704, 794, 890, 664, 382]. Year-Book [382]. Years [728, 2, 275, 448]. Yesterday [186]. York [482]. Yorkshire [128, 528]. Young [644, 892]. Ysidro [569]. Yule [441, 483, 357]. Yurovsky [563]. Yves [765].

Zahlen [264]. Zealand [865, 285].

References

Anonymous:1920:VIa

[1] Anonymous. Volume information. Journal of the Royal Statistical Society, 83 (1):i–vii, January 1920. ISSN 0952-8385.

                                             URL http://www.jstor.org/stable/
                                             2340874.

Godfrey:1920:FYC

[2] Ernest H. Godfrey. Fifty years of Canadian progress as illustrated by official statistics, 1867 to 1917. Journal of the Royal Statistical Society, 83(1):1– 85, January 1920. ISSN 0952-8385.

                                             URL http://www.jstor.org/stable/
                                             2340875.

Allen:1920:SCD

[3] J. E. Allen. Some changes in the distribution of the national income during the War. Journal of the Royal Statistical Society, 83(1):86–126, January 1920. ISSN

                                             0952-8385. URL http://www.jstor.
                                             org/stable/2340876.

Guldberg:1920:EUT

[4] Alf. Guldberg. Expansions useful in the theory of frequency distributions. Journal of the Royal Statistical Society, 83 (1):127–130, January 1920. ISSN 0952-

                                             8385. URL http://www.jstor.org/
                                             stable/2340877.

Anonymous:1920:COS

[5] Anonymous. Committee on official statistics. Journal of the Royal Statistical Society, 83(1):131–133, January 1920. ISSN 0952-8385. URL http://

                                             www.jstor.org/stable/2340878.

Anonymous:1920:RC

[6] Anonymous. Report on the census. Journal of the Royal Statistical Society, 83(1):134–139, January 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340879.

I:1920:BRBa

[7] L. I. Book review: Elementi di Statistica, by Giorgio Mortara. Journal of the Royal Statistical Society, 83(1):140– 141, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340880.

C:1920:BRBa

[8] E. C. Book review: Currency and Credit, by R. G. Hawtrey. Journal of the Royal Statistical Society, 83(1):141–143, January 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340881.

H:1920:BRBa

[9] R. G. H. Book review: Inflation, by J. Shield Nicholson. Journal of the Royal Statistical Society, 83(1):144–145, January 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340882.

P:1920:BRBa

[10] L. L. P. Book review: Money and Prices, by J. Laurence Laughlin. Journal of the Royal Statistical Society, 83(1):145– 147, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340883.

H:1920:BRBb

[11] R. G. H. Book review: Organized Banking, by E. P. Agger. Journal of the Royal Statistical Society, 83(1):147–148, January 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340884.

W:1920:BRBa

[12] A. D. W. Book review: A Primer of National Finance, by Henry Higgs. Journal of the Royal Statistical Society, 83(1): 148–149, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340885.

W:1920:BRBb

[13] A. D. W. Book review: I. Commerce and Industry: A Historical Review of the Economic Conditions of the British Empire from the Peace of Paris in 1815 to the Declaration of war in 1914, Based on Parliamentary Debates, by William Page; William Ashley; Commerce and Industry: Tables of Statistics for the British Empire from 1815 by William Page. Journal of the Royal Statistical Society, 83(1):150–151, January 1920. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340886.

M:1920:BRBa

[14] H. W. M. Book review: Co-operation and the Future of Industry, by Leonard S. Woolf; Guilds and the Social Crisis by Arthur J. Penty. Journal of the Royal Statistical Society, 83(1):151–152, January 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340887.

W:1920:BRBc

[15] F. W. Book review: War Time Control of Industry: The Experience of England, by Howard L. Gray. Journal of the Royal Statistical Society, 83(1):152–154, January 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340888.

M:1920:BRBb

[16] H. W. M. Book review: The Costing Problem, by Edward T. Elborne. Journal of the Royal Statistical Society, 83 (1):154–156, January 1920. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340889.

P:1920:BRBb

[17] L. L. P. Book review: Australia: Problems and Prospects, by Charles Wade. Journal of the Royal Statistical Society, 83(1):156–158, January 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340890.

Anonymous:1920:ONPa

[18] Anonymous. Other new publications. Journal of the Royal Statistical Society, 83(1):158–161, January 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340891.

Anonymous:1920:CNa

[19] Anonymous. Current notes. Journal of the Royal Statistical Society, 83(1): 162–177, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340892.

G:1920:FW

[20] M. G. Frances Wood, 1883–1919. Journal of the Royal Statistical Society, 83 (1):178–180, January 1920. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340893.

R:1920:SRC

[21] R. H. R. Sir Richard Crawford. Journal of the Royal Statistical Society, 83(1): 180–183, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340894.

Anonymous:1920:SEAa

[22] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 83(1): 184–188, January 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340895.

Anonymous:1920:LALa

[23] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 83(1):189–194, January 1920. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340896.

Anonymous:1920:PRa

[24] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 83(1): 195–200, January 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340897.

Knibbs:1920:OIS

[25] G. H. Knibbs. The organisation of Imperial statistics. Journal of the Royal Statistical Society, 83(2):201–224, March 1920. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341078.

Starling:1920:FSG

[26] Ernest H. Starling. The food supply of Germany during the War. Journal of the Royal Statistical Society, 83(2): 225–254, March 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341079.

Greenwood:1920:INF

[27] Major Greenwood and G. Udny Yule. An inquiry into the nature of frequency distributions representative of multiple happenings with particular reference to the occurrence of multiple attacks of disease or of repeated accidents. Journal of the Royal Statistical Society, 83(2): 255–279, March 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341080.

Brownlee:1920:DDR

[28] John Brownlee. Density and death-rate: Farr’s law. Journal of the Royal Statistical Society, 83(2):280–283, March 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341081.

Anonymous:1920:PP

[29] Anonymous. The paper pound of 1797– 1821. Journal of the Royal Statistical Society, 83(2):284–289, March 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341082.

I:1920:BRBb

[30] L. I. Book review: Calcolo delle Probabilità, by Guido Castelnuovo. Journal of the Royal Statistical Society, 83(2): 290–292, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341083.

M:1920:BRBc

[31] H. W. M. Book review: Industry and Trade, by Alfred Marshall. Journal of the Royal Statistical Society, 83(2): 292–296, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341084.

M:1920:BRBd

[32] H. O. M. Book review: The Economic Consequences of the Peace, by John Maynard Keynes. Journal of the Royal Statistical Society, 83(2):296– 298, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341085.

R:1920:BRB

[33] R. H. R. Book review: Economic Phenomena Before and After War: A Statistical Theory of Modern Wars, by Slavko Seárov. Journal of the Royal Statistical Society, 83(2):298–299, March 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341086.

P:1920:BRBc

[34] L. L. P. Book review: Industry and Finance, by Adam W. Kirkaldy. Journal of the Royal Statistical Society, 83(2): 300–302, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341087.

H:1920:BRBc

[35] R. G. H. Book review: Indian Finance and Banking, by G. Findlay Shirras. Journal of the Royal Statistical Society, 83(2):302–303, March 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341088.

S:1920:BRBa

[36] J. C. S. Book review: War Time Financial Problems, by Hartley Withers. Journal of the Royal Statistical Society, 83(2):304–305, March 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341089.

H:1920:BRBd

[37] R. G. H. Book review: The Physiology of Credit and of Money, by Oswald St. Clair. Journal of the Royal Statistical Society, 83(2):305–306, March 1920. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341090.

P:1920:BRBd

[38] L. L. P. Book review: The Skilled Labourer, 1760–1832, by J. L. Hammond; Barbara Hammond. Journal of the Royal Statistical Society, 83(2): 306–309, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341091.

P:1920:BRBe

[39] L. L. P. Book review: British Labour Conditions and Legislation during the War, by M. B. Hammond. Journal of the Royal Statistical Society, 83(2): 309–311, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341092.

B:1920:BRBa

[40] E. B. Book review: Discharged Soldiers and Sailors Pensions and Training, by Edward T. Devine. Journal of the Royal Statistical Society, 83(2): 312–313, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341093.

F:1920:BRB

[41] A. F. Book review: Economics of the Silk Industry, by Ratan C. Rawlley. Journal of the Royal Statistical Society, 83(2):313–315, March 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341094.

Anonymous:1920:ONPb

[42] Anonymous. Other new publications. Journal of the Royal Statistical Society, 83(2):316–318, March 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341095.

Anonymous:1920:CNb

[43] Anonymous. Current notes. Journal of the Royal Statistical Society, 83(2): 319–333, March 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341096.

Anonymous:1920:SEAb

[44] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 83(2): 334–336, March 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341097.

Anonymous:1920:LALb

[45] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 83(2):337–340, March 1920. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341098.

Anonymous:1920:PRb

[46] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 83(2): 341–349, March 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341099.

Birkett:1920:IST

[47] M. S. Birkett. The iron and steel trades during the War. Journal of the Royal Statistical Society, 83(3):351–400, May 1920. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340957.

Stevenson:1920:FVS

[48] T. H. C. Stevenson. The fertility of various social classes in England and Wales from the middle of the Nineteenth Century to 1911. Journal of the Royal Statistical Society, 83(3):401–444, May 1920. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340958.

Moore:1920:CCU

[49] Henry L. Moore. Crop-cycles in the United Kingdom and in France. Journal of the Royal Statistical Society, 83 (3):445–454, May 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340959.

Bennet:1920:TMC

[50] T. L. Bennet. The theory of measurement of changes in cost of living. Journal of the Royal Statistical Society, 83 (3):455–462, May 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340960.

Ottolenghi:1920:VWP

[51] Costantino Ottolenghi. The variations of wholesale prices in Italy during the Great War. Journal of the Royal Statistical Society, 83(3):463–477, May 1920. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340961.

Turin:1920:MPC

[52] S. P. Turin. Market prices and controlled prices of food in Moscow. Journal of the Royal Statistical Society, 83(3):478–479, May 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340962.

Koren:1920:BRB

[53] John Koren and G. U. Y. Book review: The History of Statistics: Their Development and Progress in Many Countries. Journal of the Royal Statistical Society, 83(3):480–482, May 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340963.

E:1920:BRB

[54] F. Y. E. Book review: The Change in the Distribution of the National Income, 1880–1913, by Arthur L. Bowley. Journal of the Royal Statistical Society, 83(3):482–484, May 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340964.

H:1920:BRBe

[55] R. G. H. Book review: Effects of the War on Money, Credit and Banking in France and the United States, by B. M. Anderson. Journal of the Royal Statistical Society, 83(3):484–487, May 1920. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340965.

W:1920:BRBd

[56] A. D. W. Book review: War Thrift, by T. N. Carver; Government Control of the Liquor Business in Great Britain and the United States, by T. N. Carver. Journal of the Royal Statistical Society, 83(3):487–489, May 1920. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340966.

P:1920:BRBf

[57] L. L. P. Book review: Economic Statesmanship, by J. Ellis Barker. Journal of the Royal Statistical Society, 83(3): 489–491, May 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340967.

P:1920:BRBg

[58] L. L. P. Book review: The History of Trade Unionism, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 83(3):491–493, May 1920. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340968.

J:1920:BRB

[59] W. H. J. Book review: Historical Sketch of State Railway Ownership, by W. M. Acworth. Journal of the Royal Statistical Society, 83(3):493–495, May 1920. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340969.

C:1920:BRBb

[60] E. C. Book review: Nationalisation of the Mines, by Frank Hodges. Journal of the Royal Statistical Society, 83(3): 495–496, May 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340970.

S:1920:BRBb

[61] J. C. S. Book review: The Case for Capitalism, by Hartley Withers. Journal of the Royal Statistical Society, 83 (3):496–498, May 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340971.

B:1920:BRBb

[62] J. B. Book review: National Guilds and the State, by S. G. Hobson. Journal of the Royal Statistical Society, 83(3): 498–499, May 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340972.

B:1920:BRBc

[63] E. B. Book review: Co-Operation in Scotland, by James Lucas; Co-Operation in Denmark by L. Smith-Gordon; C. O’Brien. Journal of the Royal Statistical Society, 83(3):499–501, May 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340973.

B:1920:BRBd

[64] J. A. B. Book review: Essays on Indian Economic Problems, by Brij Narain. Journal of the Royal Statistical Society, 83(3):501–502, May 1920. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340974.

Anonymous:1920:ONPc

[65] Anonymous. Other new publications. Journal of the Royal Statistical Society, 83(3):503–506, May 1920. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340975.

Anonymous:1920:CNc

[66] Anonymous. Current notes. Journal of the Royal Statistical Society, 83(3): 507–520, May 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340976.

Anonymous:1920:SJG

[67] Anonymous. Sir John Glover. Journal of the Royal Statistical Society, 83(3):521, May 1920. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340977.

Anonymous:1920:SHB

[68] Anonymous. Sir Henry Burdett. Journal of the Royal Statistical Society, 83 (3):521–522, May 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340978.

Anonymous:1920:SEAc

[69] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 83(3): 523–527, May 1920. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340979.

Anonymous:1920:LALc

[70] Anonymous. List of aditions to the library. Journal of the Royal Statistical Society, 83(3):528–532, May 1920. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340980.

Anonymous:1920:VIb

[71] Anonymous. Volume information. Journal of the Royal Statistical Society, 83 (4):724–733, July 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340898.

Guild:1920:VNL

[72] J. B. Guild. Variations in the numbers of live stock and in the production of meat in the United Kingdom during the War. Journal of the Royal Statistical Society, 83(4):533–571, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340899.

Shirras:1920:SEW

[73] G. Findlay Shirras. Some effects of the War on gold and silver. Journal of the Royal Statistical Society, 83(4):572–627, July 1920. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340900.

Anonymous:1920:WPC

[74] Anonymous. Wholesale prices of commodities in 1919. Journal of the Royal Statistical Society, 83(4):628–645, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340901.

Ottolenghi:1920:INW

[75] Costantino Ottolenghi. Index-numbers of wholesale prices in Italy during the Great War. Journal of the Royal Statistical Society, 83(4):646–655, July 1920. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340902.

Trachtenberg:1920:RLT

[76] H. L. Trachtenberg. The relation of lifetables to the Makeham Law. Journal of the Royal Statistical Society, 83(4):

656–661, July 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340903.

Anonymous:1920:LMA

[77] Anonymous. Letter from M. Arthur Raffalovich. Journal of the Royal Statistical Society, 83(4):662, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340904.

Anonymous:1920:RCF

[78] Anonymous. Report of the Council for the financial year ended December 31, 1919, and for the sessional year ending June 15, 1920, presented at the eighty-sixth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C. 2, on June 15, 1920. Journal of the Royal Statistical Society, 83(4):663–675, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340905.

Anonymous:1920:PES

[79] Anonymous. Proceedings of the eightysixth Annual General Meeting of the Royal Statistical Society, held on Tuesday, June 15, 1920. Journal of the Royal Statistical Society, 83(4):676, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340906.

B:1920:BRBe

[80] J. B. Book review: Free Trade, the Tariff, and Reciprocity, by F. W. Taussig. Journal of the Royal Statistical Society, 83(4):677–680, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340907.

P:1920:BRBh

[81] L. L. P. Book review: Life and Labour in the Nineteenth Century, by C. R. Fay. Journal of the Royal Statistical Society, 83(4):680–683, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340908.

C:1920:BRBc

[82] E. C. Book review: Economic Liberty, by Harold Cox. Journal of the Royal Statistical Society, 83(4):683–684, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340909.

G:1920:BRB

[83] J. B. G. Book review: The Nation’s Food, by Raymond Pearl; Food Thrift by Raymond Pearl. Journal of the Royal Statistical Society, 83(4):684–687, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340910.

M:1920:BRBe

[84] H. W. M. Book review: Food Supplies in Peace and War, by R. Henry Rew. Journal of the Royal Statistical Society, 83(4):687–689, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340911.

S:1920:BRBc

[85] J. C. S. Book review: The State Debt and the National Capital. A New Proposal for Reconstruction and Redemption. Journal of the Royal Statistical Society, 83(4):689–691, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340912.

H:1920:BRBf

[86] R. G. H. Book review: Currencies after the War. Journal of the Royal Statistical Society, 83(4):691–692, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340913.

T:1920:BRB

[87] C. T. Book review: Le Peripezie Monetaire della Guerra, by Achille Loria. Journal of the Royal Statistical Society, 83(4):692–693, July 1920. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340914.

M:1920:BRBf

[88] H. W. M. Book review: Influence of the Great War upon Shipping, by J. Russell Smith. Journal of the Royal Statistical Society, 83(4):693–694, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340915.

C:1920:BRBd

[89] C. E. C. Book review: Co-Operation at Home and Abroad: A Description and Analysis. The Progress of Co-operation in the United Kingdom (1908–1918), by C. R. Fay. Journal of the Royal Statistical Society, 83(4):694–696, July 1920. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340916.

Y:1920:BRB

[90] G. U. Y. Book review: Problems of Population and Parenthood: Being the Second Report of and the Chief Evidence taken, by the National Birth-Rate Commission, 1918–20. Journal of the Royal Statistical Society, 83(4):696–697, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340917.

Anonymous:1920:ONPd

[91] Anonymous. Other new publications. Journal of the Royal Statistical Society, 83(4):697–700, July 1920. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340918.

Anonymous:1920:CNd

[92] Anonymous. Current notes. Journal of the Royal Statistical Society, 83(4): 701–710, July 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340919.

Anonymous:1920:SEAd

[93] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 83(4): 711–713, July 1920. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340920.

Anonymous:1920:LALd

[94] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 83(4):714–717, July 1920. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340921.

Anonymous:1920:ALA

[95] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 83(4):718–723, July 1920. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340922.

Anonymous:1921:VIa

[96] Anonymous. Volume information. Journal of the Royal Statistical Society, 84 (1):i–vii, January 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340587.

Rew:1921:OS

[97] R. Henry Rew. The organization of statistics. Journal of the Royal Statistical Society, 84(1):1–21, January 1921.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340588.

Anonymous:1921:PN

[98] Anonymous. Proceedings on November 16, 1920. Journal of the Royal Statistical Society, 84(1):22–25, January 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340589.

Fraser:1921:MD

[99] Fraser. The maturing debt. Journal of the Royal Statistical Society, 84(1): 26–70, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340590.

Edgeworth:1921:MS

[100] F. Y. Edgeworth. Molecular statistics. Journal of the Royal Statistical Society, 84(1):71–89, January 1921. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340591.

Stevenson:1921:IMU

[101] T. H. C. Stevenson. The incidence of mortality upon the Rich and poor districts of Paris and London. Journal of the Royal Statistical Society, 84(1): 90–99, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340592.

I:1921:BRBa

[102] L. I. Book review: La Misura Della Vita, by Alfredo Niceforo. Journal of the Royal Statistical Society, 84(1):100– 101, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340593.

C:1921:BRBa

[103] E. C. Book review: Report from the Select Committee on Increase of Wealth (War); with Proceedings, Minutes and Appendices. Journal of the Royal Statistical Society, 84(1):101–103, January 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340594.

S:1921:BRBa

[104] J. C. S. Book review: Some Aspects of the Inequality of Incomes in Modern Communities, by Hugh Dalton. Journal of the Royal Statistical Society, 84(1): 103–106, January 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340595.

S:1921:BRBb

[105] J. C. S. Book review: English Public Finance from the Revolution of 1688. With Chapters on the Bank of England, by Harvey E. Fisk. Journal of the Royal Statistical Society, 84(1):106–108, January 1921. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340596.

G:1921:BRBa

[106] J. B. G. Book review: Prices and Price Control in Great Britain and the United States During the World War, by Simon Litman. Journal of the Royal Statistical Society, 84(1):108–109, January 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340597.

G:1921:BRBb

[107] J. B. G. Book review: Labour Conditions in Soviet Russia. Journal of the Royal Statistical Society, 84(1):109– 111, January 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340598.

T:1921:BRBa

[108] C. T. Book review: An Economic History of Rome, to the end of the Republic, by Tenney Frank. Journal of the Royal Statistical Society, 84(1):111, January 1921. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340599.

F:1921:BRBa

[109] A. F. Book review: The Early English Cotton Industry, by George W. Daniels; George Unwin. Journal of the Royal Statistical Society, 84(1):112–113, January 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340600.

Anonymous:1921:ONPa

[110] Anonymous. Other new publications. Journal of the Royal Statistical Society, 84(1):113–117, January 1921. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340601.

Anonymous:1921:CNa

[111] Anonymous. Current notes. Journal of the Royal Statistical Society, 84(1): 118–134, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340602.

Craigie:1921:LB

[112] P. G. Craigie. Luigi Bodio. Journal of the Royal Statistical Society, 84(1): 135–136, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340603.

Anonymous:1921:SEAa

[113] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 84(1): 137–144, January 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340604.

Anonymous:1921:LALa

[114] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 84(1):145–158, January 1921. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340605.

Anonymous:1921:PRa

[115] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 84(1): 159–165, January 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340606.

Flux:1921:MPC

[116] A. W. Flux. The measurement of price changes. Journal of the Royal Statistical Society, 84(2):167–215, March 1921. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340499.

Bonar:1921:MPM

[117] J. Bonar. The mint and the precious metals in Canada. Journal of the Royal Statistical Society, 84(2):216– 254, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340500.

Anonymous:1921:WPC

[118] Anonymous. Wholesale prices of commodities in 1920. Journal of the Royal Statistical Society, 84(2):255– 277, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340501.

I:1921:BRBb

[119] L. I. Book review: Lezioni di Statistica Metodologica, by Filadelfo Insolera. Journal of the Royal Statistical Society, 84(2):278–279, March 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340502.

Y:1921:BRBa

[120] G. U. Y. Book review: Readings and Problems in Statistical Methods, by Horace Secrist. Journal of the Royal Statistical Society, 84(2):279–280, March 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340503.

H:1921:BRBa

[121] W. H. H. Book review: Report on the Pandemic of Influenza, 1918–19. Journal of the Royal Statistical Society, 84 (2):280–283, March 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340504.

S:1921:BRBc

[122] J. C. S. Book review: High Prices and Deflation, by Edwin Walter Kemmerer. Journal of the Royal Statistical Society, 84(2):283–284, March 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340505.

S:1921:BRBd

[123] J. C. S. Book review: A Capital Levy and a Levy on War Wealth, by A. C. Pigou. Journal of the Royal Statistical Society, 84(2):284–285, March 1921. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340506.

S:1921:BRBe

[124] J. C. S. Book review: Labour and National Finance, by Philip Snowden. Journal of the Royal Statistical Society, 84(2):285–287, March 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340507.

B:1921:BRBa

[125] J. B. Book review: Credit-Power and Democracy, with a Draft Scheme for the Mining Industry, by C. H. Douglas. Journal of the Royal Statistical Society, 84(2):287, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340508.

B:1921:BRBb

[126] J. B. Book review: The Triumph of Nationalization, by Leo Chiozza. Journal of the Royal Statistical Society, 84(2): 288–289, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340509.

J:1921:BRB

[127] W. H. J. Book review: Industrial Control, by F. M. Lawson. Journal of the Royal Statistical Society, 84(2): 289–290, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340510.

F:1921:BRBb

[128] A. F. Book review: The Yorkshire Woollen and Worsted Industries, by Herbert Heaton. Journal of the Royal Statistical Society, 84(2):290– 293, March 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340511.

Anonymous:1921:ONPb

[129] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 84(2):293–296, March 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340512.

Anonymous:1921:CNb

[130] Anonymous. Current notes. Journal of the Royal Statistical Society, 84(2):

297–309, March 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340513.

Anonymous:1921:SEAb

[131] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 84(2): 310–313, March 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340514.

Anonymous:1921:LALb

[132] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 84(2):314–318, March 1921. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340515.

Anonymous:1921:PRb

[133] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 84(2): 319–327, March 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340516.

Wilson:1921:WW

[134] James Wilson. The world’s wheat. Journal of the Royal Statistical Society, 84 (3):329–391, May 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340795.

Pearson:1921:CPW

[135] R. J. A. Pearson. A comparison of pre-war and post-war production costs in engineering. Journal of the Royal Statistical Society, 84(3):392–425, May 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340796.

Crump:1921:RRF

[136] Norman Crump. A review of recent foreign exchange fluctuations. Journal of the Royal Statistical Society, 84(3): 426–452, May 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340797.

Crathorne:1921:BLT

[137] A. R. Crathorne and Margaret Walker.

The biology of a life-table. Journal of the Royal Statistical Society, 84(3):453–456, May 1921. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340798.

C:1921:BRBb

[138] E. C. Book review: Brussels International Financial Conference, 1920. Documents of the Conference, (1) Nos. III, IV and V, Currency Statistics, Public Finance, International Trade. ; Brussels International Financial Conference, 1920.(2) No. XIII, Memoranda of Economic Experts. Journal of the Royal Statistical Society, 84(3):457–459, May 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340799.

M:1921:BRBa

[139] B. M. Book review: The Fundamental Principles of Taxation in the Light of Modern Developments, by Josiah Stamp. Journal of the Royal Statistical Society, 84(3):459–462, May 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340800.

S:1921:BRBf

[140] J. C. S. Book review: British War Finance and the Consequences, by T. J. Kiernan; The Nation’s Financial Outlook. by A. B. Thornton; The Taxation of Excess Profits in Great Britain: A Study of the British Excess Profits Duty in Relation to the Problem of Excess Profits. Taxation in the United States. by Robert Murray Haig; George E. Holmes. Journal of the Royal Statistical Society, 84(3):462–463, May 1921. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340801.

T:1921:BRBb

[141] C. T. Book review: The Bankers’ Clearing House: What it is and What it Does, by Philip W. Matthews; R. Holland-Martin. Journal of the Royal Statistical Society, 84(3):463–464, May 1921. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340802.

S:1921:BRBg

[142] J. C. S. Book review: Commercial Goodwill–Its History, Value and Treatment in Accounts, by P. D. Leake. Journal of the Royal Statistical Society, 84 (3):464–466, May 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340803.

G:1921:BRBc

[143] J. B. G. Book review: Can Our Industrial System Survive?, by J. S. M. Ward. Journal of the Royal Statistical Society, 84(3):466, May 1921. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340804.

M:1921:BRBb

[144] H. W. M. Book review: The Economics of Welfare, by A. C. Pigou. Journal of the Royal Statistical Society, 84(3): 466–468, May 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340805.

M:1921:BRBc

[145] H. W. M. Book review: Introduction to Economics, by John Roscoe Turner.

Journal of the Royal Statistical Society, 84(3):468–470, May 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340806.

G:1921:BRBd

[146] M. G. Book review: England in Transition, 1789–1832: A Study of Movements, by William Law Mathieson. Journal of the Royal Statistical Society, 84(3):470–471, May 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340807.

Anonymous:1921:ONPc

[147] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 84(3):471–472, May 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340808.

Anonymous:1921:CNc

[148] Anonymous. Current notes. Journal of the Royal Statistical Society, 84(3): 473–488, May 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340809.

Anonymous:1921:SEAc

[149] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 84(3): 489–492, May 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340810.

Anonymous:1921:LALc

[150] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 84(3):493–496, May 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340811.

Anonymous:1921:VIb

[151] Anonymous. Volume information. Journal of the Royal Statistical Society, 84 (4):677–684, July 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341100.

Yule:1921:TCP

[152] G. Udny Yule. On the time-correlation problem, with especial reference to the variate-difference correlation method. Journal of the Royal Statistical Society, 84(4):497–537, July 1921. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341101.

Barton:1921:WMW

[153] Dorothea M. Barton. Women’s minimum wages. Journal of the Royal Statistical Society, 84(4):538–577, July 1921. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341102.

Trachtenberg:1921:AP

[154] H. L. Trachtenberg. The analysis of the periodogram. Journal of the Royal Statistical Society, 84(4):578–604, July 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341103.

Baines:1921:PTC

[155] J. A. Baines. The provisional totals of the census of India. Journal of the Royal Statistical Society, 84(4):605–606, July 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341104.

Anonymous:1921:RCF

[156] Anonymous. Report of the Council for the financial year ended December 31, 1920, and for the sessional year ending June 21, 1921, presented at the eightyseventh Annual General Meeting of the Royal Statistical Society, held at the Surveyor’s Institution, 12, Great George Street, S. W. 1, on June 21, 1921. Journal of the Royal Statistical Society, 84 (4):607–618, July 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341105.

Anonymous:1921:PES

[157] Anonymous. Proceedings of the eightseventh Annual General Meeting of the Royal Statistical Society, held on Tuesday, June 21, 1921. Journal of the Royal Statistical Society, 84(4):619–620, July 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341106.

Y:1921:BRBb

[158] G. U. Y. Book review: Principes de Statistique Théorique et Appliquée, by Armand Julin; M. Lucien March. Journal of the Royal Statistical Society, 84 (4):621–625, July 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341107.

Y:1921:BRBc

[159] G. U. Y. Book review: The Problem of Estimation: A Seventeenth-Century Controversy and its Bearing on Modern Statistical Questions, Especially Index-Numbers, by Correa Moylan. Journal of the Royal Statistical Society, 84(4): 625–626, July 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341108.

H:1921:BRBb

[160] R. G. H. Book review: Le Papier Monnaie, by G. Subercaseaux. Journal of the Royal Statistical Society, 84(4):626–628, July 1921. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341109.

B:1921:BRBc

[161] C. F. B. Book review: Financial Organization, by H. G. Moulton. Journal of the Royal Statistical Society, 84(4):628–629, July 1921. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341110.

C:1921:BRBc

[162] N. E. C. Book review: Foreign Exchange, Before, During, and After the War, by T. E. Gregory. Journal of the Royal Statistical Society, 84(4):630–631, July 1921. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341111.

B:1921:BRBd

[163] J. B. Book review: Taxation and the Cost of Living. Journal of the Royal Statistical Society, 84(4):631–633, July 1921. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341112.

C:1921:BRBd

[164] C. E. C. Book review: Is Trade Unionism Sound? A Suggestion for Outflanking the Power of Capital, by J. H. Bunting. Journal of the Royal Statistical Society, 84(4):633–634, July 1921. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341113.

S:1921:BRBh

[165] E. C. S. Book review: Government War Contracts, by J. Franklin Crowell. Journal of the Royal Statistical Society, 84 (4):634–635, July 1921. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341114.

G:1921:BRBe

[166] M. G. Book review: Poverty and Its Vicious Circles, by Jamieson B. Hurry.

Journal of the Royal Statistical Society, 84(4):636–637, July 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341115.

Anonymous:1921:ONPd

[167] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 84(4):637–640, July 1921. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341116.

Anonymous:1921:CNd

[168] Anonymous. Current notes. Journal of the Royal Statistical Society, 84(4): 641–660, July 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341117.

Anonymous:1921:SEAd

[169] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 84(4): 661–664, July 1921. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341118.

Anonymous:1921:LALd

[170] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 84(4):665–669, July 1921. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341119.

Anonymous:1921:ALA

[171] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 84(4):670–676, July 1921. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341120.

Anonymous:1922:VIa

[172] Anonymous. Volume information. Journal of the Royal Statistical Society, 85 (1):i–vii, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340517.

Rew:1922:PBA

[173] R. Henry Rew. The progress of British agriculture. Journal of the Royal Statistical Society, 85(1):1–19, January 1922. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340518.

Anonymous:1922:PM

[174] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 85(1):20–26, January 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340519.

Rusher:1922:SIM

[175] Edward A. Rusher. The statistics of industrial morbidity in Great Britain. a retrospect and a scheme for development. Journal of the Royal Statistical Society, 85(1):27–86, January 1922. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340520.

Fisher:1922:ICT

[176] R. A. Fisher. On the interpretation of χ^{2} from contingency tables, and the calculation of P. Journal of the Royal Statistical Society, 85(1):87–94, January 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340521.

Yule:1922:AMA

[177] G. Udny Yule. On the application of the χ^{2} method to association and contingency tables, with experimental illustrations. Journal of the Royal Statistical Society, 85(1):95–104, January 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340522.

Anonymous:1922:NAN

[178] Anonymous. Note on an account of the number of people in Scotland in the year 1755 by Alexander Webster, one of the Ministers of Edinburgh. Journal of the Royal Statistical Society, 85(1): 105–106, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340523.

E:1922:BRB

[179] F. Y. E. Book review: A Treatise on Probability, by John Maynard Keynes. Journal of the Royal Statistical Society, 85(1):107–113, January 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340524.

B:1922:BRBa

[180] J. B. Book review: Government and Industry, by C. Delisle Burns. Journal of the Royal Statistical Society, 85(1): 113–117, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340525.

S:1922:BRBa

[181] J. C. S. Book review: The Political Economy of War, by A. C. Pigou. Journal of the Royal Statistical Society, 85 (1):117–119, January 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340526.

S:1922:BRBb

[182] E. C. S. Book review: Prices and Wages in the United Kingdom, 1914–1920, by Arthur L. Bowley. Journal of the Royal Statistical Society, 85(1):119–121, January 1922. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340527.

A:1922:BRB

[183] J. E. A. Book review: Final Report on the Cost of Living. Journal of the Royal Statistical Society, 85(1):121–124, January 1922. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340528.

D:1922:BRBa

[184] R. D. Book review: The Health of the Industrial Worker, by Edgar L. Collis; Major Greenwood; G. Newman. Journal of the Royal Statistical Society, 85(1): 124–126, January 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340529.

T:1922:BRBa

[185] C. T. Book review: The Advancement of Science, 1921. Journal of the Royal Statistical Society, 85(1):126–127, January 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340530.

S:1922:BRBc

[186] J. C. S. Book review: Taxation: Yesterday and To-morrow, by Robert Jones. Journal of the Royal Statistical Society, 85(1):127–128, January 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340531.

W:1922:BRB

[187] J. S. M. W. Book review: The Economic Development of France and Germany, 1815 to 1914, by J. H. Clapham. Journal of the Royal Statistical Society, 85(1): 128–130, January 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2340532.

F:1922:BRB

[188] A. F. Book review: The Cotton Industry in France, by R. B. Forrester; D. H.

Macgregor. Journal of the Royal Statistical Society, 85(1):130–131, January 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340533.

Anonymous:1922:ONPa

[189] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 85(1):131–135, January 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340534.

Anonymous:1922:CNa

[190] Anonymous. Current notes. Journal of the Royal Statistical Society, 85(1): 136–150, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340535.

Anonymous:1922:SEAa

[191] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 85(1): 151–160, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340536.

Anonymous:1922:LALa

[192] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 85(1):161–169, January 1922. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340537.

Anonymous:1922:PRa

[193] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 85(1): 170–176, January 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340538.

Macrosty:1922:SCF

[194] H. W. Macrosty. Some current financial problems. Journal of the Royal Statistical Society, 85(2):177–223, March 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341164.

Hawtrey:1922:FRS

[195] R. G. Hawtrey. The Federal reserve system of the United States. Journal of the Royal Statistical Society, 85(2): 224–269, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341165.

Anonymous:1922:WPC

[196] Anonymous. Wholesale prices of commodities in 1921. Journal of the Royal Statistical Society, 85(2):270– 292, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341166.

Anonymous:1922:FG

[197] Anonymous. Francis Galton. Journal of the Royal Statistical Society, 85(2): 293–298, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341167.

S:1922:BRBd

[198] E. C. S. Book review: Prices and Wages. An Investigation of the Dynamic Forces in Social Economics, by Percy Wallis; Albert Wallis; P. S. King. Journal of the Royal Statistical Society, 85(2): 299–302, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341168.

S:1922:BRBe

[199] J. C. S. Book review: Business Statistics: Their Preparation, Compliation and Presentation, by R. W. Holland; Statistics and their Application to Commerce, by A. Lester Boddington. Journal of the Royal Statistical Society, 85 (2):302–303, March 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341169.

S:1922:BRBf

[200] J. C. S. Book review: The English Capital Market, by F. Lavington. Journal of the Royal Statistical Society, 85(2): 303–304, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341170.

T:1922:BRBb

[201] C. T. Book review: The Effects of the War on Credit, Currency, Finance, and Foreign Exchanges. Journal of the Royal Statistical Society, 85(2):305– 306, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341171.

B:1922:BRBb

[202] W. A. B. Book review: Insurance Against Unemployment, with Special Reference to British and American Conditions, by Joseph I. Cohen. Journal of the Royal Statistical Society, 85(2): 306–307, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341172.

C:1922:BRBa

[203] C. R. C. Book review: Report of the Committee Appointed, by the Secretary of State for India on the Administration and Working of Indian Railways. Journal of the Royal Statistical Society, 85 (2):308–310, March 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341173.

D:1922:BRBb

[204] R. D. Book review: The Biology of Death, by Raymond Pearl. Journal of the Royal Statistical Society, 85(2): 310–311, March 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341174.

R:1922:BRB

[205] R. H. R. Book review: English Farming, Past and Present, by Lord Ernle. Journal of the Royal Statistical Society, 85(2):311–312, March 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341175.

G:1922:BRBa

[206] J. B. G. Book review: Farming Costs, by C. S. Orwin. Journal of the Royal Statistical Society, 85(2):312–313, March 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341176.

Anonymous:1922:ONPb

[207] Anonymous. Other new publications. Journal of the Royal Statistical Society, 85(2):313–316, March 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341177.

Anonymous:1922:CNb

[208] Anonymous. Current notes. Journal of the Royal Statistical Society, 85(2): 317–329, March 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341178.

Anonymous:1922:SEAb

[209] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 85(2): 330–333, March 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341179.

Anonymous:1922:LALb

[210] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 85(2):334–339, March 1922. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341180.

Anonymous:1922:PRb

[211] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 85(2): 340–348, March 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341181.

Hart:1922:ISD

[212] J. Y. Hart. An investigation of sickness data of public elementary school teachers in London, 1904–1919. Journal of the Royal Statistical Society, 85(3):349–411, May 1922. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341182.

Beveridge:1922:WPR

[213] William H. Beveridge. Wheat prices and rainfall in Western Europe. Journal of the Royal Statistical Society, 85 (3):412–475, May 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341183.

Beveridge:1922:SN

[214] William H. Beveridge. Supplementary notes. Journal of the Royal Statistical Society, 85(3):475–478, May 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341184.

Edgeworth:1922:MS

[215] F. Y. Edgeworth. Molecular statistics.

Journal of the Royal Statistical Society, 85(3):479–487, May 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341185.

Bispham:1922:NHF

[216] J. W. Bispham. Note on a heterotypic frequency distribution. Journal of the Royal Statistical Society, 85(3):488–492, May 1922. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341186.

I:1922:BRBa

[217] L. I. Book review: A First Course in Statistics, by D. Caradog Jones. Journal of the Royal Statistical Society, 85 (3):493–494, May 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341187.

I:1922:BRBb

[218] L. I. Book review: Lezioni di Matematica Finaziaria. Vol. I, by Luigi Amoroso. Journal of the Royal Statistical Society, 85(3):494–495, May 1922. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341188.

C:1922:BRBb

[219] E. C. Book review: Wealth and Taxable Capacity: The Newmarch Lectures for 1920–1 on Current Statistical Problems in Wealth and Industry, by Josiah Stamp. Journal of the Royal Statistical Society, 85(3):495–496, May 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341189.

H:1922:BRBa

[220] C. K. H. Book review: Problèmes Actuels de l’Economique. Journal of the Royal Statistical Society, 85(3):496–499, May 1922. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341190.

H:1922:BRBb

[221] C. K. H. Book review: Reparations, Trade, and Foreign Exchange, by L. L.

B. Angas. Journal of the Royal Statistical Society, 85(3):499–500, May 1922. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341191.

B:1922:BRBc

[222] A. L. B. Book review: Wages, Prices and Profits: A Report Prepared, by the Labour Research Department by Sidney Webb. Journal of the Royal Statistical Society, 85(3):500–502, May 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341192.

B:1922:BRBd

[223] A. L. B. Book review: The Consumers’ Co-Operative Movement, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 85(3): 502–505, May 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341193.

C:1922:BRBc

[224] C. E. C. Book review: Consumers’ Co-Operative Societies, by Charles Gide; Staff of the Co-Operative Reference Library. Journal of the Royal Statistical Society, 85(3):505–507, May 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341194.

B:1922:BRBe

[225] A. L. B. Book review: Field Work and Social Research, by F. Stuart Chapin. Journal of the Royal Statistical Society, 85(3):507–508, May 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341195.

G:1922:BRBb

[226] J. B. G. Book review: The United States Department of Agriculture: A Study in Administration, by William L. Wanlass. Journal of the Royal Statistical Society, 85(3):508–509, May 1922. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341196.

G:1922:BRBc

[227] J. B. G. Book review: Esquisse d’une Histoire du Régime Agraire en Europe aux XVIIIe et XIXe Siècles, by Henri Sée. Journal of the Royal Statistical Society, 85(3):509–510, May 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341197.

Anonymous:1922:ONPc

[228] Anonymous. Other new publications. Journal of the Royal Statistical Society, 85(3):510–513, May 1922. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341198.

Anonymous:1922:CNc

[229] Anonymous. Current notes. Journal of the Royal Statistical Society, 85(3): 514–525, May 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341199.

Anonymous:1922:SEAc

[230] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 85(3): 526–529, May 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341200.

Anonymous:1922:LALc

[231] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 85(3):530–535, May 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341201.

Anonymous:1922:VIb

[232] Anonymous. Volume information. Journal of the Royal Statistical Society, 85 (4):681–688, June 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341121.

Greenwood:1922:DVL

[233] M. Greenwood. Discussion on the value of life-tables in statistical research. Journal of the Royal Statistical Society, 85 (4):537–560, June 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341122.

Verdier:1922:ISR

[234] J. W. Verdier. The interpretation of statistics relating to shipping casualties and loss of life at sea. Journal of the Royal Statistical Society, 85(4):561–596, June 1922. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341123.

Fisher:1922:GFR

[235] R. A. Fisher. The goodness of fit of regression formulae, and the distribution of regression coefficients. Journal of the Royal Statistical Society, 85(4):597–612, June 1922. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341124.

Anonymous:1922:RCF

[236] Anonymous. Report of the Council for the financial year ended December 31, 1921, and for the sessional year ending June 20, 1922, presented at the eightyeighth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June, 20, 1922. Journal of the Royal Statistical Society, 85(4):613–623, June 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341125.

Anonymous:1922:PEE

[237] Anonymous. Proceedings of the eightyeighth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 20, 1922. Journal of the Royal Statistical Society, 85(4):624–625, June 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341126.

Y:1922:BRB

[238] G. U. Y. Book review: The Population Problem: A Study in Human Evolution, by A. M. Carr-Saunders. Journal of the Royal Statistical Society, 85(4):626–629, June 1922. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341127.

G:1922:BRBd

[239] M. G. Book review: Demografia di Guerra e Altri Saggi, by Franco Savorgnan. Journal of the Royal Statistical Society, 85(4):629–630, June 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341128.

H:1922:BRBc

[240] C. K. H. Book review: British Finance During and After the War, 1914–21, by A. H. Gibson; A. W. Kirkaldy. Journal of the Royal Statistical Society, 85(4): 630–631, June 1922. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341129.

H:1922:BRBd

[241] C. K. H. Book review: British Labour: Replacement and Conciliation, 1914–21, by A. W. Kirkaldy. Journal of the Royal Statistical Society, 85(4):631–632, June 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341130.

Anonymous:1922:BRB

[242] Anonymous. Book review: Le Système Monétaire et le Change Anglais Depuis la Guerre, by Alfred Mawas. Journal of the Royal Statistical Society, 85(4):632, June 1922. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341131.

B:1922:BRBf

[243] J. A. B. Book review: Some Aspects of the Economic Consequences of the War from India, by S. G. Panandikar. Journal of the Royal Statistical Society, 85(4):633–634, June 1922. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341132.

S:1922:BRBg

[244] J. C. S. Book review: Income in the United States, 1909–19, by Wesley C. Mitchell; Willford J. King; Frederick R. Macaulay; Oswald W. Knauth. Journal of the Royal Statistical Society, 85(4): 634–635, June 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341133.

B:1922:BRBg

[245] J. B. Book review: Les Crises Economiques; Essai de Morphologie et Theorie des Crises Economiques Periodiques et de Theorie de la Conjoncture Economique, by M. Bouniatian; J. Bernard. Journal of the Royal Statistical Society, 85(4):635–636, June 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341134.

S:1922:BRBh

[246] E. C. S. Book review: The Trade Cycle, by F. Lavington. Journal of the Royal Statistical Society, 85(4):636–638, June 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341135.

P:1922:BRB

[247] W. P. Book review: Supply and Demand, by Hubert D. Henderson; Money, by D. H. Robertson. Journal of the Royal Statistical Society, 85(4):638–640, June 1922. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341136.

H:1922:BRBe

[248] C. K. H. Book review: Why Prices Rise and Fall, by F. W. Pethick Lawrence; Unemployment, by F. W. Pethick Lawrence. Journal of the Royal Statistical Society, 85(4):640–641, June 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341137.

Sp:1922:BRB

[249] J. C. Sp. Book review: The Local Government of the United Kingdom, by J. J. Clarke. Journal of the Royal Statistical Society, 85(4):641–642, June 1922. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341138.

G:1922:BRBe

[250] J. B. G. Book review: The Story of the Agricultural Club, by R. Henry Rew; Lord Bledisloe. Journal of the Royal Statistical Society, 85(4):642–643, June 1922. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341139.

Anonymous:1922:ONPd

[251] Anonymous. Other new publications. Journal of the Royal Statistical Society, 85(4):643–648, June 1922. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341140.

Anonymous:1922:CNd

[252] Anonymous. Current notes. Journal of the Royal Statistical Society, 85(4): 649–661, June 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341141.

N:1922:DJB

[253] A. N. Dr. Jacques Bertillon. Journal of the Royal Statistical Society, 85(4): 662–664, June 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341142.

Anonymous:1922:SEAd

[254] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 85(4): 665–668, June 1922. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341143.

Anonymous:1922:LALd

[255] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 85(4):669–673, June 1922. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341144.

Anonymous:1922:ALA

[256] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 85(4):674–680, June 1922. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341145.

Anonymous:1923:VIa

[257] Anonymous. Volume information. Journal of the Royal Statistical Society, 86 (1):i–vii, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341742.

Emmott:1923:RCL

[258] Lord Emmott. The relations of capital and labour. Journal of the Royal Statistical Society, 86(1):1–21, January 1923. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341743.

Anonymous:1923:PM

[259] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 86(1):22–25, January 1923. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341744.

deJastrzebski:1923:CBR

[260] T. T. S. de Jastrzebski. Changes in the birth-rate and in legitimate fertility in London, 1911–1921. Journal of the Royal Statistical Society, 86(1):26– 58, January 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341745.

Edgeworth:1923:SE

[261] F. Y. Edgeworth. Statistics of examinations. Journal of the Royal Statistical Society, 86(1):59–60, January 1923. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341746.

Newbold:1923:PMT

[262] E. M. Newbold. Professor S. Minetola’s theory of multiple frequency. Journal of the Royal Statistical Society, 86(1): 61–62, January 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341747.

Anonymous:1923:PGP

[263] Anonymous. The price of government publications. Journal of the Royal Statistical Society, 86(1):63–64, January 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341748.

E:1923:BRBa

[264] F. Y. E. Book review: Das Gesetz der Grossen Zahlen und der Stochastisch-Statistische Standpunkt in der Modernen Wissenschaft, by Al. A. Tschuprow. Journal of the Royal Statistical Society, 86(1):65–66, January 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341749.

I:1923:BRBa

[265] L. I. Book review: Frequency Arrays, by H. E. Soper. Journal of the Royal Statistical Society, 86(1):67, January 1923. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341750.

T:1923:BRBa

[266] H. L. T. Book review: Lezioni di Statistica Economica e Demografica dettate nel R. Instituto Superiore di Studi Commerciali in Roma, 1919–20, by Giorgio Mortara. Journal of the Royal Statistical Society, 86(1):67–69, January 1923. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341751.

H:1923:BRBa

[267] C. K. H. Book review: Gold und Papiergeld, by Dr. Lassar-Cohn. Journal of the Royal Statistical Society, 86(1): 69–70, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341752.

C:1923:BRBa

[268] N. E. C. Book review: Tariffs: A Study in Method, by T. E. G. Gregory. Journal of the Royal Statistical Society, 86(1): 70–72, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341753.

W:1923:BRB

[269] A. D. W. Book review: A Short Fiscal and Financial History of England, 1815–1918, by J. F. Rees. Journal of the Royal Statistical Society, 86(1):73, January 1923. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341754.

M:1923:BRBa

[270] H. W. M. Book review: The Cotton Control Board, by Hubert D. Henderson. Journal of the Royal Statistical Society, 86(1):73–75, January 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341755.

H:1923:BRBb

[271] C. K. H. Book review: Coal in International Trade, by A. J. Sargent. Journal of the Royal Statistical Society, 86 (1):76, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341756.

B:1923:BRBa

[272] A. Les. B. Book review: Costs, Merchandising Practices, Advertising and Sales in the Retail Distribution of Clothing, by North Western University School of Commerce. Journal of the Royal Statistical Society, 86(1):76–78, January 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341757.

G:1923:BRBa

[273] J. B. G. Book review: The Marketing of Whole Milk, by Henry E. Erdman. Journal of the Royal Statistical Society, 86(1):78–79, January 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341758.

C:1923:BRBb

[274] C. E. C. Book review: History of Labour in the United States, by John R. Commons; David J. Saposs; Helen L. Sumner; E. B. Mittelman; H. E. Hoagland; John B. Andrews; Selig Perlman. Journal of the Royal Statistical Society, 86 (1):79–82, January 1923. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341759.

N:1923:BRBa

[275] A. N. Book review: After Two Years:

A Study of American Prohibition, by Philip Whitwell Wilson. Journal of the Royal Statistical Society, 86(1):82– 84, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341760.

Anonymous:1923:ONPa

[276] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 86(1):84–87, January 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341761.

Anonymous:1923:CNa

[277] Anonymous. Current notes. Journal of the Royal Statistical Society, 86(1): 88–102, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341762.

Anonymous:1923:SEAa

[278] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 86(1): 103–109, January 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341763.

Anonymous:1923:LALa

[279] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 86(1):110–121, January 1923. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341764.

Anonymous:1923:PRa

[280] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 86(1): 122–128, January 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341765.

Anonymous:1923:PNA

[281] Anonymous. [photograph]: Noel Algernon Humphreys, I.S.O., Honorary Secretary from 1896 to 1905. Journal of the Royal Statistical Society, 86 (2):??, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341391.

Dudfield:1923:DRD

[282] R. Dudfield. Discussion on the registration of disease. Journal of the Royal Statistical Society, 86(2):129–153, March 1923. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341392.

Hilton:1923:SUD

[283] John Hilton. Statistics of unemployment derived from the working of the Unemployment Insurance Acts. Journal of the Royal Statistical Society, 86(2): 154–205, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341393.

ES:1923:WPC

[284] Editor of the “Statist”. Wholesale prices of commodities in 1922. Journal of the Royal Statistical Society, 86(2):206– 225, March 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341394.

Neale:1923:NZS

[285] E. P. Neale. A New Zealand study in seasonal fluctuations of external migration, with special reference to the computation of mean annual populations. Journal of the Royal Statistical Society, 86 (2):226–241, March 1923. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341395.

E:1923:BRBb

[286] W. P. E. Book review: Life Contingencies, by E. F. Spurgeon; Calculus and Probability for Actuarial Students. by Alfred Henry. Journal of the Royal Statistical Society, 86(2):242–247, March 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341396.

H:1923:BRBc

[287] R. G. H. Book review: Money and Foreign Exchange after 1914, by G. Cassel. Journal of the Royal Statistical Society, 86(2):247–248, March 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341397. See reply [482].

H:1923:BRBd

[288] R. G. H. Book review: La Monnaie, le Credit et le Change, by Auguste Arnaune. Journal of the Royal Statistical Society, 86(2):249, March 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341398.

H:1923:BRBe

[289] R. G. H. Book review: The London Money Market, by W. F. Spalding. Journal of the Royal Statistical Society, 86 (2):249–250, March 1923. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341399.

T:1923:BRBb

[290] C. T. Book review: Principles of Public Finance, by Hugh Dalton; Public Finance. Journal of the Royal Statistical Society, 86(2):251, March 1923. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341400.

H:1923:BRBf

[291] J. H. Book review: Trusts in British Industry: 1914–21. A Study of Recent Developments in Business Organization, by J. Morgan Rees. Journal of the Royal Statistical Society, 86(2): 251–254, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341401.

P:1923:BRB

[292] W. P. Book review: Labour Policy–False and True: A Study in Economic History and Industrial Economics, by Lynden Macassey. Journal of the Royal Statistical Society, 86(2):255–257, March 1923. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341402.

H:1923:BRBg

[293] J. H. Book review: Sharing Profits with Employees, by James A. Bowie. Journal of the Royal Statistical Society, 86(2): 257–260, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341403.

R:1923:BRBa

[294] C. S. R. Book review: English Prisons under Local Government, by Sidney Webb; Beatrice Webb; Bernard Shaw; English Prisons of To-day. being the Report of the Prison System Enquiry Committee, by Stephen Hobhouse; A. Fenner Brockway. Journal of the Royal Statistical Society, 86(2):260–263, March 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341404.

G:1923:BRBb

[295] M. G. Book review: Social Administration, Including the Poor Laws, by John J. Clarke. Journal of the Royal Statistical Society, 86(2):263, March 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341405.

Anonymous:1923:ONPb

[296] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 86(2):264–266, March 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341406.

Anonymous:1923:CNb

[297] Anonymous. Current notes. Journal of the Royal Statistical Society, 86(2): 267–276, March 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341407.

deJ:1923:NAH

[298] S. de J. Noel Algernon Humphreys.

Journal of the Royal Statistical Society, 86(2):277, March 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341408.

K:1923:SEC

[299] A. W. W. K. Sir Ernest Clarke. Journal of the Royal Statistical Society, 86(2): 277–279, March 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341409.

P:1923:SCS

[300] E. C. P. Sir Charles Stewart Loch. Journal of the Royal Statistical Society, 86 (2):279–280, March 1923. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341410.

Westergaard:1923:MR

[301] Harald Westergaard. Marcus Rubin. Journal of the Royal Statistical Society, 86(2):280–281, March 1923. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341411.

Anonymous:1923:SEAb

[302] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 86(2): 282–286, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341412.

Anonymous:1923:LALb

[303] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 86(2):287–293, March 1923. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341413.

Anonymous:1923:PRb

[304] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 86(2): 294–295, March 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341414.

Flux:1923:ISC

[305] A. W. Flux. International statistical comparisons. Journal of the Royal Statistical Society, 86(3):297–331, May 1923. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341620.

Snow:1923:TFP

[306] E. C. Snow. Trade forecasting and prices. Journal of the Royal Statistical Society, 86(3):332–398, May 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341621.

Wolfenden:1923:MCM

[307] Hugh H. Wolfenden. On the methods of comparing the moralities of two or more communities, and the standardization of death-rates. Journal of the Royal Statistical Society, 86(3):399–411, May 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341622.

Stead:1923:CCC

[308] H. G. Stead. The correction of correlation coefficients. Journal of the Royal Statistical Society, 86(3):412–419, May 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341623.

Hart:1923:NRS

[309] J. Y. Hart. A note on the rates of sickness of elementary school teachers recruited from London and the provinces. Journal of the Royal Statistical Society, 86(3):420–422, May 1923. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341624.

G:1923:BRBc

[310] M. G. Book review: Statistik, by Sigmund Schott. Journal of the Royal Statistical Society, 86(3):423–424, May 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341625.

Y:1923:BRB

[311] G. U. Y. Book review: The Making of Index-Numbers, by Irving Fisher. Journal of the Royal Statistical Society, 86 (3):424–430, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341626.

B:1923:BRBb

[312] J. B. Book review: Money, Credit and Commerce, by Alfred Marshall. Journal of the Royal Statistical Society, 86 (3):430–433, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341627.

C:1923:BRBc

[313] E. C. Book review: Monetary Reconstruction, by R. G. Hawtrey. Journal of the Royal Statistical Society, 86(3): 433–436, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341628.

B:1923:BRBc

[314] A. L. B. Book review: Report on an Enquiry into the Wages and Hours of Labour in the Cotton Mill Industry, by G. Findlay Shirras. Journal of the Royal Statistical Society, 86(3):436–437, May 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341629.

V:1923:BRBa

[315] J. W. V. Book review: The British Coal Mining Industry during the War, by R. A. S. Redmayne. Journal of the Royal Statistical Society, 86(3):437–439, May 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341630.

V:1923:BRBb

[316] J. W. V. Book review: Wages in the Coal Industry, by J. W. F. Rowe. Journal of the Royal Statistical Society, 86 (3):439–440, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341631.

R:1923:BRBb

[317] R. H. R. Book review: Food Production in War, by T. H. Middleton. Journal of the Royal Statistical Society, 86 (3):440–442, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341632.

G:1923:BRBd

[318] J. B. G. Book review: Supplying Britain’s Meat, by George E. Putnam. Journal of the Royal Statistical Society, 86(3):443, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341633.

M:1923:BRBb

[319] H. W. M. Book review: The Economics of Unemployment, by J. A. Hobson. Journal of the Royal Statistical Society, 86(3):443–444, May 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341634.

N:1923:BRBb

[320] E. M. N. Book review: Mental and Scholastic Tests, by Cyril Burt. Journal of the Royal Statistical Society, 86 (3):444–446, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341635.

N:1923:BRBc

[321] E. M. N. Book review: A Study of American Intelligence, by Carl C. Brigham. Journal of the Royal Statistical Society, 86(3):446, May 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341636.

Anonymous:1923:ONPc

[322] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 86(3):447–452, May 1923. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341637.

Anonymous:1923:CNc

[323] Anonymous. Current notes. Journal of the Royal Statistical Society, 86(3): 453–461, May 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341638.

Anonymous:1923:SM

[324] Anonymous. Shirley Murphy. Journal of the Royal Statistical Society, 86(3): 462–463, May 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341639.

Anonymous:1923:SEAc

[325] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 86(3): 464–467, May 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341640.

Anonymous:1923:LALc

[326] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 86(3):468–474, May 1923. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341641.

Anonymous:1923:PRc

[327] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 86 (3):475–482, May 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341642.

Anonymous:1923:VIb

[328] Anonymous. Volume information. Journal of the Royal Statistical Society, 86 (4):670–678, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341692.

Wilson:1923:SRC

[329] D. R. Wilson. On some recent contributions to the study of industrial fatigue. Journal of the Royal Statistical Society, 86(4):483–515, July 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341693.

Bowley:1923:DRD

[330] A. L. Bowley. Death-rates, density, population, and housing. Journal of the Royal Statistical Society, 86(4):516–546, July 1923. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341694.

Dunlop:1923:MAR

[331] James Craufurd Dunlop. Misstatement of age in the returns of the census of Scotland. Journal of the Royal Statistical Society, 86(4):547–569, July 1923. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341695.

Edgeworth:1923:MCW

[332] F. Y. Edgeworth. Mr. Correa Walsh on the calculation of index-numbers. Journal of the Royal Statistical Society, 86 (4):570–590, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341696.

Dudfield:1923:NCL

[333] R. Dudfield. Note on the census of London (1921). Journal of the Royal Statistical Society, 86(4):591–593, July 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341697.

Anonymous:1923:RCF

[334] Anonymous. Report of the Council for the financial year ended December 31, 1922, and for the sessional year ending June 19, 1923, presented at the eightyninth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 19 1923. Journal of the Royal Statistical Society, 86 (4):594–608, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341698.

Anonymous:1923:PEN

[335] Anonymous. Proceedings of the eightyninth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 19, 1923. Journal of the Royal Statistical Society, 86(4):609–610, July 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341699.

I:1923:BRBb

[336] L. I. Book review: Some Recent Tracts on the Mathematical Theory of Statistics. (1) On Correlation, by Alf. Guldberg; Uber die Bedeutung des Bravais-Pearsonschen Korrelations-koeffizienten. by Ernst Mally. Journal of the Royal Statistical Society, 86(4):611–612, July 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341700.

G:1923:BRBe

[337] M. G. Book review: Corso di Matematica Finanziaria, by F. Insolera. Journal of the Royal Statistical Society, 86 (4):612–613, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341701.

S:1923:BRBa

[338] J. C. S. Book review: Income in the United States — Its Amount and Distribution, 1909–19, by National Bureau of Economic Research; Willford I. King; Oswald W. Knauth; Frederick R. Macaulay; Wesley C. Mitchell; Wealth and Income of the American People — A Survey of the Economic Consequences of the War. by Walter R. Ingalls. Journal of the Royal Statistical Society, 86 (4):613–619, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341702.

M:1923:BRBc

[339] H. W. M. Book review: Money, Banking, and Exchange in India, by H. Stanley Jevon; The Future of Exchange and the Indian Currency. by H. Stanley Jevons. Journal of the Royal Statistical Society, 86(4):620–621, July 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341703.

H:1923:BRBh

[340] R. G. H. Book review: Restoration of the World’s Currencies, by R. A. Lehfeldt. Journal of the Royal Statistical Society, 86(4):621–622, July 1923. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341704.

S:1923:BRBb

[341] E. C. S. Book review: An Academic Study of Some Money Markets and Other Statistics, by E. G. Peake. Journal of the Royal Statistical Society, 86(4):622–624, July 1923. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341705.

H:1923:BRBi

[342] C. K. H. Book review: Monetary and Banking Policy of Chile, by Guillermo Subercaseaux; David Kinley. Journal of the Royal Statistical Society, 86(4): 624–626, July 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341706.

S:1923:BRBc

[343] E. C. S. Book review: Business Cycles and Unemployment, by National Bureau of Economic Research. Journal of the Royal Statistical Society, 86(4):626–628, July 1923. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341707.

H:1923:BRBj

[344] C. K. H. Book review: Effects of the War upon French Economic Life, by Charles Gide. Journal of the Royal Statistical Society, 86(4):628–629, July 1923. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341708.

H:1923:BRBk

[345] C. K. H. Book review: La Politique du Charbon, 1914–21, by M. Olivier. Journal of the Royal Statistical Society, 86 (4):629–630, July 1923. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341709.

H:1923:BRBl

[346] H. H. Book review: Report on an Inquiry into Working Class Budgets in Bombay, by G. Findlay Shirras. Journal of the Royal Statistical Society, 86(4):630–632, July 1923. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341710.

Anonymous:1923:ONPd

[347] Anonymous. Other new publications. Journal of the Royal Statistical Society, 86(4):633–638, July 1923. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341711.

Anonymous:1923:CNd

[348] Anonymous. Current notes. Journal of the Royal Statistical Society, 86(4): 639–651, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341712.

Anonymous:1923:SEAd

[349] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 86(4): 652–655, July 1923. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341713.

Anonymous:1923:LALd

[350] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 86(4):656–661, July 1923. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341714.

Anonymous:1923:ALA

[351] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 86(4):662–669, July 1923. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341715.

Anonymous:1924:VIa

[352] Anonymous. Volume information. Journal of the Royal Statistical Society, 87 (1):i–vii, January 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341287.

Baines:1924:ISI

[353] J. A. Baines. The International Statistical Institute and its fifteenth session. Journal of the Royal Statistical Society, 87(1):1–21, January 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341288.

Faber:1924:APD

[354] Harald Faber. Agricultural production in Denmark. Journal of the Royal Statistical Society, 87(1):22–75, January 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341289.

Brownlee:1924:SET

[355] John Brownlee. Some experiments to test the theory of goodness of fit. Journal of the Royal Statistical Society, 87 (1):76–82, January 1924. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341290.

Brownlee:1924:TPA

[356] John Brownlee. A test of periodogram analysis. Journal of the Royal Statistical Society, 87(1):83–88, January 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341291.

Fisher:1924:MUY

[357] Irving Fisher. Mr. Udny Yule on indexnumbers. Journal of the Royal Statistical Society, 87(1):89–98, January 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341292.

Coates:1924:LNR

[358] W. H. Coates. League of Nations report on double taxation submitted to the Financial Committee by Professors Bruins, Einaudi, Seligman, and Sir Josiah Stamp. Journal of the Royal Statistical Society, 87(1):99–102, January 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341293.

E:1924:BRBa

[359] F. Y. E. Book review: Die Statistischen Forschungsmethoden, by Emanuel Czuber. Journal of the Royal Statistical Society, 87(1):103–106, January 1924. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341294.

F:1924:BRBa

[360] A. W. F. Book review: Principes de Statistique Theorique et Appliquee, by Armand Julin. Journal of the Royal Statistical Society, 87(1):106–107, January 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341295.

M:1924:BRBa

[361] F. A. A. M. Book review: Actuarial Studies. No. 1: Sources and Characteristics of the Principal Mortality Tables. No. 2: Construction of Mortality Tables from the Records of Insured Lives. No. 4: Graduation of Mortality and Other Tables. No. 5: Total Permanent Disability Benefits in Relation to Life Insurance. Journal of the Royal Statistical Society, 87(1):108–110, January 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341296.

E:1924:BRBb

[362] F. Y. E. Book review: Economic Essays, by David Ricardo; E. C. K. Gonner. Journal of the Royal Statistical Society, 87(1):110–112, January 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341297.

C:1924:BRBa

[363] E. C. Book review: Economic Principles: An Introductory Study, by A. W. Flux. Journal of the Royal Statistical Society, 87(1):112–113, January 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341298.

C:1924:BRBb

[364] W. H. C. Book review: A Study in National Finance, by T. J. Kiernan. Journal of the Royal Statistical Society, 87 (1):113–114, January 1924. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341299.

B:1924:BRBa

[365] A. Les. B. Book review: Selling Expenses and Their Control, by H. Secrist. Journal of the Royal Statistical Society, 87(1):115–116, January 1924. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341300.

B:1924:BRBb

[366] A. Les. B. Book review: A Business Barometer for Retailers, by H. Secrist. Journal of the Royal Statistical Society, 87(1):116–117, January 1924. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341301.

T:1924:BRBa

[367] C. T. Book review: Cycles of Unemployment in the United States, 1903–22, by William A. Berridge. Journal of the Royal Statistical Society, 87(1):118, January 1924. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341302.

V:1924:BRB

[368] H. V. Book review: The Workers’ Register of Labour and Capital, 1923, by Labour Research Department. Journal of the Royal Statistical Society, 87(1): 119, January 1924. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341303.

N:1924:BRB

[369] E. M. N. Book review: Mental and Scholastic Tests among Retarded Children. Journal of the Royal Statistical Society, 87(1):119–120, January 1924. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341304.

Anonymous:1924:ONPa

[370] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 87(1):120–123, January 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341305.

Anonymous:1924:CNa

[371] Anonymous. Current notes. Journal of the Royal Statistical Society, 87(1): 124–135, January 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341306.

Anonymous:1924:SEAa

[372] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 87(1): 136–143, January 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341307.

Anonymous:1924:LALa

[373] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 87(1):144–158, January 1924. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341308.

Anonymous:1924:PRa

[374] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 87(1): 159–165, January 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341309.

Crump:1924:IDP

[375] Norman Crump. The interrelation and distribution of prices and their incidence upon price stabilization. Journal of the Royal Statistical Society, 87(2): 167–219, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341225.

Sanders:1924:FTS

[376] R. Y. Sanders. Foreign trade and shipbuilding. Journal of the Royal Statistical Society, 87(2):220–255, March 1924. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341226.

Anonymous:1924:WPC

[377] Anonymous. Wholesale prices of commodities in 1923. Journal of the Royal Statistical Society, 87(2):256– 277, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341227.

Trachtenberg:1924:WAG

[378] H. L. Trachtenberg. The wider application of the Gompertz law of mortality. Journal of the Royal Statistical Society, 87(2):278–290, March 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341228.

G:1924:BRBa

[379] M. G. Book review: The Calculus of Observations: A Treatise on Numerical Mathematics, by E. T. Whittaker; G. Robinson; A Short Course in Interpolation. by E. T. Whittaker; G. Robinson. Journal of the Royal Statistical Society, 87(2):291–293, March 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341229.

Br:1924:BRB

[380] J. Br. Book review: Statistical Method, by Truman L. Kelley. Journal of the Royal Statistical Society, 87(2):293– 294, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341230.

G:1924:BRBb

[381] M. G. Book review: Il Metodo Statistico, Teoria e Applicazioni alle Scienze naturali, alle Scienze sociali, e all’ Arte, by Alfredo Niceforo. Journal of the Royal Statistical Society, 87(2): 294–296, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341231.

R:1924:BRBa

[382] R. H. R. Book review: International Year-Book of Agricultural Statistics, 1922. Journal of the Royal Statistical Society, 87(2):296–297, March 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341232.

R:1924:BRBb

[383] R. H. R. Book review: Foundations of Agricultural Economics, by J. A. Venn. Journal of the Royal Statistical Society, 87(2):297–299, March 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341233.

R:1924:BRBc

[384] R. H. R. Book review: Mankind at the Crossroads, by Edward M. East. Journal of the Royal Statistical Society, 87(2): 299–300, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341234.

B:1924:BRBc

[385] J. B. Book review: Histoire des Doctrines de la Population, by Rene Gonnard. Journal of the Royal Statistical Society, 87(2):301–302, March 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341235.

H:1924:BRBa

[386] R. G. H. Book review: Costs and Profits: Their Relation to Business Cycles, by Hudson B. Hastings. Journal of the Royal Statistical Society, 87(2): 302–303, March 1924. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341236.

H:1924:BRBb

[387] R. G. H. Book review: Bankers and Credit, by Hartley Withers. Journal of the Royal Statistical Society, 87(2): 303–307, March 1924. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341237.

M:1924:BRBb

[388] B. M. Book review: Political Economy Club, 1821–1920. Centenary Volume. Journal of the Royal Statistical Society, 87(2):307–308, March 1924. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341238.

T:1924:BRBb

[389] C. T. Book review: Socialisation in Theory and Practice, by Heinrich Strobel; H. J. Stenning. Journal of the Royal Statistical Society, 87(2):308, March 1924. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341239.

Anonymous:1924:ONPb

[390] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 87(2):309–313, March 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341240.

Anonymous:1924:CNb

[391] Anonymous. Current notes. Journal of the Royal Statistical Society, 87(2): 314–325, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341241.

Anonymous:1924:SEAb

[392] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 87(2): 326–333, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341242.

Anonymous:1924:LALb

[393] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 87(2):334–339, March 1924. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341243.

Anonymous:1924:PRb

[394] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 87(2): 340–349, March 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341244.

Flux:1924:CP

[395] A. W. Flux. The census of production. Journal of the Royal Statistical Society, 87(3):351–390, May 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341146.

Lowes-Dickinson:1924:PIA

[396] Arthur Lowes-Dickinson. Publicity in industrial accounts, with a comparison of English and American methods. Journal of the Royal Statistical Society, 87 (3):391–432, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341147.

VanIsselt:1924:CIP

[397] E. W. Van Dam Van Isselt. The census of industrial production in the Netherlands. Journal of the Royal Statistical Society, 87(3):433–441, May 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341148.

Fisher:1924:CUW

[398] R. A. Fisher. The conditions under which χ^{2} measures the discrepancey between observation and hypothesis. Journal of the Royal Statistical Society, 87 (3):442–450, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341149.

G:1924:BRBc

[399] M. G. Book review: Statistica Metodologica, by Felice Vinici. Journal of the Royal Statistical Society, 87(3): 451–452, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341150.

G:1924:BRBd

[400] M. G. Book review: Mathematische Bevölkerungstheorie, by E. Czuber. Journal of the Royal Statistical Society, 87(3):452, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341151.

S:1924:BRBa

[401] J. C. S. Book review: A Tract on Monetary Reform, by John Maynard Keynes. Journal of the Royal Statistical Society, 87(3):452–455, May 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341152.

F:1924:BRBb

[402] A. W. F. Book review: The Inversion of Science and a Scheme of Scientific Reformation, by Frederick Soddy. Journal of the Royal Statistical Society, 87(3): 455–456, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341153.

S:1924:BRBb

[403] J. C. S. Book review: Current Economic Affairs, by Walter Renton Ingalls. Journal of the Royal Statistical Society, 87 (3):456–458, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341154.

C:1924:BRBc

[404] W. H. C. Book review: Land-Value Policy, by J. Dundas White. Journal of the Royal Statistical Society, 87(3):458–460, May 1924. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341155.

A:1924:BRB

[405] W. M. A. Book review: Railway Amalgamation in Great Britain, by W. R. Simnett. Journal of the Royal Statistical Society, 87(3):460–462, May 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341156.

C:1924:BRBd

[406] C. E. C. Book review: The British Trade Boards System, by Dorothy Sells. Journal of the Royal Statistical Society, 87 (3):462–463, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341157.

H:1924:BRBc

[407] C. K. H. Book review: Employment, Hours and Earnings in Prosperity and Depression: United States, 1920–22, by Willford I. King. Journal of the Royal Statistical Society, 87(3):463–465, May 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341158.

H:1924:BRBd

[408] C. K. H. Book review: Losses of Life Caused, by War by Samuel Dumas; K. O. Vedel-Petersen; Harald Westergaard. Journal of the Royal Statistical Society, 87(3):465–466, May 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341159.

Anonymous:1924:ONPc

[409] Anonymous. Other new publications. Journal of the Royal Statistical Society, 87(3):466–469, May 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341160.

Anonymous:1924:CNc

[410] Anonymous. Current notes. Journal of the Royal Statistical Society, 87(3): 470–480, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341161.

Anonymous:1924:SEAc

[411] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 87(3): 481–484, May 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341162.

Anonymous:1924:LALc

[412] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 87(3):485–492, May 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341163.

Anonymous:1924:VIb

[413] Anonymous. Volume information. Journal of the Royal Statistical Society, 87 (4):666–672, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341459.

Greenwood:1924:VSS

[414] Major Greenwood. The vital statistics of Sweden and England and Wales: An essay in international comparison. Journal of the Royal Statistical Society, 87 (4):493–543, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341460.

Hilton:1924:ESE

[415] John Hilton. Enquiry by sample: An experiment and its results. Journal of the Royal Statistical Society, 87(4):544–570, July 1924. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341461.

Edgeworth:1924:UMR

[416] F. Y. Edgeworth. Untried methods of representing frequency. Journal of the Royal Statistical Society, 87(4):571–594, July 1924. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341462.

Baines:1924:CI

[417] J. A. Baines. The census of India, 1921.

Journal of the Royal Statistical Society, 87(4):595–598, July 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341463.

Anonymous:1924:RCF

[418] Anonymous. Report of the Council for the financial year ended December 31, 1923, and for the sessional year ending June 17, 1924, presented at the ninetienth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 17, 1924. Journal of the Royal Statistical Society, 87(4):599–613, July 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341464.

Anonymous:1924:PNA

[419] Anonymous. Proceedings of the ninetieth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 17, 192. Journal of the Royal Statistical Society, 87(4):614–616, July 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341465.

I:1924:BRB

[420] L. I. Book review: The Mathematical Theory of Probabilities, by Arne Fisher. Journal of the Royal Statistical Society, 87(4):617–618, July 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341466.

D:1924:BRB

[421] R. D. Book review: The Elements of Vital Statistics in their Bearing on Social and Public Health Problems, by A. Newsholme. Journal of the Royal Statistical Society, 87(4):618–619, July 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341467.

F:1924:BRBc

[422] A. W. F. Book review: The Mathematical Groundwork of Economics, by A. L. Bowley. Journal of the Royal Statistical Society, 87(4):619–621, July 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341468.

S:1924:BRBc

[423] J. C. S. Book review: The Foreign Exchanges, by A. W. Flux. Journal of the Royal Statistical Society, 87(4):621–622, July 1924. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341469.

H:1924:BRBe

[424] C. K. H. Book review: War Finances of the Netherlands up to 1918, by M. J. van der Flier. Journal of the Royal Statistical Society, 87(4):622–623, July 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341470.

T:1924:BRBc

[425] C. T. Book review: The Inter-Ally Debts: An Analysis of War and Post-War Public Finance, 1914–1923, by Harvey E. Fisk. Journal of the Royal Statistical Society, 87(4):623–624, July 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341471.

T:1924:BRBd

[426] R. J. T. Book review: The Agricultural Crisis, 1920–23, by R. R. Enfield. Journal of the Royal Statistical Society, 87 (4):624–625, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341472.

B:1924:BRBd

[427] W. A. B. Book review: Labour Supply and Regulation, by Humbert Wolfe.

Journal of the Royal Statistical Society, 87(4):625–626, July 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341473.

B:1924:BRBe

[428] W. A. B. Book review: The Twelve-Hour Shift in Industry, by Committee of Work-Periods in Continuous Industry of the Federated American Engineering Societies. Journal of the Royal Statistical Society, 87(4):626–627, July 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341474.

T:1924:BRBe

[429] C. T. Book review: Economics of the Household, by Benjamin Andrews. Journal of the Royal Statistical Society, 87 (4):628, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341475.

Anonymous:1924:ONPd

[430] Anonymous. Other new publications. Journal of the Royal Statistical Society, 87(4):628–632, July 1924. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341476.

Anonymous:1924:CNd

[431] Anonymous. Current notes. Journal of the Royal Statistical Society, 87(4): 633–646, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341477.

Anonymous:1924:SEAd

[432] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 87(4): 647–650, July 1924. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341478.

Anonymous:1924:LALd

[433] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 87(4):651–657, July 1924. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341479.

Anonymous:1924:ALA

[434] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 87(4):658–665, July 1924. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341480.

Anonymous:1925:VIa

[435] Anonymous. Volume information. Journal of the Royal Statistical Society, 88 (1):i–vii, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341574.

Anonymous:1925:GPF

[436] Anonymous. The growth of population and the factors which control it. Journal of the Royal Statistical Society, 88 (1):1–58, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341575.

Anonymous:1925:PM

[437] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 88(1):58–62, January 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341576.

Stevenson:1925:LGP

[438] T. H. C. Stevenson. The laws governing population. Journal of the Royal Statistical Society, 88(1):63–90, January 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341577.

Tschuprow:1925:AFD

[439] Al. A. Tschuprow. On the asymptotic frequency distribution of the arithmetic means of n correlated observations for very great values of n. Journal of the Royal Statistical Society, 88(1):91– 104, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341578.

Brownlee:1925:ECD

[440] John Brownlee. On the error in the correlation due to random sampling when proportionate mortalities are used. Journal of the Royal Statistical Society, 88(1):105–106, January 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341579.

D:1925:BRB

[441] R. D. Book review: The Function of Statistical Method in Scientific Investigation, by G. Udny Yule. Journal of the Royal Statistical Society, 88(1):107– 108, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341580.

N:1925:BRBa

[442] E. M. N. Book review: Statistical Method, by Harry Jerome. Journal of the Royal Statistical Society, 88(1):108, January 1925. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341581.

E:1925:BRBa

[443] F. Y. E. Book review: Matematica de la Mortalidad con Elementos de Probabilades, by Isidoro Rubio; Introduccion a la Metodologia Estadistica (Fundamentos de Estadistica Matematica), by Antonio Flores de Lemus. Journal of the Royal Statistical Society, 88(1):

109–110, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341582.

E:1925:BRBb

[444] W. P. E. Book review: Mortality Experience of Government Life Annuitants, 1900–20, by Government Actuary and the Actuary to the National Debt Commissioners. Journal of the Royal Statistical Society, 88(1):110–112, January 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341583.

M:1925:BRB

[445] F. A. A. M. Book review: The Mortality of Annuitants, 1900–20, by W. Palan Elderton; H. J. P. Oakley. Journal of the Royal Statistical Society, 88(1):112– 115, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341584.

Z:1925:BRBa

[446] A. Z. Book review: On Periodicities in Climatic and Economic Phenomena and their Covariation, by Axel F. Enstrom. Journal of the Royal Statistical Society, 88(1):115–116, January 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341585.

Z:1925:BRBb

[447] A. Z. Book review: Generating Economic Cycles, by H. L. Moore. Journal of the Royal Statistical Society, 88(1): 116–118, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341586.

Z:1925:BRBc

[448] A. Z. Book review: A Study of the Cyclical Fluctuations Occurring in the National Bank System during the Years 1903–21, by Lincoln W. Hall. Journal of the Royal Statistical Society, 88(1): 118–119, January 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341587.

S:1925:BRBa

[449] J. C. S. Book review: Financial Reform, by Henry Higgs. Journal of the Royal Statistical Society, 88(1):119–121, January 1925. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341588.

H:1925:BRBa

[450] H. H. Book review: The Science of Public Finance, by G. Findlay Shirras. Journal of the Royal Statistical Society, 88 (1):121–123, January 1925. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341589.

P:1925:BRBa

[451] G. A. P. Book review: Russian Debts and Russian Reconstruction: A Study of the Relation of Russia’s Foreign Debt to her Economic Recovery, by Leo Pasvolsky; Harold G. Moulton. Journal of the Royal Statistical Society, 88(1):123– 124, January 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341590.

R:1925:BRBa

[452] R. H. R. Book review: Annuaire International de Statistique Agricole, 1923. Journal of the Royal Statistical Society, 88(1):124–125, January 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341591.

B:1925:BRBa

[453] J. B. Book review: I Fondamenti Scientifici della Riforma Economica, by Achille Loria. Journal of the Royal Statistical Society, 88(1):125–128, January 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341592.

S:1925:BRBb

[454] E. C. S. Book review: Is Unemployment Inevitable? An Analysis and a Forecast. Journal of the Royal Statistical Society, 88(1):128–131, January 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341593.

Anonymous:1925:ONPa

[455] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 88(1):131–135, January 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341594.

Anonymous:1925:CNa

[456] Anonymous. Current notes. Journal of the Royal Statistical Society, 88(1): 136–151, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341595.

B:1925:AM

[457] J. B. Alfred Marshall. Journal of the Royal Statistical Society, 88(1):152– 156, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341596.

Anonymous:1925:SEAa

[458] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 88(1): 157–164, January 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341597.

Anonymous:1925:LALa

[459] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 88(1):165–176, January 1925. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341598.

Anonymous:1925:PRa

[460] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 88(1): 177–183, January 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341599.

Thompson:1925:AC

[461] R. J. Thompson. An agricultural census. Journal of the Royal Statistical Society, 88(2):185–220, March 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341269.

Russell:1925:FSE

[462] E. S. Russell and T. Edser. The fishery statistics of England and Wales. Journal of the Royal Statistical Society, 88(2): 221–259, March 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341270.

Anonymous:1925:WPC

[463] Anonymous. Wholesale prices of commodities in 1924. Journal of the Royal Statistical Society, 88(2):260– 278, March 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341271.

R:1925:BRBb

[464] E. C. R. Book review: An Introduction to the Mathematical Analysis of Statistics, by C. H. Forsyth. Journal of the Royal Statistical Society, 88(2): 279–280, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341272.

Z:1925:BRBd

[465] A. Z. Book review: The Nature and Purpose of the Measurement of Social Phenomena, by A. L. Bowley. Journal of the Royal Statistical Society, 88 (2):281, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341273.

Z:1925:BRBe

[466] A. Z. Book review: Smithsonian Mathematical Formulae and Tables of Elliptic Functions. Journal of the Royal Statistical Society, 88(2):281–282, March 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341274.

S:1925:BRBc

[467] J. C. S. Book review: Malthus and His Work, by James Bonar. Journal of the Royal Statistical Society, 88(2): 282–284, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341275.

B:1925:BRBb

[468] J. A. B. Book review: The Population of India: A Comparative Study, by Brij Narain. Journal of the Royal Statistical Society, 88(2):284–286, March 1925. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341276.

C:1925:BRBa

[469] W. H. C. Book review: Current Problems in Finance and Government, by Josiah Stamp. Journal of the Royal Statistical Society, 88(2):286–289, March 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341277.

B:1925:BRBc

[470] J. B. Book review: Le Peripezie Monetarie del Dopo Guerra, by A. Loria. Journal of the Royal Statistical Society, 88(2):289–290, March 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341278.

S:1925:BRBd

[471] E. C. S. Book review: The Problem of Business Forecasting, by Warren M. Persons; William Trufant Foster; Albert J. Hettinger. Journal of the Royal Statistical Society, 88(2):290–292, March 1925. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341279.

H:1925:BRBb

[472] C. W. H. Book review: The Role of the State in the Provision of Railways, by H. M. Jagtiani; William Acworth. Journal of the Royal Statistical Society, 88(2): 292–293, March 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341280.

H:1925:BRBc

[473] C. W. H. Book review: The Economics of Road Transport, by K. G. Fenelon. Journal of the Royal Statistical Society, 88(2):294, March 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341281.

Anonymous:1925:ONPb

[474] Anonymous. Other new publications. Journal of the Royal Statistical Society, 88(2):295–297, March 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341282.

Anonymous:1925:CNb

[475] Anonymous. Current notes. Journal of the Royal Statistical Society, 88(2): 298–308, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341283.

Anonymous:1925:SEAb

[476] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 88(2): 309–312, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341284.

Anonymous:1925:LALb

[477] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 88(2):313–324, March 1925. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341285.

Anonymous:1925:PRb

[478] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 88(2): 325–334, March 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341286.

Hurcomb:1925:ORS

[479] C. W. Hurcomb. Official railway statistics in Great Britain. Journal of the Royal Statistical Society, 88(3):335–378, May 1925. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341415.

Hoare:1925:BLU

[480] Alfred Hoare. The bearing of labour unrest upon the path to be taken to sound currency. Journal of the Royal Statistical Society, 88(3):379–402, May 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341416.

Coates:1925:DTT

[481] W. H. Coates. Double taxation and tax evasion. Journal of the Royal Statistical Society, 88(3):403–427, May 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341417.

Crump:1925:NYE

[482] Norman Crump. The New York Exchange: A reply to Professor Cassel. Journal of the Royal Statistical Society, 88(3):428–432, May 1925. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341418. See [287].

E:1925:BRBc

[483] F. Y. E. Book review: A Mathematical Theory of Evolution Based on the Conclusions of Dr. J. C. Willis, F.R.S, by G. Udny Yule. Journal of the Royal Statistical Society, 88(3):433–436, May 1925. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341419.

Y:1925:BRB

[484] G. U. Y. Book review: The Life, Letters and Labours of Francis Galton, by Karl Pearson. Journal of the Royal Statistical Society, 88(3):436–442, May 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341420.

Z:1925:BRBf

[485] A. Z. Book review: An Introduction to Mathematical Probability, by J. L. Coolidge. Journal of the Royal Statistical Society, 88(3):442–443, May 1925. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341421.

N:1925:BRBb

[486] E. M. N. Book review: Complementi di Matematiche Generali, by Filadelfo Insolera. Journal of the Royal Statistical Society, 88(3):443, May 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341422.

R:1925:BRBc

[487] E. C. R. Book review: Statistical Methods Applied to Economics and Business, by Frederick Cecil Mills. Journal of the Royal Statistical Society, 88(3):444, May 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341423.

Z:1925:BRBg

[488] A. Z. Book review: Social Consequences of Business Cycles, by Maurice B. Hexter. Journal of the Royal Statistical Society, 88(3):444–445, May 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341424.

H:1925:BRBd

[489] C. W. H. Book review: The Elements of Railway Economics, by William M. Acworth. Journal of the Royal Statistical Society, 88(3):445–448, May 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341425.

B:1925:BRBd

[490] W. A. B. Book review: Miners’ Wages and the Cost of Coal, by Isador Lubin; Council and Staff of the Institute of Economics. Journal of the Royal Statistical Society, 88(3):448–451, May 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341426.

R:1925:BRBd

[491] H. R. Book review: Experiments in State Control at the War Office and the Ministry of Food, by E. M. H. Lloyd. Journal of the Royal Statistical Society, 88 (3):451–453, May 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341427.

Anonymous:1925:ONPc

[492] Anonymous. Other new publications. Journal of the Royal Statistical Society, 88(3):453–457, May 1925. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341428.

Anonymous:1925:CNc

[493] Anonymous. Current notes. Journal of the Royal Statistical Society, 88(3): 458–466, May 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341429.

Paish:1925:SWM

[494] George Paish. Sir William Mitchell Acworth. Journal of the Royal Statistical Society, 88(3):467–470, May 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341430.

Y:1925:RD

[495] G. U. Y. Reginald Dudfield. Journal of the Royal Statistical Society, 88(3): 470–473, May 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341431.

Anonymous:1925:SEAc

[496] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 88(3): 474–479, May 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341432.

Anonymous:1925:LALc

[497] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 88(3):480–487, May 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341433.

Anonymous:1925:VIb

[498] Anonymous. Volume information. Journal of the Royal Statistical Society, 88 (4):664–671, July 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341548.

Shaw:1925:WMI

[499] Napier Shaw. The week or month as an intermediate time-unit for statistical purposes. Journal of the Royal Statistical Society, 88(4):489–512, July 1925. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341549.

Shirras:1925:TCB

[500] G. Findlay Shirras. Taxable capacity and the burden of taxation and public debt. Journal of the Royal Statistical Society, 88(4):513–556, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341550.

Edgeworth:1925:EPI

[501] F. Y. Edgeworth. The element of probability in index-numbers. Journal of the Royal Statistical Society, 88(4):557–575, July 1925. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341551.

Rhodes:1925:GLE

[502] E. C. Rhodes. On the generalized law of error. Journal of the Royal Statistical Society, 88(4):576–580, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341552.

Anonymous:1925:RCF

[503] Anonymous. Report of the Council for the financial year ended December 31, 1924, and for the sessional year ending June 16, 1925, presented at the ninety-first Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 16, 1925. Journal of the Royal Statistical Society, 88(4):581–598, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341553.

Anonymous:1925:PNF

[504] Anonymous. Proceedings of the ninetyfirst Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, on Tuesday, June 16, 1925. Journal of the Royal Statistical Society, 88(4):599–600, July 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341554.

S:1925:BRBe

[505] C. P. S. Book review: Papers Relating to Political Economy, by F. Y. Edgeworth. Journal of the Royal Statistical Society, 88(4):601–603, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341555.

N:1925:BRBc

[506] E. M. N. Book review: Causal Factors in Infant Mortality, by R. M. Woodbury. Journal of the Royal Statistical Society, 88(4):603–605, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341556.

R:1925:BRBe

[507] R. H. R. Book review: Milk and Milk Products: Statistical Survey of Production and Trade. Journal of the Royal Statistical Society, 88(4):605–606, July 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341557.

R:1925:BRBf

[508] R. H. R. Book review: Report of the Committee on Stabilization of Agricultural Prices. Journal of the Royal Statistical Society, 88(4):606–608, July 1925. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341558.

R:1925:BRBg

[509] R. H. R. Book review: Co-Operative Marketing of Agricultural Produce in England and Wales. Journal of the Royal Statistical Society, 88(4):608–609, July 1925. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341559.

H:1925:BRBe

[510] R. G. H. Book review: The Discount Policy of the Federal Reserve System, by B. H. Beckhart; The Federal Reserve System in Operation, by E. A. Goldenweiser. Journal of the Royal Statistical Society, 88(4):609–611, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341560.

C:1925:BRBb

[511] N. E. C. Book review: The Return to Gold, by T. E. Gregory. Journal of the Royal Statistical Society, 88(4):611–612, July 1925. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341561.

F:1925:BRB

[512] A. W. F. Book review: Canada’s Balance of International Indebtedness, 1900–13: An Inductive Study on the Theory of International Trade, by Jacob Viner. Journal of the Royal Statistical Society, 88(4):612–613, July 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341562.

P:1925:BRBb

[513] G. A. P. Book review: The Austrian Crown, its Depreciation and Stabilization, by J. van Walre de Bordes; Henry Strakosch. Journal of the Royal Statistical Society, 88(4):614–615, July 1925. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341563.

Z:1925:BRBh

[514] A. Z. Book review: Russian Currency and Banking, 1914–1924, by S. S. Katzenellenbaum. Journal of the Royal Statistical Society, 88(4):615–616, July 1925. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341564.

R:1925:BRBh

[515] R. H. R. Book review: England on the Eve of the Industrial Revolution, by L. W. Moffit. Journal of the Royal Statistical Society, 88(4):616–617, July 1925. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341565.

B:1925:BRBe

[516] M. B. Book review: Iron and Steel in the Industrial Revolution, by T. S. Ashton. Journal of the Royal Statistical Society, 88(4):618–619, July 1925. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341566.

T:1925:BRB

[517] C. T. Book review: The Bases of Indian Economy, by B. G. Bhatnagar. Journal of the Royal Statistical Society, 88 (4):619–620, July 1925. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341567.

A:1925:BRB

[518] J. E. A. Book review: The Building of Satellite Towns, by C. B. Purdom. Journal of the Royal Statistical Society, 88 (4):620–622, July 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341568.

Anonymous:1925:ONPd

[519] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 88(4):622–627, July 1925. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341569.

Anonymous:1925:CNd

[520] Anonymous. Current notes. Journal of the Royal Statistical Society, 88(4): 628–640, July 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341570.

Anonymous:1925:SEAd

[521] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 88(4): 641–645, July 1925. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341571.

Anonymous:1925:LALd

[522] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 88(4):646–653, July 1925. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341572.

Anonymous:1925:ALA

[523] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 88(4):654–663, July 1925. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341573.

Anonymous:1926:VIa

[524] Anonymous. Volume information. Journal of the Royal Statistical Society, 89 (1):i–vii, January 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341481.

Yule:1926:WDW

[525] G. Udny Yule. Why do we sometimes get nonsense-correlations between timeseries? — A study in sampling and the nature of time-series. Journal of the Royal Statistical Society, 89(1):1– 63, January 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341482.

Anonymous:1926:PM

[526] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 89(1):64–69, January 1926. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341483.

Rew:1926:ISI

[527] Henry Rew. The International Statistical Institute and its sixteenth session. Journal of the Royal Statistical Society, 89(1):70–95, January 1926. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341484.

Shimmin:1926:DEW

[528] Arnold N. Shimmin. Distribution of employment in the wool textile industry of the West Riding of Yorkshire. Journal of the Royal Statistical Society, 89(1): 96–120, January 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/

2341485. See discussion [529].

Anonymous:1926:DMS

[529] Anonymous. Discussion on Mr. Shimmin’s paper. Journal of the Royal Statistical Society, 89(1):120–128, January 1926. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2341486. See

[528].

Edgeworth:1926:MRC

[530] F. Y. Edgeworth. Mr. Rhodes’ curve and the method of adjustment. Journal of the Royal Statistical Society, 89(1): 129–143, January 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341487.

I:1926:BRBa

[531] L. I. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 89(1):144–145, January 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341488.

R:1926:BRBa

[532] E. C. R. Book review: The Fundamentals of Statistics, by L. L. Thurstone. Journal of the Royal Statistical Society, 89(1):145–146, January 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341489.

R:1926:BRBb

[533] E. C. R. Book review: Elements of Statistics, by Frederick C. Kent. Journal of the Royal Statistical Society, 89(1): 146–147, January 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341490.

S:1926:BRBa

[534] F. S. Book review: Annals of Eugenics, by Karl Pearson; Ethel M. Elderton.

Journal of the Royal Statistical Society, 89(1):147–151, January 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341491.

N:1926:BRBa

[535] E. M. N. Book review: Studies in Human Biology, by Raymond Pearl. Journal of the Royal Statistical Society, 89 (1):151–152, January 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341492.

A:1926:BRBa

[536] J. E. A. Book review: The Confessions of a Capitalist, by E. J. P. Benn. Journal of the Royal Statistical Society, 89(1): 152–154, January 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341493.

C:1926:BRBa

[537] L. R. C. Book review: Labour and Housing in Bombay: A Study in the Economic Conditions of the Wage-Earning Classes in Bombay, by A. R. Burnett-Hurst; Stanley Reed. Journal of the Royal Statistical Society, 89(1):154– 156, January 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341494.

R:1926:BRBc

[538] R. A. S. R. Book review: American Petroleum: Supply and Demand. Journal of the Royal Statistical Society, 89 (1):156–158, January 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341495.

M:1926:BRBa

[539] H. W. M. Book review: The Present Position and Prospects in the Wool Textile Industry, by Arnold N. Shimmin. Journal of the Royal Statistical Society, 89 (1):158–159, January 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341496.

C:1926:BRBb

[540] H. F. C. Book review: The Economic Development of the British Overseas Empire, by L. C. A. Knowles. Journal of the Royal Statistical Society, 89 (1):160–161, January 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341497.

S:1926:BRBb

[541] J. C. S. Book review: The Economic Laws of Art Production, by Hubert Llewellyn Smith. Journal of the Royal Statistical Society, 89(1):161–163, January 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341498.

Anonymous:1926:ONPa

[542] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 89(1):163–167, January 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341499.

Anonymous:1926:CNa

[543] Anonymous. Current notes. Journal of the Royal Statistical Society, 89(1): 168–181, January 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341500.

R:1926:SAB

[544] R. H. R. and S. de J. Sir Athelstane Baines. Journal of the Royal Statistical Society, 89(1):182–184, January 1926. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341501.

Anonymous:1926:SEAa

[545] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 89(1): 185–193, January 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341502.

Anonymous:1926:LALa

[546] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 89(1):194–209, January 1926. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341503.

Anonymous:1926:PRa

[547] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 89(1): 210–216, January 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341504.

Thompson:1926:PBD

[548] R. J. Thompson. The productivity of British and Danish farming. Journal of the Royal Statistical Society, 89(2): 217–255, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341310.

Scott:1926:SAA

[549] H. C. Scott. Some administrative aspects of the Rignano scheme of inheritance taxation. Journal of the Royal Statistical Society, 89(2):256–282, March 1926. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341311.

Anonymous:1926:WPC

[550] Anonymous. Wholesale prices of commodities in 1925. Journal of the Royal Statistical Society, 89(2):283– 299, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341312.

Bowley:1926:IPI

[551] A. L. Bowley. The influence on the precision of index-numbers of correlation between the prices of commodities. Journal of the Royal Statistical Society, 89(2): 300–319, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341313.

Greenwood:1926:PTT

[552] Major Greenwood. Professor Tschuprow on the theory of correlation. Journal of the Royal Statistical Society, 89(2): 320–325, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341314.

Soper:1926:MHS

[553] H. E. Soper. The moments of the hypergeometric series. Journal of the Royal Statistical Society, 89(2):326– 328, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341315.

P:1926:BRBa

[554] L. L. P. Book review: Memorials of Alfred Marshall, by A. C. Pigou. Journal of the Royal Statistical Society, 89(2): 329–333, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341316.

P:1926:BRBb

[555] G. A. P. Book review: Nationalokonomie und Statistik, by Rudolf Meerwarth. Journal of the Royal Statistical Society, 89(2):333–335, March 1926. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341317.

T:1926:BRB

[556] H. L. T. Book review: Statistics and their Application to Commerce, by A. Lester Boddington. Journal of the Royal Statistical Society, 89(2):335– 336, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341318.

N:1926:BRBb

[557] E. M. N. Book review: Om Seksualproporsjonen ved Fedselen. Journal of the Royal Statistical Society, 89(2): 336–339, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341319.

Z:1926:BRBa

[558] A. Z. Book review: Economics of Business Cycles, by Arthur B. Adams. Journal of the Royal Statistical Society, 89 (2):339–340, March 1926. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341320.

C:1926:BRBc

[559] W. H. C. Book review: Monetary Stability, by J. R. Bellerby. Journal of the Royal Statistical Society, 89(2): 340–342, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341321.

H:1926:BRBa

[560] R. G. H. Book review: Financial Reconstruction in England, 1815–22, by A. W. Acworth. Journal of the Royal Statistical Society, 89(2):342–343, March 1926. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341322.

H:1926:BRBb

[561] R. G. H. Book review: The Financial Crisis of France, by George Peel. Journal of the Royal Statistical Society, 89 (2):343–346, March 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341323.

C:1926:BRBd

[562] L. R. C. Book review: Wealth and Taxable Capacity of India, by K. T. Shah; K. J. Khambata; The Wealth of India, by P. A. Wadia; G. N. Joshi. Journal of the Royal Statistical Society, 89(2): 346–348, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341324.

P:1926:BRBc

[563] G. A. P. Book review: Currency Problems and Policy of the Soviet Union, by L. N. Yurovsky. Journal of the Royal Statistical Society, 89(2):348– 349, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341325.

H:1926:BRBc

[564] A. B. H. Book review: Has Poverty Diminished? A Sequel to Livelihood and Poverty, by A. L. Bowley; Margaret H. Hogg. Journal of the Royal Statistical Society, 89(2):350–352, March 1926. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341326.

C:1926:BRBe

[565] C. E. C. Book review: Family Allowances. Journal of the Royal Statistical Society, 89(2):352–353, March 1926. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341327.

C:1926:BRBf

[566] C. E. C. Book review: Wages and the Family, by Paul H. Douglas. Journal of the Royal Statistical Society, 89(2): 353–354, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341328.

Anonymous:1926:ONPb

[567] Anonymous. Other new publications. Journal of the Royal Statistical Society, 89(2):355–359, March 1926. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341329.

Anonymous:1926:CNb

[568] Anonymous. Current notes. Journal of the Royal Statistical Society, 89(2): 360–370, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341330.

P:1926:FYE

[569] L. L. P. Francis Ysidro Edgeworth. Journal of the Royal Statistical Society, 89 (2):371–377, March 1926. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341331.

Anonymous:1926:SEAb

[570] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 89(2): 378–382, March 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341332.

Anonymous:1926:LALb

[571] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 89(2):383–393, March 1926. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341333.

Anonymous:1926:PRb

[572] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 89(2): 394–403, March 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341334.

Edge:1926:GMD

[573] P. Granville Edge. The growth of mortality due to motor vehicles in England and Wales, 1904–23. Journal of the Royal Statistical Society, 89(3):405–451, May 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341716.

Macrosty:1926:SBS

[574] Henry W. Macrosty. Statistics of British shipping. Journal of the Royal Statistical Society, 89(3):452–543, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341717.

Rhodes:1926:CTS

[575] E. C. Rhodes. The comparison of two sets of observations. Journal of the Royal Statistical Society, 89(3):544–552, May 1926. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341718.

Connor:1926:FMP

[576] L. R. Connor. Fertility of marriage and population growth. Journal of the Royal Statistical Society, 89(3):553–566, May 1926. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341719.

Ashton:1926:GTB

[577] T. S. Ashton. The growth of textile businesses in the Oldham District, 1884– 1924. Journal of the Royal Statistical Society, 89(3):567–583, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341720.

R:1926:BRBd

[578] E. C. R. Book review: An Introduction to Statistical Methods, by Horace Secrist. Journal of the Royal Statistical Society, 89(3):584–585, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341721.

R:1926:BRBe

[579] E. C. R. Book review: Handbook of Mathematical Statistics, by H. L. Rietz. Journal of the Royal Statistical Society, 89(3):585, May 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341722.

H:1926:BRBd

[580] A. B. H. Book review: Population Problems in the United States and Canada, by Louis I. Dublin; The Natural Increase of Mankind, by J. Shirley Sweeney. Journal of the Royal Statistical Society, 89(3):586–587, May 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341723.

C:1926:BRBg

[581] L. R. C. Book review: The Evolution of Provincial Finance in British India, by B. R. Ambedkar; Edwin R. A. Seligman. Journal of the Royal Statistical Society, 89(3):587–588, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341724.

S:1926:BRBc

[582] E. C. S. Book review: The Secret of High Wages, by Bertram Austin; W. Francis Lloyd; Walter T. Layton. Journal of the Royal Statistical Society, 89(3):588–591, May 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341725.

C:1926:BRBh

[583] C. E. C. Book review: Family Allowances in Practice, by Hugh H. R. Vibart. Journal of the Royal Statistical Society, 89(3):591–592, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341726.

R:1926:BRBf

[584] R. H. R. Book review: A State Trading Adventure, by Frank H. Coller. Journal of the Royal Statistical Society, 89 (3):592–593, May 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341727.

R:1926:BRBg

[585] R. H. R. Book review: Report Upon Large-Scale Co-Operative Marketing in the United States of America, by R. B. Forrester; The Co-Operative Purchase of Agricultural Requisites. Journal of the Royal Statistical Society, 89 (3):593–594, May 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341728.

R:1926:BRBh

[586] R. A. S. R. Book review: Coal, by Edward T. Devine. Journal of the Royal Statistical Society, 89(3):594–599, May 1926. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341729.

H:1926:BRBe

[587] A. B. H. Book review: Variations in the Composition of Milk, by J. F. Tocher. Journal of the Royal Statistical Society, 89(3):599–600, May 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341730.

R:1926:BRBi

[588] R. H. R. Book review: Report on the Trade in Refrigerated Beef, Mutton and Lamb. Journal of the Royal Statistical Society, 89(3):600, May 1926. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341731.

B:1926:BRBa

[589] D. M. B. Book review: Germany’s Industrial Revival, by Philip Dawson. Journal of the Royal Statistical Society, 89(3):600–601, May 1926. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341732.

N:1926:BRBc

[590] E. M. N. Book review: La Salute Pubblica in Italia Durante e Dopo la Guerra, by Giorgio Mortara. Journal of the Royal Statistical Society, 89(3):601–603, May 1926. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341733.

P:1926:BRBd

[591] G. A. P. Book review: The Economic Development of Russia, 1905–1914, with Special Reference to Trade, Industry and Finance, by Margaret S. Miller. Journal of the Royal Statistical Society, 89 (3):604–605, May 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341734.

C:1926:BRBi

[592] C. E. C. Book review: Factory Legislation and its Administration, 1891–1924, by H. A. Mess. Journal of the Royal Statistical Society, 89(3):605–606, May 1926. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341735.

A:1926:BRBb

[593] P. A. Book review: Local Government in Many Lands, by G. Montagu Harris. Journal of the Royal Statistical Society, 89(3):606–607, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341736.

Anonymous:1926:ONPc

[594] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 89(3):608–610, May 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341737.

Anonymous:1926:CNc

[595] Anonymous. Current notes. Journal of the Royal Statistical Society, 89(3): 611–618, May 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341738.

I:1926:AAT

[596] L. I. Alexander Alexandrovitch Tschuprow. Journal of the Royal Statistical Society, 89(3):619–622, May 1926. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341739.

Anonymous:1926:SEAc

[597] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 89(3): 623–627, May 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341740.

Anonymous:1926:LALc

[598] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 89(3):628–634, May 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341741.

Anonymous:1926:VIb

[599] Anonymous. Volume information. Journal of the Royal Statistical Society, 89 (4):807–815, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341245.

Snow:1926:SOT

[600] E. C. Snow. Some observations on trade forecasting and prices. Journal of the Royal Statistical Society, 89(4):635–679, July 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341246.

Jones:1926:SSC

[601] D. Caradog Jones. Some statistics concerning occupational invalidity– economic, mental and moral. Journal of the Royal Statistical Society, 89(4): 680–702, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341247.

Gini:1926:CIM

[602] Corrado Gini. The contributions of Italy to modern statistical methods. Journal of the Royal Statistical Society, 89(4): 703–724, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341248.

Anonymous:1926:RCF

[603] Anonymous. Report of the Council for the financial year ended December 31, 1925, and for the sessional year ending June 15, 1926, presented at the ninetysecond Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 15, 1926. Journal of the Royal Statistical Society, 89(4):725–742, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341249.

Anonymous:1926:PNS

[604] Anonymous. Proceedings of the ninetysecond Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, on June 15, 1926. Journal of the Royal Statistical Society, 89(4):743, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341250.

I:1926:BRBb

[605] L. I. Book review: A Contribution to the Study of the Human Factor in the Causation of Accidents, by E. M. Newbold. Journal of the Royal Statistical Society, 89(4):744–746, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341251.

H:1926:BRBf

[606] A. B. H. Book review: A Study in Vocational Guidance, by Industrial Fatigue Research Board; National Institute of Industrial Psychology. Journal of the Royal Statistical Society, 89(4):746–748, July 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341252.

F:1926:BRB

[607] A. W. F. Book review: The Christian Ethic as an Economic Factor: The Social Service Lecture, 1926, by Josiah Stamp. Journal of the Royal Statistical Society, 89(4):748–749, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341253.

M:1926:BRBb

[608] H. W. M. Book review: The Economics of Private Enterprise, by J. H. Jones.

Journal of the Royal Statistical Society, 89(4):749–750, July 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341254.

H:1926:BRBg

[609] R. G. H. Book review: British War Budgets, by F. W. Hirst; J. E. Allen. Journal of the Royal Statistical Society, 89 (4):750–752, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341255.

M:1926:BRBc

[610] H. W. M. Book review: The Amalgamation Movement in English Banking, 1825–1924, by Joseph Sykes. Journal of the Royal Statistical Society, 89(4): 752–753, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341256.

C:1926:BRBj

[611] N. E. C. Book review: Banking Policy and the Price Level, by D. H. Robertson. Journal of the Royal Statistical Society, 89(4):753–754, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341257.

C:1926:BRBk

[612] C. E. C. Book review: Wages and the State, by E. M. Burns. Journal of the Royal Statistical Society, 89(4):754–756, July 1926. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341258.

C:1926:BRBl

[613] L. R. C. Book review: Co-Operation at Home and Abroad, by C. R. Fay. Journal of the Royal Statistical Society, 89 (4):756–757, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341259.

Z:1926:BRBb

[614] A. Z. Book review: The Cotton Growing Countries, Present and Potential. Journal of the Royal Statistical Society, 89 (4):757–758, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341260.

W:1926:BRB

[615] A. D. W. Book review: Memoires et Documents pour servir a l’Histoire du Commerce et de l’Industrie en France: Le Commerce Maritime de la Bretagne au XVII Siecle, by Henri See; Quelques Professions Connues, Inconnues et Meconnues. by Julien Hayem. Journal of the Royal Statistical Society, 89(4):758–759, July 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341261.

B:1926:BRBb

[616] C. F. B. Book review: Constitution, Functions and Finance of Indian Municipalities, by K. T. Shah; G. H. Bahadurji. Journal of the Royal Statistical Society, 89(4):759–761, July 1926. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341262.

Anonymous:1926:ONPd

[617] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 89(4):761–765, July 1926. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341263.

Anonymous:1926:CNd

[618] Anonymous. Current notes. Journal of the Royal Statistical Society, 89(4): 766–784, July 1926. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341264.

M:1926:HST

[619] B. M. The Hon. Sir Timothy Coghlan, K.C.M.G., 1857–1926. Journal of the Royal Statistical Society, 89(4):785, July 1926. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341265.

Anonymous:1926:SEAd

[620] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 89(4): 786–789, July 1926. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341266.

Anonymous:1926:LALd

[621] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 89(4):790–796, July 1926. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341267.

Anonymous:1926:ALA

[622] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 89(4):797–806, July 1926. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341268.

Anonymous:1927:VIa

[623] Anonymous. Volume information. Journal of the Royal Statistical Society, 90 (1):i–vii, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341335.

Anonymous:1927:FM

[624] Anonymous. Front matter. Journal of the Royal Statistical Society, 90(1):ii–iii, ???? 1927. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341336.

Anonymous:1927:GCC

[625] Anonymous. German currency: Its collapse and recovery, 1920–26. Journal of the Royal Statistical Society, 90(1):1–40, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341337.

Anonymous:1927:PM

[626] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 90(1):41–44, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341338.

Macrosty:1927:IDU

[627] Henry W. Macrosty. Inflation and deflation in the United States and the United Kingdom, 1919–23. Journal of the Royal Statistical Society, 90(1):45–134, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341339.

Rhodes:1927:PMS

[628] E. C. Rhodes. The precision of means and standard deviations when the individual errors are correlated. Journal of the Royal Statistical Society, 90(1): 135–143, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341340.

R:1927:BRBa

[629] E. C. R. Book review: A First Course in Statistical Method, by G. Irving Gavett. Journal of the Royal Statistical Society, 90(1):144–145, ???? 1927. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341341.

C:1927:BRBa

[630] L. R. C. Book review: An Introduction to the Methods of Economic Statistics, by William Leonard Crum; Alson Currie Patton. Journal of the Royal Statistical Society, 90(1):145–146, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341342.

S:1927:BRBa

[631] F. S. Book review: Sur un Problème d’Économie Pure, by Ragnar Frisch. Journal of the Royal Statistical Society, 90(1):146–147, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341343.

C:1927:BRBb

[632] L. R. C. Book review: Population Problems of the Age of Malthus, by G. Talbot Griffith. Journal of the Royal Statistical Society, 90(1):148–149, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341344.

W:1927:BRB

[633] H. M. W. Book review: Poverty, Nutrition and Growth, by D. Noel Paton; Leonard Findlay. Journal of the Royal Statistical Society, 90(1):149–151, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341345.

S:1927:BRBb

[634] J. C. S. Book review: The Economic Problem, by R. G. Hawtrey. Journal of the Royal Statistical Society, 90(1): 151–153, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341346.

S:1927:BRBc

[635] J. C. S. Book review: National Economics for Britain’s Day of Need–The Solution of the Unemployment Problem, by Edward Batten. Journal of the Royal Statistical Society, 90(1):153–155, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341347.

B:1927:BRBa

[636] C. F. B. Book review: The Economics of Wages and Labour, by Nora Milnes. Journal of the Royal Statistical Society, 90(1):156–158, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341348.

A:1927:BRB

[637] J. E. A. Book review: To-Day and To-Morrow, by Henry Ford; Samuel Crowther. Journal of the Royal Statistical Society, 90(1):158–160, ???? 1927. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341349.

V:1927:BRBa

[638] H. V. Book review: Business Economics, by William Ashley. Journal of the Royal Statistical Society, 90(1): 160–161, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341350.

S:1927:BRBd

[639] E. C. S. Book review: The Economics of The Leather Industry, by B. Ramachandra Rau. Journal of the Royal Statistical Society, 90(1):161–163, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341351.

H:1927:BRBa

[640] R. G. H. Book review: A Discourse Upon Usury, by Thomas Wilson; R. H. Tawney. Journal of the Royal Statistical Society, 90(1):163–164, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341352.

S:1927:BRBe

[641] J. C. S. Book review: Income in the Various States–Its Sources and Distribution, 1919, 1920 and 1921, by Maurice Lever. Journal of the Royal Statistical Society, 90(1):164–166, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341353.

C:1927:BRBc

[642] L. R. C. Book review: Money and the Money Market in India, by P. A. Wadia; G. N. Joshi. Journal of the Royal Statistical Society, 90(1):166, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341354.

C:1927:BRBd

[643] L. R. C. Book review: The Cost of Living, by Norman Dearle. Journal of the Royal Statistical Society, 90(1):166–167, ???? 1927. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341355.

R:1927:BRBb

[644] E. C. R. Book review: The Young Delinquent, by Cyril Burt. Journal of the Royal Statistical Society, 90(1):167–168, ???? 1927. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341356.

Anonymous:1927:ONPa

[645] Anonymous. Other new publications. Journal of the Royal Statistical Society, 90(1):169–171, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341357.

Anonymous:1927:CNa

[646] Anonymous. Current notes. Journal of the Royal Statistical Society, 90(1): 172–192, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341358.

M:1927:LE

[647] B. M. Lord Emmott. Journal of the Royal Statistical Society, 90(1):193, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341359.

R:1927:SJW

[648] H. R. Sir James Wilson. Journal of the Royal Statistical Society, 90(1):194, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341360.

Anonymous:1927:SEAa

[649] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 90(1): 195–202, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341361.

Anonymous:1927:LALa

[650] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 90(1):203–217, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341362.

Anonymous:1927:PRa

[651] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 90(1): 218–224, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341363.

Flux:1927:IIP

[652] A. W. Flux. Indices of industrial productive activity. Journal of the Royal Statistical Society, 90(2):225–271, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341766.

Wood:1927:ESS

[653] George H. Wood. An examination of some statistics relating to the wool textile industry. Journal of the Royal Statistical Society, 90(2):272–328, ???? 1927. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341767.

Anonymous:1927:WPC

[654] Anonymous. Wholesale prices of commodities in 1926. Journal of the Royal Statistical Society, 90(2):329–346, ???? 1927. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341768.

Greenwood:1927:HNP

[655] M. Greenwood and L. Isserlis. A historical note on the problem of small samples. Journal of the Royal Statistical Society, 90(2):347–352, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341769.

Coates:1927:RCN

[656] W. H. Coates. Report of the committee on national debt and taxation. Journal of the Royal Statistical Society, 90(2): 353–364, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341770.

Macrosty:1927:SIB

[657] Henry W. Macrosty. Submerged information: Banking. Journal of the Royal Statistical Society, 90(2):365–371, ???? 1927. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341771.

N:1927:BRBa

[658] E. M. N. Book review: Elements of Statistics, by A. L. Bowley. Journal of the Royal Statistical Society, 90(2):

372–373, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341772.

N:1927:BRBb

[659] E. M. N. Book review: A Simple Formula for Welding Curves in Graduating Observational Data, by John Rice Miner; Raymond Pearl. Journal of the Royal Statistical Society, 90(2):373–374, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341773.

R:1927:BRBc

[660] E. C. R. Book review: Elements of Mathematics for Students of Economics and Statistics, by D. Caradog Jones; G. W. Daniels. Journal of the Royal Statistical Society, 90(2):374, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341774.

B:1927:BRBb

[661] A. L. B. Book review: Cyclical Fluctuations. Retail and Wholesale Trade, United States, 1919–1925, by Simon S. Kuznets; Wesley C. Mitchell. Journal of the Royal Statistical Society, 90(2): 374–376, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341775.

M:1927:BRBa

[662] H. W. M. Book review: The National Income, 1924, by Arthur L. Bowley; Josiah Stamp. Journal of the Royal Statistical Society, 90(2):376–378, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341776.

C:1927:BRBe

[663] L. R. C. Book review: Das Deutsche Volkseinkommen, by Erich Rogowski.

Journal of the Royal Statistical Society, 90(2):378–380, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341777.

H:1927:BRBb

[664] R. G. H. Book review: The First Year of the Gold Standard, by T. E. Gregory. Journal of the Royal Statistical Society, 90(2):380–382, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341778.

H:1927:BRBc

[665] A. B. H. Book review: The Need for Eugenic Reform, by Leonard Darwin. Journal of the Royal Statistical Society, 90 (2):382–384, ???? 1927. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341779.

H:1927:BRBd

[666] A. B. H. Book review: Labour Migration in England, 1800–50, by Arthur Redford. Journal of the Royal Statistical Society, 90(2):384–386, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341780.

C:1927:BRBf

[667] C. E. C. Book review: World Migration and Labour, by John W. Brown. Journal of the Royal Statistical Society, 90(2): 386–388, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341781.

B:1927:BRBc

[668] W. A. B. Book review: Some Problems of Wages and Their Regulation in Great Britain Since 1918, by Alan G. B. Fisher. Journal of the Royal Statistical Society, 90(2):388–389, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341782.

Anonymous:1927:ONPb

[669] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 90(2):390–392, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341783.

Anonymous:1927:CNb

[670] Anonymous. Current notes. Journal of the Royal Statistical Society, 90(2): 393–404, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341784.

G:1927:JB

[671] M. G. John Brownlee. Journal of the Royal Statistical Society, 90(2):405–407, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341785.

Anonymous:1927:SEAb

[672] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 90(2): 408–412, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341786.

Anonymous:1927:LALb

[673] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 90(2):413–422, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341787.

Anonymous:1927:PRb

[674] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 90(2): 423–432, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341788.

Watson:1927:NHI

[675] Alfred W. Watson. National health insurance: A statistical review. Journal of the Royal Statistical Society, 90(3): 433–486, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341202.

Newbold:1927:PAS

[676] Ethel M. Newbold. Practical applications of the statistics of repeated events’ particularly to industrial accidents. Journal of the Royal Statistical Society, 90(3):487–547, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341203.

Anderson:1927:LDS

[677] Oskar Anderson. On the logic of the decomposition of statistical series into separate components. Journal of the Royal Statistical Society, 90(3):548–569, ???? 1927. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341204.

Yule:1927:RS

[678] G. Udny Yule. On reading a scale. Journal of the Royal Statistical Society, 90 (3):570–587, ???? 1927. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341205.

B:1927:BRBd

[679] J. B. Book review: Official Papers, by Alfred Marshall. Journal of the Royal Statistical Society, 90(3):588–590, ???? 1927. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341206.

Z:1927:BRBa

[680] A. Z. Book review: An Economic History of Modern Britain, by J. H.

Clapham. Journal of the Royal Statistical Society, 90(3):590–592, ???? 1927. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341207.

P:1927:BRB

[681] G. A. P. Book review: Diminishing Returns in Agriculture, by F. Lester Patton. Journal of the Royal Statistical Society, 90(3):592–594, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341208.

H:1927:BRBe

[682] R. G. H. Book review: La Stabilisation du Franc, by Fernand Leurence. Journal of the Royal Statistical Society, 90(3): 594–595, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341209.

M:1927:BRBb

[683] H. W. M. Book review: Mouvement des Prix et des Salaires Pendant la Guerre, by Lucien March. Journal of the Royal Statistical Society, 90(3):595–596, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341210.

M:1927:BRBc

[684] H. W. M. Book review: Industrial Fluctuations, by A. C. Pigou. Journal of the Royal Statistical Society, 90(3):596–598, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341211.

Z:1927:BRBb

[685] A. Z. Book review: An Academic Study of Some Money Market and Other Statistics, by E. G. Peake. Journal of the Royal Statistical Society, 90(3):598–599, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341212.

Z:1927:BRBc

[686] A. Z. Book review: Business Annals, by Willard Long Thorp. Journal of the Royal Statistical Society, 90(3):599–600, ???? 1927. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341213.

S:1927:BRBf

[687] E. C. S. Book review: The Marketing Problem: How it is Being Tackled in the U.S.A, by Edward T. Elbourne; Josiah Stamp. Journal of the Royal Statistical Society, 90(3):600–601, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341214.

C:1927:BRBg

[688] L. R. C. Book review: The Widening Retail Market and Consumers’ Buying Habits. Journal of the Royal Statistical Society, 90(3):602–603, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341215.

M:1927:BRBd

[689] H. W. M. Book review: The Finance of Foreign Trade, by William F. Spalding. Journal of the Royal Statistical Society, 90(3):603–604, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341216.

V:1927:BRBb

[690] H. V. Book review: Capitalism is Socialism, by J. Taylor Peddie; Gilbert C. Vyle. Journal of the Royal Statistical Society, 90(3):604–605, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341217.

M:1927:BRBe

[691] F. A. A. M. Book review: The Origin and Early History of Insurance, by C. F. Trenerry. Journal of the Royal Statistical Society, 90(3):605–606, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341218.

H:1927:BRBf

[692] W. H. H. Book review: The Vaccination Controversy, by H. Dennis Taylor. Journal of the Royal Statistical Society, 90(3):607–608, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341219.

C:1927:BRBh

[693] L. R. C. Book review: Prohibition at Its Worst, by Irving Fisher. Journal of the Royal Statistical Society, 90(3):608–609, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341220.

Anonymous:1927:ONPc

[694] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 90(3):609–611, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341221.

Anonymous:1927:CNc

[695] Anonymous. Current notes. Journal of the Royal Statistical Society, 90(3): 612–619, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341222.

Anonymous:1927:SEAc

[696] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 90(3): 620–625, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341223.

Anonymous:1927:LALc

[697] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 90(3):626–635, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341224.

Anonymous:1927:VIb

[698] Anonymous. Volume information. Journal of the Royal Statistical Society, 90 (4):835–842, ???? 1927. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341364.

Elbourne:1927:TAS

[699] E. T. Elbourne. Trade association statistics. Journal of the Royal Statistical Society, 90(4):637–684, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341365.

Jones:1927:PWP

[700] D. Caradog Jones. Pre-war and post-war taxation. Journal of the Royal Statistical Society, 90(4):685–728, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341366.

Reed:1927:SLC

[701] Lowell J. Reed and Raymond Pearl. On the summation of logistic curves. Journal of the Royal Statistical Society, 90 (4):729–746, ???? 1927. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341367.

Lehfeldt:1927:WSM

[702] R. A. Lehfeldt. The world’s stock of money. Journal of the Royal Statistical Society, 90(4):747–749, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341368.

Frisch:1927:PEP

[703] Ragnar Frisch. “Sur un problème d’Économie pure”. Journal of the Royal Statistical Society, 90(4):750, ???? 1927. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341369.

Anonymous:1927:RCF

[704] Anonymous. Report of the Council for the financial year ended December 31, 1926, and for the sessional year ending June 21, 1927, presented at the ninetythird Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 21, 1927. Journal of the Royal Statistical Society, 90(4):751–768, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341370.

Anonymous:1927:PNT

[705] Anonymous. Proceedings of the ninetythird Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, on Tuesday, June 21, 1927. Journal of the Royal Statistical Society, 90(4):769–774, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341371.

G:1927:BRB

[706] M. G. Book review: Interpolation, by J. F. Steffensen. Journal of the Royal Statistical Society, 90(4):775–776, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341372.

C:1927:BRBi

[707] E. C. Book review: The Theory of International Prices: History, Criticism, and Re-Stateement, by James W. Angell. Journal of the Royal Statistical Society, 90(4):777–778, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341373.

V:1927:BRBc

[708] H. V. Book review: The Road to Prosperity, by George Paish; Josiah Stamp. Journal of the Royal Statistical Society, 90(4):779–780, ???? 1927. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341374.

C:1927:BRBj

[709] W. H. C. Book review: The Income Tax in Great Britain and the United States, by Harrison B. Spaulding. Journal of the Royal Statistical Society, 90(4):780–782, ???? 1927. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341375.

S:1927:BRBg

[710] J. C. S. Book review: Making the Tariff in the United States, by Thomas Walker Page; Council and Staff of the Institute of Economics. Journal of the Royal Statistical Society, 90(4):782–784, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341376.

H:1927:BRBg

[711] R. G. H. Book review: American Foreign Investments, by Robert W. Dunn. Journal of the Royal Statistical Society, 90(4):784, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341377.

V:1927:BRBd

[712] H. V. Book review: The Industry and Trade of Japan, by S. Uyehara. Journal of the Royal Statistical Society, 90(4):

785–786, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341378.

S:1927:BRBh

[713] E. C. S. Book review: Business Forecasting and its Practical Application, by William Wallace; W. T. Layton. Journal of the Royal Statistical Society, 90(4): 786–788, ???? 1927. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341379.

S:1927:BRBi

[714] E. C. S. Book review: Industrial Combinations in England, by Patrick Fitzgerald. Journal of the Royal Statistical Society, 90(4):788–790, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341380.

H:1927:BRBh

[715] A. B. H. Book review: Health, Wealth, and Population in the Early Days of the Industrial Revolution, by M. C. Buer. Journal of the Royal Statistical Society, 90(4):790–791, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341381.

C:1927:BRBk

[716] L. R. C. Book review: A Study on the Minimum Wage, by J. H. Richardson. Journal of the Royal Statistical Society, 90(4):792–793, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341382.

V:1927:BRBe

[717] H. V. Book review: A Study in Social Economics: The Hunter River Valley, by F. R. E. Mauldon. Journal of the Royal Statistical Society, 90(4):794–795, ???? 1927. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341383.

R:1927:BRBd

[718] R. H. R. Book review: English Farming, Past and Present, by Lord Ernle. Journal of the Royal Statistical Society, 90(4):795, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341384.

Anonymous:1927:ONPd

[719] Anonymous. Other new publications. Journal of the Royal Statistical Society, 90(4):796–801, ???? 1927. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341385.

Anonymous:1927:CNd

[720] Anonymous. Current notes. Journal of the Royal Statistical Society, 90(4): 802–810, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341386.

Higgs:1927:JSN

[721] Henry Higgs. Joseph Shield Nicholson. Journal of the Royal Statistical Society, 90(4):811–813, ???? 1927. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341387.

Anonymous:1927:SEAd

[722] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 90(4): 814–818, ???? 1927. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341388.

Anonymous:1927:LALd

[723] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 90(4):819–825, ???? 1927. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341389.

Anonymous:1927:ALA

[724] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 90(4):826–834, ???? 1927. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341390.

Anonymous:1928:VIa

[725] Anonymous. Volume information. Journal of the Royal Statistical Society, 91 (1):v–vii, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341643.

Anonymous:1928:FM

[726] Anonymous. Front matter. Journal of the Royal Statistical Society, 91(1):i–iii, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341644.

Vigor:1928:CEE

[727] H. D. Vigor. Crop estimates in England. Journal of the Royal Statistical Society, 91(1):1–49, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341645.

Connor:1928:CAD

[728] L. R. Connor. On certain aspects of the distribution of income in the United Kingdom in the years 1913 and 1924. Journal of the Royal Statistical Society, 91(1):50–78, ???? 1928. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341646.

Greenwood:1928:GP

[729] M. Greenwood. Graunt and Petty. Journal of the Royal Statistical Society, 91 (1):79–85, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341647.

S:1928:BRBa

[730] F. S. Book review: Frequency Curves and Correlation, by W. Palin Elderton. Journal of the Royal Statistical Society, 91(1):86–87, ???? 1928. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341648.

H:1928:BRBa

[731] D. H. Book review: A Study of British Genius, by Havelock Ellis. Journal of the Royal Statistical Society, 91(1):87–89, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341649.

H:1928:BRBb

[732] A. B. H. Book review: Criminal Intelligence, by Carl Murchison. Journal of the Royal Statistical Society, 91 (1):89–90, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341650.

W:1928:BRBa

[733] H. M. W. Book review: Infant Mortality and its Causes. With an Appendix on the Trend of Maternal Mortality Rates in the United States, by Robert Morse Woodbury. Journal of the Royal Statistical Society, 91(1):90–92, ???? 1928. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341651.

H:1928:BRBc

[734] W. H. H. Book review: Health Problems in Organized Society, by Arthur Newsholme. Journal of the Royal Statistical Society, 91(1):92–93, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341652.

H:1928:BRBd

[735] R. G. H. Book review: An Economist’s Protest, by Edwin Cannan. Journal of the Royal Statistical Society, 91(1): 93–95, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341653.

C:1928:BRBa

[736] L. R. C. Book review: Die Industriekonjunktur–ein Rechenfehler!, by F. Schmidt. Journal of the Royal Statistical Society, 91(1):95–98, ???? 1928. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341654.

W:1928:BRBb

[737] D. W. W. Book review: Advertising Fluctuations: Seasonal and Cyclical, by William Leonard Crum. Journal of the Royal Statistical Society, 91(1):98–100, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341655.

H:1928:BRBe

[738] R. G. H. Book review: The Stabilisation of the Mark, by Hjalmar Schacht. Journal of the Royal Statistical Society, 91(1):100–102, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341656.

C:1928:BRBb

[739] L. R. C. Book review: Die Sakulare Entwicklung der Kaufkraft des Geldes, by Emil Notz. Journal of the Royal Statistical Society, 91(1):103, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341657.

G:1928:BRBa

[740] J. B. G. Book review: A Primer of Agricultural Economics, by Henry Rew.

Journal of the Royal Statistical Society, 91(1):103–104, ???? 1928. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341658.

B:1928:BRBa

[741] D. M. B. Book review: The American Wool Manufacture, by Arthur Harrison Cole. Journal of the Royal Statistical Society, 91(1):104–105, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341659.

M:1928:BRBa

[742] J. M. Book review: The Iron Industry in Prosperity and Depression, by Homer Bews Vanderblue; William Leonard Crum. Journal of the Royal Statistical Society, 91(1):105–106, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341660.

M:1928:BRBb

[743] J. M. Book review: Machinery and Labour, by George E. Barnett. Journal of the Royal Statistical Society, 91(1): 106–107, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341661.

Anonymous:1928:ONPa

[744] Anonymous. Other new publications. Journal of the Royal Statistical Society, 91(1):107–110, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341662.

Anonymous:1928:CNa

[745] Anonymous. Current notes. Journal of the Royal Statistical Society, 91(1): 111–122, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341663.

M:1928:OLG

[746] B. M. Obituary: Lord George Hamilton, G.C.S.I. Journal of the Royal Statistical Society, 91(1):123, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341664.

Anonymous:1928:SEAa

[747] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 91(1): 124–131, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341665.

Anonymous:1928:LALa

[748] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 91(1):132–144, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341666.

Anonymous:1928:PRa

[749] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 91(1): 145–151, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341667.

Daniels:1928:PWD

[750] G. W. Daniels and J. Jewkes. The postwar depression in the Lancashire cotton industry. Journal of the Royal Statistical Society, 91(2):153–206, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341529.

Stevenson:1928:VSW

[751] T. H. C. Stevenson. The vital statistics of wealth and poverty. Journal of the Royal Statistical Society, 91(2):207–230, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341530.

deJastrezebski:1928:CSA

[752] S. de Jastrezebski. Changes in sex and age constitutions of some representative European populations. Journal of the Royal Statistical Society, 91(2):231–249, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341531.

R:1928:BRBa

[753] E. C. R. Book review: Mathematical Statistics, by Henry Lewis Rietz. Journal of the Royal Statistical Society, 91 (2):250–251, ???? 1928. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341532.

S:1928:BRBb

[754] E. C. S. Book review: Business Cycles: The Problem and its Setting, by Wesley C. Mitchell. Journal of the Royal Statistical Society, 91(2):251–253, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341533.

R:1928:BRBb

[755] L. R. C. Book review: Statistics and their Application to Commerce, by A. Lester Boddington. Journal of the Royal Statistical Society, 91(2):253, ???? 1928. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341534.

R:1928:BRBc

[756] L. R. C. Book review: A Survey of the Social Structure of England and Wales, as Illustrated, by Statistics. by A. M. Carr-Saunders; D. Caradog Jones. Journal of the Royal Statistical Society, 91 (2):254–255, ???? 1928. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341535.

M:1928:BRBc

[757] H. O. M. Book review: Monopolies, Cartels and Trusts in British Industry, by Hermann Levy. Journal of the Royal Statistical Society, 91(2):255–256, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341536.

H:1928:BRBf

[758] A. B. H. Book review: The Physique of Women in Industry: A Contribution towards the Determination of the Optimum Load. Journal of the Royal Statistical Society, 91(2):256–257, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341537.

P:1928:BRBa

[759] G. A. P. Book review: Industry Prosperity and the Farmer, by Russell C. Engberg; Council and Staff of the Institute of Economics. Journal of the Royal Statistical Society, 91(2):257–259, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341538.

M:1928:BRBd

[760] H. W. M. Book review: The International Accounts, by Cleona Lewis. Journal of the Royal Statistical Society, 91 (2):259, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341539.

A:1928:BRB

[761] J. E. A. Book review: The War Finance of France: The War Expenditure of France, by Gaston Jeze; How France Met her War Expenditure, by Henri Truchy. Journal of the Royal Statistical Society, 91(2):259–262, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341540.

M:1928:BRBe

[762] H. W. M. Book review: On Stimulus in Economic Life, by Josiah Stamp. Journal of the Royal Statistical Society, 91 (2):262–263, ???? 1928. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341541.

Anonymous:1928:ONPb

[763] Anonymous. Other new publications. Journal of the Royal Statistical Society, 91(2):263–267, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341542.

Anonymous:1928:CNb

[764] Anonymous. Current notes. Journal of the Royal Statistical Society, 91(2): 268–277, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341543.

Anonymous:1928:YGS

[765] Anonymous. Yves Guyot (Sept. 6, 1843– Feb. 21, 1928). Journal of the Royal Statistical Society, 91(2):278–279, ???? 1928. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341544.

Anonymous:1928:SEAb

[766] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 91(2): 280–285, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341545.

Anonymous:1928:LALb

[767] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 91(2):286–292, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341546.

Anonymous:1928:PRb

[768] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 91(2): 293–302, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/

2341547. See corrigenda [769].

Anonymous:1928:CPR

[769] Anonymous. Corrigenda: Periodical returns. Journal of the Royal Statistical Society, 91(3):??, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341600. See [768].

Macrosty:1928:TGS

[770] Henry W. Macrosty. Trade and the gold standard. Journal of the Royal Statistical Society, 91(3):303–345, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341601.

Edge:1928:VRE

[771] P. Granville Edge. Vital registration in Europe. The development of official statistics and some differences in practice. Journal of the Royal Statistical Society, 91(3):346–393, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341602.

ES:1928:WPC

[772] Editor of the “Statist”. Wholesale prices of commodities in 1927. Journal of the Royal Statistical Society, 91(3):394–411, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341603.

C:1928:BRBc

[773] L. R. C. Book review: A Study in Public Finance, by A. C. Pigou. Journal of the Royal Statistical Society, 91(3):412–415, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341604.

H:1928:BRBg

[774] R. G. H. Book review: International Trade, by F. W. Taussig. Journal of the Royal Statistical Society, 91(3):415–416, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341605.

M:1928:BRBf

[775] H. W. M. Book review: Shipping Problems, 1916–21, by W. Palin Elderton; Cyclical Fluctuations in the Shipping and Shipbuilding Industries. by F. Cyril James. Journal of the Royal Statistical Society, 91(3):416–418, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341606.

J:1928:BRB

[776] W. H. J. Book review: Railway Statistics, their Compilation and Use, by A. E. Kirkus. Journal of the Royal Statistical Society, 91(3):418–420, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341607.

L:1928:BRB

[777] H. M. L. Book review: Central Banks: A Study of the Constitution of Banks of Issue, with an Analysis of Representative Charters, by C. H. Kisch; W. A. Elkin. Journal of the Royal Statistical Society, 91(3):420–422, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341608.

C:1928:BRBd

[778] N. E. C. Book review: Modern Monetary Systems, by Bertrand Nogaro. Journal of the Royal Statistical Society, 91(3): 422–423, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341609.

M:1928:BRBg

[779] J. M. Book review: The Migration of British Capital, to 1875, by Leland Hamilton Jenks. Journal of the Royal Statistical Society, 91(3):423–424, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341610.

B:1928:BRBb

[780] W. A. B. Book review: The British Employment Exchange, by John Barton Seymour. Journal of the Royal Statistical Society, 91(3):424–426, ???? 1928. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341611.

C:1928:BRBe

[781] L. R. C. Book review: Thr Properity of Australia: An Economic Analysis, by Frederic C. Benham. Journal of the Royal Statistical Society, 91(3):426, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341612.

H:1928:BRBh

[782] R. G. H. Book review: The Economic Impact of America, by George Peel. Journal of the Royal Statistical Society, 91(3):427–428, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341613.

P:1928:BRBb

[783] G. A. P. Book review: Russian Economic Development since the Revolution, by Maurice Dobb; H. S. Stevens. Journal of the Royal Statistical Society, 91(3):428–430, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341614.

Anonymous:1928:ONPc

[784] Anonymous. Other new publications. Journal of the Royal Statistical Society, 91(3):430–434, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341615.

Anonymous:1928:CNc

[785] Anonymous. Current notes. Journal of the Royal Statistical Society, 91(3): 435–446, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341616.

R:1928:LE

[786] R. H. R. Lord Eversley. Journal of the Royal Statistical Society, 91(3):447–448, ???? 1928. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341617.

Anonymous:1928:SEAc

[787] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 91(3): 449–455, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341618.

Anonymous:1928:LALc

[788] Anonymous. List of additions to library. Journal of the Royal Statistical Society, 91(3):456–462, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341619.

Anonymous:1928:VIb

[789] Anonymous. Volume information. Journal of the Royal Statistical Society, 91 (4):627–632, ???? 1928. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341668.

Jones:1928:CLS

[790] D. Caradog Jones. The cost of living of a sample of middle-class families. Journal of the Royal Statistical Society, 91(4): 463–518, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341669.

Hilton:1928:SFE

[791] John Hilton. Some further enquiries by sample. Journal of the Royal Statistical Society, 91(4):519–540, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341670.

Jensen:1928:PS

[792] Adolph Jensen. Purposive selection.

Journal of the Royal Statistical Society, 91(4):541–547, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341671.

Rhodes:1928:NCF

[793] E. C. Rhodes. On the normal correlation function as an approximation to the distribution of paired drawings. Journal of the Royal Statistical Society, 91(4): 548–550, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341672.

Anonymous:1928:RCF

[794] Anonymous. Report of the Council for the financial year ending June 19, 1928, presented at the ninety-fourth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 19, 1928. Journal of the Royal Statistical Society, 91(4):551–566, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341673.

Anonymous:1928:PNF

[795] Anonymous. Proceedings of the ninetyfourth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 19, 1928. Journal of the Royal Statistical Society, 91(4):567–568, ???? 1928. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341674.

R:1928:BRBd

[796] E. C. R. Book review: Probability and its Engineering Uses, by Thornton C. Fry. Journal of the Royal Statistical Society, 91(4):569–572, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341675.

H:1928:BRBi

[797] J. H. Book review: The Cost of Living in Foreign Countries. Journal of the Royal Statistical Society, 91(4):572, ???? 1928. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341676.

B:1928:BRBc

[798] W. A. B. Book review: Wages and Hours of Work in the Coal-Mining Industry. Journal of the Royal Statistical Society, 91(4):573–574, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341677.

T:1928:BRBa

[799] C. T. Book review: The British Coal Dilema, by Isador Lubin; Helen Everett. Journal of the Royal Statistical Society, 91(4):574–575, ???? 1928. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341678.

G:1928:BRBb

[800] J. B. G. Book review: British Food Control, by William H. Beveridge. Journal of the Royal Statistical Society, 91(4): 575–578, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341679.

M:1928:BRBh

[801] H. W. M. Book review: The Reserve Banks and the Money Market, by W. Randolph Burgess. Journal of the Royal Statistical Society, 91(4):578–580, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341680.

C:1928:BRBf

[802] L. R. C. Book review: Economics of Instalment Selling: A Study in Consumers’ Credit, with Special Reference to the Automobile, by Edwin R. A. Seligman. Journal of the Royal Statistical Society, 91(4):581–584, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341681.

R:1928:BRBe

[803] E. C. R. Book review: La Statistique et sa Methode, by Lucien March; Differences et Correlation en Statistique, by Lucien March. Journal of the Royal Statistical Society, 91(4):584–585, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341682.

C:1928:BRBg

[804] N. E. C. Book review: Behaviour of Prices, by F. C. Mills. Journal of the Royal Statistical Society, 91(4):585–586, ???? 1928. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341683.

T:1928:BRBb

[805] C. T. Book review: Politique et Theories Monetaires Abglaises d’Apres-Guerre, by Jean-Pierre Lazard. Journal of the Royal Statistical Society, 91(4): 586–587, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341684.

C:1928:BRBh

[806] C. E. C. Book review: Lancashire under the Hammer, by B. Bowker; Great Britain from Adam Smith to the Present Day: An Economic and Social Survey, by C. R. Fay. Journal of the Royal Statistical Society, 91(4):587–588, ???? 1928. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341685.

G:1928:BRBc

[807] A. P. L. G. Book review: The Evolution of Industrial Organization, by B. F. Shields. Journal of the Royal Statistical Society, 91(4):588–589, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341686.

Anonymous:1928:ONPd

[808] Anonymous. Other new publications. Journal of the Royal Statistical Society, 91(4):589–592, ???? 1928. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341687.

Anonymous:1928:CNd

[809] Anonymous. Current notes. Journal of the Royal Statistical Society, 91(4): 593–605, ???? 1928. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341688.

Anonymous:1928:SEAd

[810] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 91(4): 606–610, ???? 1928. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341689.

Anonymous:1928:LALd

[811] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 91(4):611–616, ???? 1928. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341690.

Anonymous:1928:ALA

[812] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 91(4):617–625, ???? 1928. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341691.

Anonymous:1929:FMa

[813] Anonymous. Front matter. Journal of the Royal Statistical Society, 92(1):i–viii, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341434.

Flux:1929:NI

[814] A. W. Flux. The national income. Journal of the Royal Statistical Society, 92 (1):1–25, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341435.

Anonymous:1929:PM

[815] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 92(1):26–33, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341436.

Soper:1929:IPD

[816] H. E. Soper. The interpretation of periodicity in disease prevalence. Journal of the Royal Statistical Society, 92 (1):34–73, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341437.

Clark:1929:GAU

[817] C. G. Clark. A graphical analysis of the unemployment position. 1920–1928. Journal of the Royal Statistical Society, 92(1):74–99, ???? 1929. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341438.

G:1929:BRBa

[818] M. G. Book review: Statistique Mathematique, by G. Darmois. Journal of the Royal Statistical Society, 92(1):100–101, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341439.

I:1929:BRB

[819] J. O. I. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 92(1):101–103, ???? 1929. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341440.

E:1929:BRBa

[820] W. P. E. Book review: Population Statistics and their Compilation, by Hugh H. Wolfenden. Journal of the Royal Statistical Society, 92(1):103–104, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341441.

C:1929:BRBa

[821] L. R. C. Book review: Statistical Laws of Demand and Supply, with Special Application to Sugar, by Henry Schultz. Journal of the Royal Statistical Society, 92 (1):104–106, ???? 1929. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341442.

H:1929:BRBa

[822] A. B. H. Book review: Studies in Deceit, by Hugh Hartshorne; Mark A. May. Journal of the Royal Statistical Society, 92(1):106–107, ???? 1929. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341443.

N:1929:BRB

[823] E. M. N. Book review: Twin Births and Twins from a Hereditary Point of View, by Gunnar Dahlberg. Journal of the Royal Statistical Society, 92(1):108–109, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341444.

H:1929:BRBb

[824] W. H. H. Book review: Report of the Committee on Vaccination. Ministry of Health, by His Majesty’s Stationary Office. Journal of the Royal Statistical Society, 92(1):109–110, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341445.

A:1929:BRBa

[825] J. E. A. Book review: The Economics of Rail Transport in Great Britain, by C. E. R. Sherrington; Guy Granet. Journal of the Royal Statistical Society, 92(1): 110–112, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341446.

H:1929:BRBc

[826] A. B. H. Book review: Industrial Tuneside, by Henry A. Mess. Journal of the Royal Statistical Society, 92(1):112–114, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341447.

B:1929:BRBa

[827] W. A. B. Book review: British Engineering Wages, by Robert S. Spicer. Journal of the Royal Statistical Society, 92(1):115–116, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341448.

B:1929:BRBb

[828] W. A. B. Book review: The Wages of Unskilled Labour in Manufacturing Industries in the United States, 1890– 1924, by Whitney Coombs. Journal of the Royal Statistical Society, 92(1): 116–117, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341449.

C:1929:BRBb

[829] L. R. C. Book review: Buying Power of Labour and Post-War Cycles, by Asher Achinstein. Journal of the Royal Statistical Society, 92(1):117–118, ???? 1929. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341450.

B:1929:BRBc

[830] W. A. B. Book review: Coal and its Conflicts, by John R. Raynes. Journal of the Royal Statistical Society, 92(1):118, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341451.

E:1929:BRBb

[831] P. G. E. Book review: La Population et les Traqces de Voies a Paris depuis un Siecle, by Maurice Halbwachs. Journal of the Royal Statistical Society, 92 (1):119, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341452.

Anonymous:1929:ONPa

[832] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 92(1):119–123, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341453.

Anonymous:1929:CNa

[833] Anonymous. Current notes. Journal of the Royal Statistical Society, 92(1): 124–139, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341454.

Anonymous:1929:SEAa

[834] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 92(1): 140–146, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341455.

Anonymous:1929:LALa

[835] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 92(1):147–155, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341456.

Anonymous:1929:PRa

[836] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 92(1): 156–162, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341457.

Anonymous:1929:BMa

[837] Anonymous. Back matter. Journal of the Royal Statistical Society, 92(1):??, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341458.

Anonymous:1929:FMb

[838] Anonymous. Front matter. Journal of the Royal Statistical Society, 92(2):i–viii, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341505.

Flux:1929:DNI

[839] A. W. Flux. Discussion on the national income. Journal of the Royal Statistical Society, 92(2):163–182, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341506. See corrigendum

[882].

Hill:1929:ISV

[840] A. Bradford Hill. An investigation of sickness in various industrial occupations. Journal of the Royal Statistical Society, 92(2):183–238, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341507. See corrigendum

[906].

Anonymous:1929:WPC

[841] Anonymous. Wholesale prices of commodities in 1928. Journal of the Royal Statistical Society, 92(2):239–255, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341508.

Griffith:1929:RSS

[842] G. Talbot Griffith. Rickman’s second series of Eighteenth-Century population figures. Journal of the Royal Statistical Society, 92(2):256–263, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341509.

Irwin:1929:NTG

[843] J. O. Irwin. Note on the χ^{2} test for goodness of fit. Journal of the Royal Statistical Society, 92(2):264–266, ????

1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341510.

G:1929:BRBb

[844] M. G. Book review: Inleiding tot de Beoefening der Statistiek, by C. A. Verrijn Stuart. Journal of the Royal Statistical Society, 92(2):267–268, ???? 1929. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341511.

B:1929:BRBd

[845] A. L. B. Book review: Cours de Statistique, professé en 1927–8 à la Faculté de Droite. Journal of the Royal Statistical Society, 92(2):268–269, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341512.

M:1929:BRBa

[846] H. W. M. Book review: Some Economic Factors in Modern Life, by Josiah Stamp. Journal of the Royal Statistical Society, 92(2):269–271, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341513.

W:1929:BRB

[847] A. W. W. Book review: The Balance of Births and Deaths. Vol. I, Western and Northern Europe, by Robert R. Kuczynski. Journal of the Royal Statistical Society, 92(2):271–272, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341514.

H:1929:BRBd

[848] A. B. H. Book review: American Marriage and Family Relationships, by E. R. Groves; W. R. Ogburn. Journal of the Royal Statistical Society, 92(2):272–273, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341515.

F:1929:BRB

[849] A. W. F. Book review: The Board of Trade, by Hubert Llewellyn Smith. Journal of the Royal Statistical Society, 92(2):273–274, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341516.

M:1929:BRBb

[850] H. W. M. Book review: A History of Prices and of the State of the Circulation from 1792 to 1856, by Thomas Tooke; William Newmarch; An Introduction to Tooke and Newmarch’s A History of Prices, by T. E. Gregory. Journal of the Royal Statistical Society, 92(2):274–275, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341517.

M:1929:BRBc

[851] H. W. M. Book review: Select Statutes, Documents and Reports Relating to British Banking, 1832–1928, by T. E. Gregory; E. Cassel. Journal of the Royal Statistical Society, 92(2):275–277, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341518.

K:1929:BRB

[852] A. E. K. Book review: The Railway Policy of South Africa, by S. Herbert Frankel. Journal of the Royal Statistical Society, 92(2):277–278, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341519.

A:1929:BRBb

[853] J. E. A. Book review: The Problem of Motor Transport, by Christopher T. Brunner. Journal of the Royal Statistical Society, 92(2):278–280, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341520.

Anonymous:1929:ONPb

[854] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 92(2):280–283, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341521.

Anonymous:1929:CNb

[855] Anonymous. Current notes. Journal of the Royal Statistical Society, 92(2): 284–296, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341522.

Anonymous:1929:SRH

[856] Anonymous. Sir R. Henry Rew. Journal of the Royal Statistical Society, 92(2): 297–299, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341523.

Anonymous:1929:AS

[857] Anonymous. Augustus Sauerbeck. Journal of the Royal Statistical Society, 92 (2):300, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341524.

Anonymous:1929:SEAb

[858] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 92(2): 301–304, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2341525.

Anonymous:1929:LALb

[859] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 92(2):305–311, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341526.

Anonymous:1929:PRb

[860] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 92(2): 312–321, ???? 1929. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2341527.

Anonymous:1929:BMb

[861] Anonymous. Back matter. Journal of the Royal Statistical Society, 92(2):??, ???? 1929. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2341528.

Anonymous:1929:FMc

[862] Anonymous. Front matter. Journal of the Royal Statistical Society, 92(3):i–viii, ???? 1929. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2342235.

Snow:1929:LIE

[863] E. C. Snow. The limits of industrial employment. Journal of the Royal Statistical Society, 92(3):323–373, ???? 1929. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342236.

Bresciani-Turroni:1929:MWG

[864] C. Bresciani-Turroni. The movement of wages in Germany during the depreciation of the mark and after stabilization. Journal of the Royal Statistical Society, 92(3):374–427, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342237.

Fraser:1929:NZI

[865] Malcolm Fraser. New Zealand — infant mortality rates and still-births. Journal of the Royal Statistical Society, 92(3): 428–444, ???? 1929. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2342238.

Soper:1929:GSD

[866] H. E. Soper. The general sampling distribution of the multiple correlation coefficient. Journal of the Royal Statistical Society, 92(3):445–447, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342239.

B:1929:BRBe

[867] J. B. Book review: The Shadow of the World’s Future, or the Earth’s Population Possibilities and the Consequences of the Present Rate of Increase of the Earth’s Inhabitants, by George Handley Knibbs. Journal of the Royal Statistical Society, 92(3):448–449, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342240.

C:1929:BRBc

[868] L. R. C. Book review: Trade and Credit, by R. G. Hawtrey. Journal of the Royal Statistical Society, 92(3):450–451, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342241.

C:1929:BRBd

[869] L. R. C. Book review: Considerazioni sui “Barometri Economici.”, by C. Bresciani Turroni. Journal of the Royal Statistical Society, 92(3):451–452, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342242.

A:1929:BRBc

[870] J. E. A. Book review: A Study of Interest Rates, by Karin Kock. Journal of the Royal Statistical Society, 92(3):452–453, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2342243.

H:1929:BRBe

[871] J. H. Book review: The Problem of Trust and Monopoly Control, by A. P. L. Gordon. Journal of the Royal Statistical Society, 92(3):453–454, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342244.

V:1929:BRBa

[872] J. W. V. Book review: The Problem of Industrial Relations and Other Lectures, by Henry Clay. Journal of the Royal Statistical Society, 92(3):454–457, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342245.

B:1929:BRBf

[873] W. A. B. Book review: Wages in Practice and Theory, by J. W. F. Rowe. Journal of the Royal Statistical Society, 92 (3):457–458, ???? 1929. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342246.

V:1929:BRBb

[874] J. W. V. Book review: The Control of Wages, by Walton Hamilton; Stacey May. Journal of the Royal Statistical Society, 92(3):458, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342247.

B:1929:BRBg

[875] W. A. B. Book review: The Unemployed: Old Policies and New, by Ronald C. Davison. Journal of the Royal Statistical Society, 92(3):458–460, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342248.

V:1929:BRBc

[876] J. W. V. Book review: The Clothing Workers of Great Britain, by S. P.

Dobbs; Sidney Webb. Journal of the Royal Statistical Society, 92(3):460–461, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2342249.

Anonymous:1929:ONPc

[877] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 92(3):461–463, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342250.

Anonymous:1929:CNc

[878] Anonymous. Current notes. Journal of the Royal Statistical Society, 92(3): 464–473, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2342251.

E:1929:FGP

[879] W. P. E. Francis G. P. Neison. Journal of the Royal Statistical Society, 92(3):473, ???? 1929. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2342252.

Anonymous:1929:SEAc

[880] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 92(3): 474–478, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2342253.

Anonymous:1929:LALc

[881] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 92(3):479–485, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342254.

Anonymous:1929:CDN

[882] Anonymous. Corrigendum: Discussion on the national income. Journal of the Royal Statistical Society, 92(3):??, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342255. See

[839].

Anonymous:1929:VI

[883] Anonymous. Volume information. Journal of the Royal Statistical Society, 92 (4):663–vii, ???? 1929. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2342212.

Anonymous:1929:FMd

[884] Anonymous. Front matter. Journal of the Royal Statistical Society, 92(4):i–viii, ???? 1929. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2342213.

Loveday:1929:MTL

[885] A. Loveday. The measurement of tariff levels. Journal of the Royal Statistical Society, 92(4):487–529, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342214.

Llewellyn-Smith:1929:NSL

[886] H. Llewellyn-Smith. The new survey of London life and labour. Journal of the Royal Statistical Society, 92(4):530–558, ???? 1929. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2342215.

Faber:1929:NMC

[887] Harald Faber. A new method of comparing the productivity of crops on arable land in England and Wales, Scotland and Denmark. Journal of the Royal Statistical Society, 92(4):559–579, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342216.

Irwin:1929:FDN

[888] J. O. Irwin. On the frequency distribution of any number of deviates from the mean of a sample from a normal population and the partial correlations between them. Journal of the Royal Statistical Society, 92(4):580–584, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342217.

Coates:1929:DTT

[889] W. H. Coates. Double taxation and tax evasion. Journal of the Royal Statistical Society, 92(4):585–593, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342218.

Anonymous:1929:RCF

[890] Anonymous. Report of the Council for the financial year ended December 31, 1928, and for the sessional year ending June 18, 1929, presented at the ninety-fifth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 18, 1929. Journal of the Royal Statistical Society, 92(4):594–609, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342219.

Anonymous:1929:PNF

[891] Anonymous. Proceedings of the ninetyfifth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 18, 1929. Journal of the Royal Statistical Society, 92(4):610, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342220.

H:1929:BRBf

[892] A. B. H. Book review: Introduction to the Mathematics of Statistics, by R. W. Burgess; J. Wesley Young. Journal of the Royal Statistical Society, 92(4): 611–612, ???? 1929. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2342221.

S:1929:BRB

[893] E. C. S. Book review: Migration and Business Cycles, by H. Jerome; Wesley C. Mitchell. Journal of the Royal Statistical Society, 92(4):612–614, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342222.

C:1929:BRBe

[894] N. E. C. Book review: The Variability of Prices, by Lodewijk Hamburger. Journal of the Royal Statistical Society, 92(4):614–616, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342223.

B:1929:BRBh

[895] C. F. B. Book review: Taxation in the Modern State, by Alzada Comstock. Journal of the Royal Statistical Society, 92(4):616–618, ???? 1929. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342224.

G:1929:BRBc

[896] A. P. L. G. Book review: Germany Under the Dawes Plan, by Max Sering; Dr. Rer. Polit. Journal of the Royal Statistical Society, 92(4):618–619, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342225.

A:1929:BRBd

[897] J. E. A. Book review: Transfert des Reparations et le Plan Dawes, by Harry D. Gideonse. Journal of the Royal Statistical Society, 92(4):619–620, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342226.

P:1929:BRB

[898] G. A. P. Book review: The Mexican Agrarian Revolution, by Frank Tannenbaum; Die Bauernbewegung in der Russischen Revolution, 1917, by von S. Dubrowski. Journal of the Royal Statistical Society, 92(4):621–622, ???? 1929. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342227.

Anonymous:1929:ONPd

[899] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 92(4):622–625, ???? 1929. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342228.

Anonymous:1929:CNd

[900] Anonymous. Current notes. Journal of the Royal Statistical Society, 92(4): 626–640, ???? 1929. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2342229.

Anonymous:1929:SAE

[901] Anonymous. Sir Alfred Edmund Bateman, K.C.M.G. Journal of the Royal Statistical Society, 92(4):641–643, ???? 1929. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342230.

Anonymous:1929:SEAd

[902] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 92(4): 644–647, ???? 1929. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2342231.

Anonymous:1929:LALd

[903] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 92(4):648–651, ???? 1929. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342232.

Anonymous:1929:ALA

[904] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 92(4):652–661, ???? 1929. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342233.

Anonymous:1929:BMc

[905] Anonymous. Back matter. Journal of the Royal Statistical Society, 92(4):??, ???? 1929. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2342234.

Anonymous:1930:CIS

[906] Anonymous. Corrigendum: An investigation of sickness in various industrial occupations. Journal of the Royal Statistical Society, 93(1):??, ???? 1930. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342144. See [840].